A Complete Bibliography of Publications in the Journal of Fish Biology (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.06

Title word cross-reference

3 [619]. _{2} [894]. _{4} [414]. α [365]. β [545]. × [130].

-Glucose [76]. -methyltestosterone [365].

0-group [681, 297, 196].

1254 [754]. 1970 [48]. 1974 [204].

2 [805]. 222 [561].

3- [564].

4 [805]. 4-D [805]. 48/80 [775].

5-hydroxydopamine [755].

6-hydroxydopamine [755].

71 [94].

= [56]. =Pseudocrenilabrus [387].

A. [860, 387]. Abate [917, 776]. Aberdeenshire [441]. abilities [105]. ability [245, 387, 354]. Abnormalities [478, 318]. Abramis [511, 318, 85, 766, 767]. Absence [194]. absorption [194]. Abstracts [48]. abundance [575, 600, 420, 254]. acadja [102]. Acanthocephala [523]. acanthocephalan [303, 317, 740]. Acanthocolpidae [357]. accessory [815]. acclimation [330]. acclimatized [799]. Accumulation [776, 741, 669, 512, 728]. acetate [243]. acetylcholine [772]. acid [902, 650, 913, 812, 562, 696, 808, 848, 536]. acid-Schiff [650]. acids [15, 611]. Acquired [19, 288]. action [507, 832]. activated [741]. activities [386, 637, 839, 884]. activity [835, 707, 746, 932, 593, 520, 405, 648, 886, 504, 58, 708, 621, 711, 623, 842, 334, 329, 449, 411, 309, 310, 311, 527, 56]. aculeatus [499, 736, 109, 228, 555, 413, 501, 563, 598, 597, 858, 167, 238, 456, 680]. acus [41]. Acute [805, 175, 539, 635, 789, 27, 574, 419, 817, 73]. acutely [657]. Adaptation [621, 859, 898, 903]. adaptations [594, 431]. adapted [634]. adjacent [151]. adjuvants [883]. administration [873]. Adrenaline [106]. adult [713, 593, 55, 901, 465, 703, 568, 229, 873, 582]. advantages [80]. Advisory [233, 248]. aeglefinus [722, 461, 631, 263]. aeromonad [142]. Aeromonas [908, 414, 647]. affecting [2]. affinis [608, 224]. Africa [399, 102, 322, 379, 299, 500, 355]. African [375, 150, 152, 831, 332, 257, 869, 453]. after [170, 792, 606, 817, 930]. afzeliusi [446]. against [675]. Age [280, 587, 803, 441, 401, 554, 757, 355, 412, 742, 495, 850, 454, 325, 514, 901, 240, 276, 435, 443, 855, 52, 223, 408]. aggregata [758]. aggregations [304]. Aggressive [919]. Agonistic [901]. Agosia [789]. air [68, 815, 899, 849, 840, 573]. air-breathing [815, 899, 849, 840, 573]. air-economizer [68]. Al [746]. Alaw [133, 139, 140]. albella [332]. albida [91]. albino [60]. albolineatus [700]. albula [82, 862]. Alburnus [805, 269, 762]. Alectis [301]. aleuticm [356]. alevins [783, 829]. Alexandria [601]. alimentary [177, 88, 89, 263, 654]. Alkaline [886]. alleviation [871]. Allocreadium [800]. allograft [126]. along [368]. Alosa [394, 668, 710]. alpinus [797, 600, 719, 761, 661, 259, 369, 345, 265, 567]. alterations [31]. altivelis [308]. ambivalent [36]. America [162]. American [394, 454, 710, 465, 688]. americanus [56]. Amino [562, 848, 696, 536]. ammocoete [842]. Ammodytes [309, 310, 311]. ammonia [935, 466, 657]. among [417]. amongst [916]. Amouncements [712]. amount [697]. Amphipnoidae [573]. Amphipnous [573]. Amylase [859]. Anabas [788]. anadromous [369, 345]. anaerobic [913]. anaesthesia [768, 529]. anaesthetic [532, 682, 927, 928, 929]. analogues [798]. analyses [255, 503]. Analysis [900, 664, 282, 104, 583, 197, 51, 570, 739, 610, 215, 919, 745, 874]. anatomical [620]. anatomy [88, 654, 814]. anchoita [201]. anchovy [201]. andersoni [818]. androgenetic [516]. androgens [11]. Anglesey [133, 139, 140]. angling [139]. Anguilla [860, 571, 601, 770, 501, 540, 378, 605, 521, 912, 645, 185]. anguillarum [583, 521]. angustimanus [339]. animals [227]. anionic [507]. anisakids [931]. Anisakis [151, 649]. Announcement [157, 249, 333]. Announcments [306]. Annual [835, 746, 680, 863, 136, 933, 311]. Antarctic [692]. antibiotic [873]. antibodies [99]. antibody [890, 219]. antibody-producing [890]. antigen [504, 822]. antigen-stimulated [822]. antigenic [189]. Antigenicity [125]. antigens [526, 606]. Apeltes [858]. aphyiine [198]. Aplocheilus [800]. apoda [72]. apparatus [68, 546, 538]. apparent [219, 832]. application [231, 149]. appraisal [233, 248]. Approaches [618, 610]. April [48]. aquaria [409, 769]. Aquarium [351, 227]. aquatic [279, 907, 351, 34]. Arab [746]. arasites [733]. architecture [550]. Archosargus [342]. Arctic [761, 901, 259, 369, 345]. arcticus [901]. area [418, 215, 259, 369, 345, 201]. areas [223]. arenatus [216]. Argentine [280, 201]. Argulus [61]. Argyll [485, 488]. Argyripnus [127]. Aristichthys [656]. armatus [251]. Aroclor [754]. arsenate [440]. artedii [920]. arteriosus [343, 489]. arthropod [579]. Arthur [692]. artificial [19, 176, 595]. artificially [179]. Ascanius [587]. Ascaridata [151]. Ascaridida [649]. Aspects [669, 622, 379, 918, 356, 461, 479, 729, 660, 239, 906, 343, 724, 542, 141, 357, 367, 525, 704, 524, 153, 858, 242, 16, 124]. assessing [643]. assessment [155, 390, 927, 928, 929]. assimilation [260]. associated [343, 549, 67, 230, 756]. Association [13]. assumption [283]. asynchronous [716]. atherinoides [35]. Atlantic [459, 642, 421, 234, 315, 625, 495, 759, 419, 786, 423, 32, 736, 631, 704, 703, 120, 672, 695, 368, 595, 702, 773, 706, 804, 833, 821, 148, 149, 888, 551]. atlanticus [127, 398]. ATP [929]. atraria [273]. attempt [45]. Audernaerde [446]. aurata [659, 930]. auratus [517, 171, 472, 385, 195, 895, 666, 619]. aurea [634, 388]. Australia [569, 693, 860]. Australian [872, 8]. australis [860]. auto [99]. auto-antibodies [99]. automated [874]. automatic [784]. autumn [934, 905, 80]. Availability [520, 309]. Avoidance [809, 19]. Avon [383, 317, 523]. axial [651]. ayu [308].

B [798]. baccatum [357]. bacterial [570]. Bacteriological [724]. bacteriophage [883]. Baffin [259, 369, 345]. Bagrus [299, 401]. Baird [608, 224]. baits [19]. Bala [346, 78]. balance [308, 45]. Balaton [122]. Bald [14]. Balistes [765]. ballan [587]. Baltic [452]. band [486]. banded [430]. Bangladesh [852]. barbatulus [397, 383]. barbel [473]. Barbus [835, 746, 473, 277, 278, 372, 176]. Barents [882]. based [407]. basic [283, 879]. Basilichthys [280]. basis [484, 501]. Bass [904, 810, 686, 161, 105, 896, 493, 458, 596, 584, 738, 832]. Batoidei [876]. batrachus [815, 899]. Bay [297, 493, 774]. be [185]. beach [655]. Bear [761]. Behaviour [692, 231, 473, 324, 889, 901, 253, 919, 672, 548, 738, 242, 103, 36, 77]. Behavioural [574, 309, 310, 311, 620, 192]. Beneden [461]. benthic [431, 374]. benzocaine [529]. bergylta [587]. Bering [674]. between [222, 735, 591, 351, 402, 125, 126, 170, 8, 720, 540, 120, 623, 695, 617, 502]. bibliography [119, 552]. big [216]. big-eye [216]. bighead [656]. bile [611]. Bimodal [595]. binding [915]. bio [548]. bio-energetics [548]. Biochemical [780, 148, 149, 589, 188, 506, 730, 832]. bioelectric [843]. Biological [13, 96, 726, 685, 882, 631]. Biology [35, 865, 557, 384, 844, 866, 669, 18, 128, 868, 622, 379, 715, 434, 830, 491, 150, 37, 683, 918, 588, 251, 337, 498, 454, 317, 749, 750, 751, 542, 141, 639, 685, 661, 882, 533, 431, 461, 357, 367, 292, 525, 841, 345, 479, 743, 487, 560, 422, 543, 334, 335, 336, 729, 660, 239, 906]. biomass [618]. Biometric [799]. Biosociology [266]. biphenyls [753]. bis [574]. bisexual [716]. bispinis [692]. blackfly [917, 776]. Blackwater [427]. bladder [416]. bleak [805, 269, 762]. Bleeker [332]. blennies [477]. Blennim [477]. Blennius [682, 731, 848]. Blgr [491]. Bloch [608, 22, 7, 840, 924, 892, 41, 11, 550, 536, 284, 326]. Blood [537, 26, 414, 817, 244, 894, 565, 652, 810, 528, 250, 462, 152, 561, 811, 506, 798, 849, 840, 159, 353, 307, 836, 647, 327, 529, 745, 513, 250]. blue [806, 588, 535]. Bluefish [904, 315]. bluegill [648, 129, 889]. Blyth [700]. bodies [658]. Body [468, 4, 591, 405, 820, 886, 547, 237, 474, 756]. Bola [114, 97]. bonariensis [280]. bones [742]. Book [564, 12, 24, 38, 49, 57, 74, 83, 95, 117, 134, 144, 156, 172, 191, 203, 232, 247, 262, 274, 305, 319, 344, 361, 377, 404, 415, 428, 451, 460, 469, 494, 519, 530, 576, 616, 644, 684, 727, 795, 925, 936, 681, 755]. borne [539]. Borth [199]. Bothnian [774]. Bothriocephalus [406, 62]. bottom [819]. bottom-living [819]. Boulenger [387, 630, 591, 723, 713, 574, 531, 841, 480]. Bowdich [150]. box [166]. Brachydanio [718, 700, 533, 891, 721]. bradycardia [411]. brama [511, 318, 85, 766, 767]. branchialis [791]. Branchiura [61]. branding [348]. bream [511, 318, 85, 766, 767, 342, 930]. breathing [843, 815, 899, 849, 840, 573]. bred [659, 721]. Breeding [434, 725, 520, 901, 841, 252]. breviceps [334, 335, 336]. brevis [286]. brief [112]. brill [62]. Britain [364, 282, 525, 28]. British [47, 94, 880, 115, 421, 151, 726, 313, 739, 420, 328, 447, 113, 164, 350]. broodfish [168]. brook [652, 110, 7, 635, 922, 522, 417, 647]. brown [507, 222, 76, 735, 50, 599, 475, 914, 419, 197, 512, 728, 133, 526, 606, 822, 445, 739, 740, 137, 779, 881, 919, 883, 510, 467, 672, 265, 572, 689, 548, 426, 813, 505, 80, 142, 267, 268, 441]. Brunswick [394]. bubalis [71]. bubble [490]. Buchanan [559, 891]. budget [474, 45]. buffer [590]. bulbus [343, 489]. bullhead [646, 383]. bully [334, 335, 336]. burbot [28]. Burchell [252, 355]. Burma [97]. Burn [599]. burrows [486]. butyltin [574]. Bystrice [41].

C [924]. C. [920]. Cadmium [783, 747, 397, 175, 648, 889, 472, 563, 598, 829]. caesium [512, 728]. calcitonin [892]. calcium [892, 510]. California [338, 339, 420, 410]. Caligidae [442, 556]. Caligidea [161]. Callionymus [918]. camssim [130]. Canada [35, 733, 137, 579]. canal [89, 601, 654]. candidates [375]. canicula [933]. capacity [231, 590]. capelin [882]. capensis [296, 293, 537, 217]. capito [58]. captivity [722]. capture [283, 370]. capture-recapture [283]. Carangidae [301]. Caranx [301]. Carassim [195]. Carassius [517, 188, 130, 171, 472, 385, 895, 666, 619]. carbohydrate [90]. carcinoma [788]. cardiac [548]. Cardiganshire [199, 128]. Caribbean [165, 479, 660]. carnivorous [88]. carolinensis [765]. carotenoid [665]. carp [283, 463, 364, 321, 30, 852, 227, 409, 478, 218, 270, 287, 288, 110, 325, 514, 188, 633, 130, 801, 838, 893, 406, 200, 734, 603, 691, 754, 884, 75, 516, 374, 235, 656, 594, 604]. carpio [283, 463, 351, 270, 287, 288, 110, 325, 633, 130, 801, 838, 406, 200, 734, 754, 884, 374]. cartilaginous [673]. case [414]. Caspian [122]. catch [1, 900]. catching [2]. catecholamines [772, 558]. catfish [564, 437, 464, 798, 815, 899, 11, 439, 316, 550, 869, 814, 401, 284, 757, 617, 513]. catfishes [531, 688]. Catla [559]. Catostomus [775, 825]. caudal [846]. caught [185]. causative [498]. caused [419, 862]. cavity [4]. CCl [414]. Cell [849, 825, 343, 798, 922, 195]. Cells [652, 890, 517, 634, 433, 250, 197, 910, 926, 662, 771, 896, 265, 567]. cellular [526, 606, 513]. cellulase [623, 329, 903]. Celtic [436]. Celyn [132]. centracanthid [53]. central [122, 854, 551]. cepedianum [861, 807]. cephalus [205, 865, 209, 31, 443, 549, 743, 487, 560, 211, 210, 212, 330, 396, 762, 663]. cephatus [44]. Cepola [486]. cercaria [357]. cerebralis [498, 534, 535]. certain [652, 675]. Cestoda [4, 685, 406, 461, 367, 62]. cestodarian [578]. cestode [882]. chair [797]. chalk [505]. chalkstream [705]. chamber [256]. Changes [818, 887, 237, 647, 223, 327, 289, 756, 835, 746, 780, 591, 43, 593, 106, 667, 20, 481, 55, 749, 750, 606, 188, 523, 905, 566, 67, 699, 540, 298, 508, 572, 440, 284, 709, 65]. Channa [849, 840, 924, 892]. Channel [42, 564, 437, 798, 439, 316, 814, 809, 617, 513, 292]. Chanos [730]. Chaparrudo [427]. chaptis [114, 97]. char [600, 719, 761, 259, 369, 345, 265, 567]. characin [162]. Characoidei [162]. Characteristics [640, 195, 599, 250, 365, 726]. characters [446]. Charr [629]. check [136, 757]. check’formation [643]. checklist [313]. chelo [359]. chemical [677, 652, 108, 743, 614]. chemosenses [827]. Chesapeake [493]. Cheshire [687]. Chew [59]. Chile [280, 931]. Chimaera [673]. Chinese [325, 514]. Chinook [569]. Chloramine [246]. Chloramine-T [246]. chloramines [635]. Chloride [771, 65, 634, 720]. chlorine [534]. cholesterol [611]. Chondrichthyes [876]. chorion [538]. chorionic [290, 308]. Chromis [551]. chromium [648]. Chromosome [93, 366]. chromosomes [854, 261]. chronic [229, 930]. chryselis [53]. Chrysichthys [464, 531]. chrysogaster [789]. chrysops [161]. Chu [91]. chub [443, 273, 762, 663]. chum [772]. Cichlasoma [725]. cichlid [844, 866, 387, 723, 794, 285, 725, 851, 554]. Cichlidae [65, 322, 379, 299, 399, 542, 81, 782]. cichlids [33]. ciliary [304]. Ciliata [669]. cinematographic [843]. Circadian [710]. circuit [784, 482]. circular [603]. circulating [671]. Cirrhinus [337]. cisco [920]. Citharinus [471]. citrate [801]. Cladocera [914]. Clarias [815, 899, 252, 355]. Clariidae [355]. clarki [168]. class [687, 803]. clavata [715, 639]. Clinostomum [686]. clomiphene [801]. closed [784, 295]. Clupea [452, 151, 314, 649]. clupeaformis [920]. Clupeidae [151]. clupeoid [371]. clupeoides [78]. cmentatus [10]. coagulation [565]. coalfield [813]. coarse [390, 916, 9, 178]. Coast [575, 918, 588, 628, 239, 500]. coastal [865, 202, 328, 845, 102]. cobitis [37]. cod [376, 275, 621]. Coho [772, 365, 915, 756, 824]. coibor [91]. Coitus [356]. cold [768]. Colisa [817]. collection [42]. Collett [407]. colour [253]. Columbia [420]. Comment [615]. commercially [151]. commersoni [775, 825]. Commission [248, 233]. Common [886, 318, 878, 325, 514, 130]. Communal [7]. communication [614, 373]. Comparative [462, 687, 506, 105, 633, 672, 327, 358]. comparing [492]. Comparison [771, 617, 349, 670, 512, 420, 181]. compensation [921]. Compensatory [121]. competition [610]. components [583, 832, 874]. Composition [34, 591, 468, 850, 237, 108, 254, 756]. concentration [642, 243, 481, 822, 307, 883, 439, 416]. concentrations [507, 864, 234, 648, 889, 110, 657, 817]. condition [497, 591, 751, 857, 332, 426, 543]. conditions [205, 792, 395, 70, 595, 241]. Conference [25, 204]. confined [569]. consequences [509]. Considerations [810, 453]. conspecific [8]. conspecifics [827]. constituents [840, 529]. construction [818]. consumption [84, 499, 682, 227, 129, 893, 336, 477]. containing [741]. contaminated [747]. contamination [549]. content [902, 222, 792, 665, 438, 21, 747, 536, 553]. contents [497, 676]. continuing [679]. Continuous [290]. Contribution [64, 137]. Control [534, 535, 246, 79, 32, 776, 760, 933, 311, 112]. conversion [314, 611, 720, 605, 586, 585]. convict [725]. Cooper [77]. Copepod [834, 748, 258, 429]. Copepoda [442, 556, 161]. Copper [521, 27, 175, 789, 778]. coregonid [282]. coregonids [202]. Coregoninae [670]. Coregonus [18, 82, 670, 481, 78, 862, 14, 403, 920]. corpuscular [929]. Corrib [885]. corticosteroid [11]. cortisol [481, 308]. Corvina [91]. Corynopoma [846]. Coryphaena [497]. Cost [456]. costs [610]. cottid [71]. Cottus [646, 383, 71]. counter [54, 123]. counters [784]. Course [200]. Cove [566, 577]. Cow [676]. crassum [661]. Creek [323, 693]. Creplin [600, 719, 187, 837]. Critical [358]. cross [351]. crucian [130, 188]. cruciata [161]. crude [724, 549]. Crustacea [61, 560]. Cryptobia [828]. Cryptocotyle [187, 837]. crysoleucas [545, 553, 527]. Ctenopharyngodon [30, 351, 227, 893, 603, 516, 235, 594, 604]. cuchia [573]. cuckoo [436]. culture [81, 671, 603, 691]. cultured [450, 21, 617]. Cumberland [259, 369, 345]. curves [929]. cut [168]. cut-throat [168]. Cuv. [280]. Cuvier [359, 337, 58, 286, 431, 216, 91]. cyanellus [440]. cyanocobalamin [808]. cycle [69, 436, 481, 294, 933, 311, 680, 863]. cycles [872, 886, 22, 112]. cyclical [746]. Cymatogaster [758]. Cynoscion [769]. cyprinacea [258]. cyprini [321]. cyprinid [718, 589, 219, 454]. Cyprinm [406, 884]. Cyprinodon [483]. Cyprinodontidae [312]. Cyprinus [283, 463, 351, 270, 287, 288, 110, 325, 633, 130, 801, 838, 200, 734, 754, 374]. cyst [31]. Cystidicola [774]. Cytogenetic [400]. Cytological [716]. Cytophotometric [400]. Czechoslovakia [41].

D [562, 805, 76]. D-Amino [562]. D. [600, 719]. dab [749, 750, 751, 547]. dace [189, 789, 276, 136, 273]. Dahomey [102]. Daily [582]. Dam [818]. damaged [826]. damnosum [776]. damsel [551]. Danio [721, 700]. Danish [522]. dark [621]. data [900, 879, 649]. date [69]. day [593]. decline [59]. Dee [408]. deep [245, 387, 354]. deepwater [254]. Defence [424]. deficiency [808, 924]. deformations [659]. deformities [721]. dehydrogenase [625]. demersal [520, 674, 544]. denaturation [503]. dendriticum [600, 719]. Dendrophysa [91]. dense [922]. Densities [277]. Density [638, 412, 870, 767, 911, 342, 514, 893, 467, 690, 242, 241, 476]. Density-dependent [638]. dentition [272]. dependant [594]. dependent [638]. depletion [713]. deposited [166]. deposition [405, 72]. depression [891]. Depth [794, 851, 256]. dermal [67, 300, 29, 46, 70, 99, 426]. dermis [161]. Dermocystidiosis [321]. Dermocystidium [321]. description [127, 402, 97]. desert [483]. detect [827, 99]. detection [309]. detergent [507]. detergents [279, 772]. determination [256, 803, 836, 757]. determining [592]. developed [819]. developing [671]. Development [253, 775, 806, 409, 22, 406, 209, 33, 665, 896, 558, 678, 332, 875, 169, 212, 73, 581]. developmental [805]. Devon [916, 445]. diadromous [434]. diagnostic [414]. diaphanus [430]. Dicentrachus [896]. Dicentrarchus [458]. Diclidophora [100]. dieffenbachii [185]. Diel [20, 708, 842]. Diet [208, 859, 898, 903, 741, 594, 488]. Dietary [785, 392, 720]. diets [365, 691, 34]. Differences [417, 609, 8, 716, 927, 928, 341, 160]. Different [838, 392, 877, 646, 781, 620, 223]. Digenea [357, 371, 187]. digeneans [263]. Digestion [907, 604, 71]. digestive [910, 926, 859, 898, 903, 782, 594]. dilution [745]. dimethylnitrosamine [539]. Dioctophymidae [257]. Diphyllobothrium [600, 719]. Diplostomum [715, 79, 639]. discharge [220, 876]. disease [452, 87, 321, 419, 101, 498, 694, 534, 535, 448, 14, 490, 16]. diseased [99]. diseases [674, 552, 493, 271]. disinfectants [675]. disinfection [534]. dispersion [766]. display [919, 261]. dissolved [935, 864, 228, 555]. distinction [222]. distinctions [502]. distinguishing [302]. Distribution [464, 896, 369, 412, 592, 82, 818, 575, 600, 303, 313, 445, 882, 109, 297, 674, 257, 485, 265, 429, 254, 595, 821]. disturbing [179]. ditremum [600, 719].

Diurnal [553, 405, 701]. DL [700]. DL-thyroxine [700]. DMN [539]. docmac [299]. dogfish [933]. dolphin [497]. dolphin-fish [497]. dominance [916]. Donovan [459]. Dorosoma [781, 861, 807, 827]. Dorset [324, 443]. dosages [854]. doses [930]. downstream [475]. dragonet [918]. drift [20]. drifting [475]. drugs [207]. drying [534]. duckweed [604]. duct [611]. due [321, 861]. during [713, 376, 382, 905, 465, 237, 716, 582, 411]. Dwythwch [672]. dynamic [832]. dynamics [719, 867, 339, 522, 335, 476]. dyscrasia [817]. dysfunction [287].

E. [661]. early [805, 806, 635, 518]. East [299, 399, 631, 500, 872, 831, 332, 257]. eastern [315, 551]. eating [653]. Ecological [399, 153]. ecology [693, 266, 686, 356, 199, 3, 338, 339, 479, 660, 752, 787]. economizer [68]. ecosystem [549, 738]. ectoparasites [362, 834]. ectoparasitic [459]. ectopic [170]. ectoprocta [853]. Editorial [698]. eel [571, 770, 573, 180, 378, 521, 912, 185]. Eels [501, 412, 448, 540]. Effect [565, 89, 665, 596, 690, 619, 935, 392, 812, 354, 630, 732, 806, 370, 561, 535, 138, 194, 801, 6, 892, 187, 444, 563, 567, 211, 466, 235, 778, 73, 309, 310, 311, 167, 476]. Effects [25, 590, 39, 308, 342, 700, 893, 654, 472, 90, 416, 912, 529, 449, 745, 581, 586, 585, 527, 917, 349, 866, 775, 790, 805, 772, 676, 870, 405, 545, 682, 207, 820, 722, 534, 514, 425, 696, 816, 547, 856, 188, 109, 228, 413, 678, 854, 510, 909, 875, 385, 637, 839, 753, 754, 927, 928, 929, 271, 457, 895, 515, 714, 544, 362, 666, 242, 480, 791, 755]. Efficacy [675]. efficiencies [492]. efficiency [596, 336, 586, 585]. effluent [752]. Egg [923, 166, 325, 167, 456]. eggs [69, 783, 290, 812, 806, 879, 72, 475, 166, 671, 235, 396, 201]. Egypt [601, 799]. elasmobranch [470]. elasmobranchs [570]. electric [370, 492, 876, 869, 143, 482]. electrically [86]. electrically-operated [86]. electricus [869]. electroanaesthesia [509]. electrolyte [472, 154]. Electron [597, 352, 770, 899, 316, 261]. electronic [123]. Electrophoresis [296, 929]. Electrophoretic [511, 503, 282, 670, 302, 53]. electrophoretically [255]. elements [513]. elevated [179]. elver [540]. elvers [601, 605]. Embiotocidae [352]. embryonic [875]. embryos [352, 170, 558, 584]. emerald [35]. emphasis [414]. emptying [820, 697]. Encapsulated [326]. endemic [338]. endocrine [933]. endothelium [343]. energetic [610]. energetics [548]. energy [713, 438, 474, 45]. England [427, 642, 646, 687, 850, 277, 278, 372, 240, 276, 435, 656]. English [292, 737, 762, 663]. Engraulis [201]. Enophrys [71]. enter [245, 387, 354]. entirely [923]. environment [41]. Environmental [112, 902, 894, 58, 510, 271, 311]. enzymatic [255]. Enzyme [386, 637, 839, 215, 884]. enzyme-gene [215]. enzymes [859, 898, 903, 414, 647, 753]. ephippiatus [186]. epidemic [142]. Epidermal [612, 662, 378]. epidermis [197, 265, 572, 567, 689]. Epithelial [788, 99]. epithelium [304, 389, 922]. equilibration [794, 851]. erinaceus [461, 631]. Erne [98, 679]. erosion [724]. Error [243]. erythrocyte [110, 21]. erythrocytes [76, 764, 296]. erythrodermatitis [364]. erythrophthalmus [511, 859, 898, 903, 298, 502]. esculenta [81]. esmarkii [557]. Esox [19, 87, 360, 88, 425, 696, 816, 435, 26, 909, 457]. español [118, 135, 146, 158]. Essex [427]. established [195]. establishment [192]. este [118, 135, 146, 158]. Esthwaite [502]. estimate [129]. estimating [618, 166]. Estimation [267]. estrogen [11]. estuaries [234, 706]. estuarine [921, 175, 704, 549, 501, 782, 358, 77, 837]. Estuary [427, 295, 368, 668, 484, 785, 791, 906]. etheostomid [215]. Ethmalosa [150]. ethylenediamine [243]. Etmopterus [651]. Eubothrium [685, 661, 882]. Euphrasen [71]. Europe [122, 153, 398, 688]. European [233, 248, 325, 514, 120, 378, 912, 10]. euryhaline [771]. eurythermal [921]. Eustrongylides [445, 257]. eutrophic [850]. evacuation [547, 856]. Evaluation [137, 102]. Evermann [664]. evidence [716, 400]. evolution [198]. examination [770]. examining [180]. exchange [312, 657]. excretion [564, 626, 596]. excretory [777]. exercise [780, 188, 444]. existence [421, 819, 120]. exogenous [11, 211, 553]. exotics [656]. experiment [182, 81]. Experimental [283, 476, 60, 395, 70]. Experiments [455, 360, 809]. exploitation [43]. exploited [151]. exposed [499, 921, 724, 889, 884, 657, 440, 289, 666, 477]. exposing [714]. exposure [65, 890, 472, 466, 817]. external [895]. extract [210, 154, 75]. eye [216, 53]. eye-lens [53]. eyefluke [715, 639].

F. [104]. Fabricius [427, 367]. factor [497, 751, 521]. factors [495, 58, 643, 895, 2]. failure [689, 99]. falcatus [31]. fallax [668]. false [643]. fam [315]. FAO [25]. farionis [774]. farm [395]. farmed [583]. fasciatusi [800]. fasdatus [817]. faster [182]. fat [392]. fate [424]. fathead [790, 690, 738]. fatty [15]. fauna [880, 122, 934, 138, 132, 761, 566, 577, 181, 223, 10, 160]. favourable [126]. features [589]. Fecundity [238, 844, 870, 221]. fed [105, 691]. feeder [86]. Feeding [693, 860, 71, 677, 471, 868, 609, 877, 520, 852, 587, 63, 183, 628, 697, 722, 781, 190, 531, 425, 602, 286, 577, 876, 793, 672, 869, 690, 334, 449, 821, 787]. female [351, 933, 167, 680, 863]. Fernworthy [445]. Ferox [797, 885]. fertile [290, 400]. fertility [870]. fertilization [176]. Fibre [673, 932, 620, 381]. fibrils [290]. field [486, 79, 341]. filamentosus [531]. file [765]. file-fish [765]. Filey [297]. Film [346]. filter [868, 781]. filtrates [300]. fimbriata [150]. fin [724, 549]. Fine [758]. fingerling [351]. fins [665]. First [185, 376, 315, 398]. Fischer [376]. Fish [366, 119, 261, 279, 68, 4, 486, 84, 13, 69, 255, 497, 784, 900, 155, 244, 748, 54, 390, 88, 916, 921, 470, 764, 733, 234, 592, 664, 391, 43, 618, 393, 676, 370, 520, 82, 667, 424, 624, 104, 343, 546, 557, 575, 588, 628, 363, 843, 101, 583, 219, 152, 380, 537, 561, 794, 934, 125, 126, 170, 179, 662, 194, 822, 559, 340, 313, 673, 455, 8, 431, 304, 389, 371, 849, 840, 566, 9, 64, 626, 395, 178, 438, 678, 21, 891, 747, 307, 643, 711, 131, 81, 508, 725, 721, 221, 658, 851]. fish [70, 99, 229, 671, 154, 753, 75, 657, 927, 928, 929, 518, 614, 2, 358, 97, 362, 153, 326, 845, 453, 765, 10, 398, 752, 291, 586, 103, 744, 785, 837, 145, 115]. fished [19]. Fisheries [115, 233, 248, 13, 50, 533, 47, 94]. fishery [233, 248, 346, 491, 206, 220, 137, 402]. Fishes [375, 655, 650, 693, 872, 609, 532, 718, 387, 405, 183, 462, 282, 562, 266, 93, 51, 808, 602, 506, 674, 420, 323, 716, 501, 579, 165, 479, 552, 831, 257, 53, 623, 541, 853, 271, 329, 660, 554, 71, 42, 113, 164, 350, 787, 551]. fishing [370, 492, 655, 909, 143, 482, 102, 25]. fistulated [611]. five [152, 215, 347]. flatfishes [177]. flavescens [427, 208]. flavolineatum [193]. flesh [365]. flesus [484, 328, 834]. flock [399]. Floodplain [818]. flora [570]. Flounder [96, 897, 328]. flounders [484]. flow [1, 385, 836, 289]. fltuviatilis [735]. fluctuating [330]. fluctuations [138, 378]. fluid [154]. flush [890]. fluviatilis [707, 880, 122, 593, 850, 22, 926, 7, 612, 52, 582]. foetus [125, 126, 170]. folds [662]. folic [808]. following [564, 780, 43]. fontinalis [652, 110, 635, 922, 417]. Food [471, 609, 559, 897, 259, 787, 231, 84, 260, 280, 877, 299, 314, 351, 89, 207, 20, 63, 183, 629, 129, 850, 628, 78, 66, 51, 147, 728, 342, 133, 531, 893, 286, 720, 577, 240, 276, 435, 443, 626, 827, 484, 501, 467, 623, 605, 869, 334, 336, 474, 833, 309, 167, 586, 585]. force [425]. force-feeding [425]. form [927, 928, 929]. formation [136]. formosa [400]. formosus [422]. forms [127]. formulation [532]. forskahlii [286]. Forssk [299, 301]. Forssk. [294]. Forsskål [653]. fossilis [608, 11, 550, 284]. Fouquet [246, 200]. four [693, 622, 376, 850, 215, 327]. fractionated [210, 154, 75]. fragility [929]. free [536]. Freetown [655]. Freeze [348]. frequency [921, 697, 674]. Frescon [79]. fresh [923, 929]. freshwater [835, 872, 746, 155, 379, 634, 640, 152, 380, 537, 313, 465, 509, 771, 438, 552, 510, 623, 853, 927, 928, 614, 185, 71, 164, 744, 13]. Fries [683]. friesii [683]. Frome [324]. fry [87, 360, 390, 302, 514, 923, 638, 235]. function [873, 763, 490, 709]. Functional [489, 620, 508]. Fundulus [430, 830, 312, 539, 104, 538, 410, 581]. Fundulusheteroclitus [290]. fungal [612]. Further [98].

Gadidae [28, 181]. gadoid [304, 389]. Gadus [100, 275, 541]. gairdneri [677, 890, 902, 392, 908, 783, 260, 565, 574, 468, 528, 206, 386, 697, 792, 826, 194, 633, 740, 905, 811, 610, 828, 654, 509, 159, 353, 558, 701, 489, 541, 414, 637, 839, 184, 741, 754, 466, 657, 529, 474, 289, 714, 778, 763, 745, 874, 411, 709, 586, 585]. Galaxias [434]. galaxiid [434]. Galaxiidae [72]. Galaxioidei [272]. Galeus [338, 651]. Gambmia [224]. Gambusia [608]. gariepinus [252, 355]. Garry [55]. gas [823, 836, 657, 490]. gas-bubble [490]. gasterostei [639]. Gasterosteus [499, 736, 109, 228, 555, 413, 501, 563, 598, 597, 858, 167, 238, 456, 680]. Gastric [856, 820, 697, 547]. gastrointestinal [439]. gayi [931]. gels [255]. gene [215]. general [453]. generator [143, 482]. generic [97]. Genetic [760, 282, 730]. genetics [148, 149]. genus [718, 664, 667, 272, 716, 53, 551]. George [399]. Georgia [376]. Germany [406]. Gervais [907, 250, 81]. giant [37]. Gibbons [352]. gibbosa [332]. Gila [273]. Gilbert [356]. gill [902, 321, 922, 793, 550, 466, 846]. Gillichthys [77]. gills [507, 100, 772, 634, 667, 792, 826, 899, 896, 228]. Gilquinia [367]. gilthead [930]. Girard [608, 737, 224, 273, 875, 410]. gland [846, 422]. globins [458]. glucose [131, 76]. Glugea [395]. Glukhova [3]. glycogen [913, 106]. glycolysis [633]. Gmelin [869]. goat [479]. gobies [198]. gobiid [664]. Gobiidae [198]. Gobio [141, 646, 383, 71]. Gobioidea [186]. Gobioidei [407]. Gobius [427, 37, 10]. goby [427, 122, 886, 878, 37, 683, 819, 186, 77]. goldfish [517, 472, 385, 195, 895, 373]. gonad [365, 325, 665, 332]. Gonadal [154, 844, 22, 553]. gonadotrophin [308]. gonadotrophs [527]. gonadotropin [895, 666, 930, 553, 527]. gonadotropin-induced [895]. gonads [409, 478, 754, 619]. gonionotus [154]. gonosomatic [751]. goosefish [174, 108]. gorbuscha [772]. Gorge [818]. gossypol [39]. Gracilisentis [861]. Graham [81]. granular [775, 650]. grass [30, 227, 893, 603, 691, 516, 235, 594, 604]. grasses [34]. gravel [166, 515]. Gray [605, 185]. grayling [66, 901, 5, 408]. graylings [272]. Great [364, 375]. Green [676, 440]. greenback [877]. Greenland [36]. grey [877, 607, 865, 359, 58, 209, 782, 743, 211, 210, 212]. Grillotia [461, 631]. grounds [636]. group [681, 43, 227, 297, 196, 145, 115]. groupers [660]. Growth [668, 641, 365, 915, 32, 136, 691, 594, 845, 392, 412, 295, 742, 681, 322, 280, 50, 723, 877, 376, 314, 587, 818, 182, 886, 687, 767, 850, 911, 575, 454, 147, 423, 722, 342, 514, 464, 194, 893, 209, 887, 240, 276, 435, 443, 420, 876, 259, 711, 467, 855, 912, 596, 52, 656, 773, 441, 401, 554, 449, 778, 804, 833, 757, 408, 586, 585, 769, 355, 140, 372]. gudgeon [141]. Guernsey [42]. Guichenot [931]. guirdneri [39]. Gulf [338, 339]. Günther [872, 72, 629, 464, 531, 286, 401]. guppy [870, 192, 242, 241, 341, 476]. Gut [782, 497, 207]. gwyniad [78]. gynogenesis [734]. gynogenetic [516]. Gyrodactylus [3]. Gyrometra [578].

H [619]. H-thymidine [619]. habit [431]. habitat [723, 685]. habitats [780, 877, 153]. habits [280, 877, 63, 531, 286, 897, 869, 653, 785, 252]. haddock [722, 461, 631, 263]. Haeckel [125, 126, 170]. haematocrit [243, 327, 289]. Haematological [573, 927, 928, 929, 457, 155, 244, 590, 574, 694, 811, 506]. haematology [565, 119, 905, 802, 744]. haemodynamics [823, 836]. haemoglobin [296, 110, 21, 929, 518]. haemoglobins [503, 298, 508, 458, 217]. haemopoietic [632]. Haemulon [193]. Hagenmüller [395]. hake [931]. halibut [36]. Ham [852, 573]. Hamilton [559, 891, 114, 91, 97]. Hamilton-Buchanan [559]. Hampshire [383]. handling [567]. Haplochromis [299, 399]. Harbour [692]. hard [397]. hardness [778]. Hardwicke [605]. harengus [452, 151, 314, 649]. Harpagifer [692]. hatchery [569, 704, 659, 813, 773, 824]. hatchery-confined [569]. hatchery-reared [704, 813, 824]. hatching [812, 514]. Hawaiian [330, 358, 396]. Haweswater [18]. Hay [454]. HCG [308]. head [29]. healing [349, 597, 70]. health [155]. hearing [786]. Heart [541]. heat [535]. heated [220]. Heavy [830, 669, 741, 581]. Heckel [835, 746]. hellerii [125, 126, 170]. Helminth [579, 128, 504, 383, 111, 429]. helminths [208, 931, 194]. hematopoietic [124]. Hemihaplochromis [387]. Henle [436]. Henneguya [667]. hepatic [753, 884]. hepato [450]. hepato-renal [450]. hepatocytes [440]. hepatosomatic [751]. herbicide [853]. herbivorous [483]. Herefordshire [66, 147]. hermaphrodite [403]. herring [452, 151, 314, 649]. heteroclitus [830, 312, 539, 104, 538, 581]. heterophils [775]. Heterophyidae [31, 187]. Heteropneustes [608, 11, 284]. Heteropneutes [550]. high [241, 143, 482]. Hilgendorf [299]. hinds [660]. hippoglossoides [36]. Hippopotamyms [622]. hippurus [497]. histamine [775]. Histochemical [197, 770, 620, 347]. histochemistry [846, 31, 567]. histological [517, 574, 645]. Histology [807, 88, 777]. Histopathological [667, 31, 328, 754]. Histopathology [861, 556, 444, 213, 229, 230, 426]. Histophathology [316]. historical [13]. history [394, 356, 41, 307, 334]. Holarctic [670]. Holmberg [162]. holubi [226, 331, 295, 176]. homing [809]. Hong [566, 577]. hooghliensis [91]. hook [19]. hook-avoidance [19]. hooks [326]. hormone [915, 211, 884]. hormones [866, 645]. horse [292, 525]. Host [880, 60, 748, 442, 79, 31, 371, 223, 524, 45]. host-parasite [45]. Hosts [257, 740]. House [393]. Hubbs [201]. Hudson [430]. human [308, 930]. hybrid [130]. Hybrids [762, 663, 511, 589, 351, 120, 695]. hybriin [502]. hydrated [534]. hydration [866, 154]. Hydrocephalus [87]. Hydrocynus [286]. Hydrogen [439]. hydrostatic [455]. hydroxydopamine [755]. hyperactivity [648]. hyperbaric [416]. hypercapnia [590]. Hypophthalmichthys [656]. hypophyseal [284]. hypophysectomized [11]. hypophysectomy [510]. hypophysis [645]. hypothalamic [527]. hypothalamo [284]. hypothalamo-hypophyseal-ovarian [284]. hypothalamus [645]. hypoxia [382, 289, 411]. hypoxic [205, 792].

I. [688]. ichthyo [566, 577]. ichthyo-fauna [566, 577]. Ichthyophthiriasis [218, 270, 287, 288]. Ichthyophthirim [246]. Ichthyophthirius [200]. ICR [60]. Ictalurus [437, 798, 316, 814, 617, 688, 513]. idella [30, 351, 227, 893, 603, 516, 235, 594, 604]. Identification [695, 500, 511, 130, 149]. identified [620]. ifluencing [58]. ignobilis [301]. IHN [124]. II [65, 218, 126, 140, 278, 382, 353, 488, 46, 229, 169, 836, 335, 149, 310]. III [270, 170, 372, 187, 70, 230, 336, 311]. images [477]. Immature [263]. Immersion [9]. Immune [40, 125, 126, 170, 391, 189, 526, 798, 883]. immunity [288]. immunogen [883]. immunoglobulin [822]. immunohistochemical [496]. importance [723]. important [845]. impoundment [676]. impoundments [375]. improvement [50]. inability [219]. inadequate [243]. Inbreeding [891]. incidence [151, 378]. including [570]. incorporation [848, 619]. increase [455, 476]. Increased [477]. increases [921]. increasing [854]. increments [365]. index [902, 751, 21]. India [552]. Indian [520, 920, 578]. Indiana [129]. indicated [282, 53]. indicator [537, 535, 548, 658, 518]. indicus [301]. Individual [341]. indomethacin [801]. Induced [44, 210, 75, 388, 106, 648, 444, 647, 895]. inducing [571]. Induction [890]. indus [360, 816]. industrial [291]. industrially [747]. infected [4, 667, 32, 40, 161, 828, 316]. Infection [395, 837, 556, 719, 200, 612, 414, 647]. infections [257, 230]. infectious [632, 60, 675, 271, 124]. infestation [269]. inflammatory [92, 161, 187]. Influence [889, 423, 474, 65, 260, 495, 307, 689, 773, 411]. influencing [643]. infra [784]. infra-red [784]. infrastructure [343]. initiating [521]. injected [424, 189, 154]. injection [564, 526]. injections [606]. Inland [233, 248]. innervation [755]. inorganic [462, 635]. inrelation [481]. inshore [557, 575, 588, 628, 181, 239]. insulin [696]. intake [314, 720]. integration [221]. intensities [781]. intensity [310]. Intensive [603, 909, 691, 617]. Interaction [735, 385]. intermediate [524]. International [204]. interrelation [69]. interrelationship [183]. interrenal [517]. intersex [857]. interspecies [927]. Interspecific [916, 609]. intertidal [682]. Intestinal [194, 807, 208]. intestinalis [4, 515, 524]. intestine [814]. intolerance [131]. intraovarian [352]. intraperitoneally [424]. intraspecies [928]. Intraspecific [373, 610]. Introduced [164]. introduction [375, 153]. introgression [215]. invertebrate [20]. invertebrates [279, 178]. Investigation [734, 681, 843, 679, 80]. Investigations [571, 657, 13, 260]. ion [308, 307, 439]. ionconcentrations [65]. Iraq [835, 746]. Ireland [642, 98, 640, 282, 540]. Irish [670, 141, 313, 111, 236, 447, 765]. irradiation [73]. Island [761, 259, 369, 345]. islands [761, 661, 193, 254, 42]. Isle [181, 223]. Isles [94, 880, 115, 421, 47, 164]. isocitrate [625]. isoenzymes [651]. isolated [772, 848]. isolation [718]. issue [118, 135, 146, 158]. IV [287, 444, 99].

Jägerskiöld [459]. Jan [661]. jar [671]. jarbua [653]. John [394]. Johnels [446]. Jordan [664]. journals [796]. juvenile [226, 331, 295, 852, 641, 365, 915, 759, 613, 710, 348, 789, 263, 919, 695, 594, 193, 330, 595, 773, 804, 833, 490, 500]. juveniles [821, 769].

Kafue [818]. Kainji [471, 286, 622, 609, 742]. Kamloops [159]. kannume [294]. Kariba [412]. karyotype [430]. Karyotypic [664, 104]. Katsuwonus [932]. Kenya [900, 543]. keta [772]. Ketone [658]. kidney [777]. killifish [430, 410, 581]. kinetics [930]. King [756]. Kinross [267, 268]. Kirk [599]. Kirtland [854]. kisutch [772, 894, 915, 756, 824]. known [338, 113, 350]. Kokanee [121]. Kong [566, 577]. krausii [542]. Krøyer [886, 878]. kunduchi [578].

L [919, 108, 895, 214, 481, 857, 547, 896, 862]. L. [507, 707, 935, 864, 450, 880, 452, 4, 486, 844, 866, 18, 497, 427, 913, 283, 19, 128, 463, 571, 76, 87, 360, 590, 499, 511, 932, 681, 88, 318, 871, 50, 599, 797, 205, 421, 225, 580, 593, 315, 151, 503, 397, 314, 865, 82, 682, 89, 207, 182, 312, 20, 475, 424, 601, 726, 539, 104, 820, 914, 646, 904, 302, 55, 495, 85, 636, 687, 850, 911, 575, 628, 473, 22, 197, 189, 269, 504, 275, 66, 147, 600, 719, 270, 287, 288, 910, 926, 722, 32, 859, 342, 110, 749, 750, 751]. L. [7, 325, 277, 278, 372, 425, 696, 816, 606, 923, 174, 856, 188, 130, 199, 736, 838, 141, 445, 406, 209, 31, 109, 899, 297, 237, 461, 263, 292, 525, 3, 240, 276, 435, 443, 28, 136, 228, 555, 171, 187, 395, 413, 137, 200, 259, 369, 345, 484, 501, 638, 522, 26, 67, 496, 734, 903, 883, 510, 699, 467, 909, 855, 540, 563, 598, 672, 743, 298, 458, 378, 265, 612, 572, 567, 689, 548, 385, 487, 560, 196, 195, 14, 107, 169, 426, 823, 836, 383, 328, 912, 258, 403, 731, 848, 52, 44, 211]. L. [210, 212, 111, 181, 223, 429, 457, 645, 505, 695, 374, 933, 515, 330, 396, 80, 142, 267, 268, 595, 582, 702, 441, 706, 858, 833, 821, 73, 236, 885, 71, 160, 148, 149, 888, 59, 477, 408, 167, 238, 456, 680, 863, 930, 62, 239, 649, 906]. Labeo [742, 852, 163, 296, 293, 537, 802, 217]. Laboratory [314, 290, 342, 209, 725, 410, 809, 13]. laboratory-reared [342, 410]. Labrador [733, 202]. labrax [896, 458]. Labrus [587]. Lacéepède [832]. Lacépède [825, 668, 531, 584, 251]. Lack [126]. lacking [290]. lactic [913]. laevis [303, 317, 523, 740]. Lagoon [150, 183, 464, 531, 865, 845, 63]. lagoons [102, 752]. Lagos [183, 63, 150]. Lake [346, 129, 78, 374, 544, 920, 128, 379, 747, 909, 334, 335, 336, 863, 471, 412, 900, 122, 868, 622, 609, 742, 322, 379, 280, 35, 299, 43, 491, 618, 399, 206, 799, 356, 542, 286, 841, 855, 851, 294, 543, 544, 401, 554, 787, 59]. Lakes [739, 233, 248, 82, 687, 850, 179, 8, 257, 375]. lamellae [922]. Lampetra [22, 7]. Lampetru [22]. lamprey [22, 237, 842]. lampreys [7]. landings [900]. landlocked [237]. Lanka [865]. large [233, 248, 375]. largemouth [810, 686, 105, 596, 584, 738, 832]. larvae [4, 592, 806, 151, 342, 445, 209, 558, 257, 410, 842, 584, 396, 649]. Larval [931, 406, 127, 105, 237, 263, 169]. larvicidal [776]. larvicide [917, 776]. late [934]. lateral [780, 347, 490]. Lates [491]. latex [196]. latipinna [416]. Laurel [323]. lauryl [517]. lavaretus [18, 481, 14, 403]. LDH [651]. lead [817]. learning [231]. Lekki [464, 531]. Lemna [604]. length [497, 280, 376, 855, 543, 254, 595, 441]. lens [53]. Leone [655]. leonensis [446]. leopard [186]. leopard-spotted [186]. Lepeophtheirus [442, 556]. Lepomis [648, 889, 440]. Lernaea [161, 258]. Lernaeocera [791]. lesion [444, 300, 213, 229]. lesions [612, 46]. lesser [431, 933, 309, 310, 311]. Lesueur [920, 688]. Lesueurigobius [683]. Lethal [179, 507, 917, 27, 657, 193]. leuciscm [189, 136]. Leuciscus [189, 276, 136, 762, 663]. leucocyte [650]. Leucocytes [433, 624, 432]. leucosticta [543]. Leukocyte [218]. leukocytes [775]. level [722, 138, 596, 411, 586, 585]. levels [392, 696, 472, 184, 154, 753, 884, 929, 873, 666]. Leven [267, 268, 582]. Ley [916]. liberator [775]. Life [394, 376, 22, 356, 635, 237, 41, 307, 334]. light [784, 781, 770, 793, 310]. Ligula [269, 4, 515, 524]. like [419]. Limanda [749, 750, 751, 547]. Limandaferruginea [897]. lime [534]. limi [854]. limitations [804]. line [195, 490]. lingua [187, 837]. lipid [405, 545]. liquid [764]. literature [155, 366, 533, 153]. little [338, 113, 350]. little-known [113, 350]. littoral [138, 132]. liver [405, 437, 188, 414, 754]. Living [25, 819]. Liza [877, 607]. Llyn [880, 346, 78, 133, 132, 139, 140, 672, 408]. LM [825]. loach [397, 383]. loading [694]. localization [496]. location [181]. loch [275, 602, 481, 403, 267, 268, 582, 441]. lochs [50, 797, 485, 488, 52]. loci [215]. Locomotor [648, 385, 103]. locomotory [673]. Lomond [481, 403]. Long [523, 649]. Long-term [523, 649]. longfin [789]. longirostris [454]. Lophius [174, 108]. loss [811, 555]. Lota [28]. Lough [98, 679, 540]. Loughs [885]. low [864, 65, 109, 417, 330, 930]. Lowe [186]. lower [493, 785]. lowered [235]. lowland [680]. Lucia [868, 787]. lucidum [216]. Luciolates [491]. lucius [19, 87, 88, 425, 696, 435, 26, 909, 457]. Lugg [66, 147]. Lumley [655]. luteus [835, 746]. Lymnaea [79]. Lymphocystis [452, 328]. lysozyme [496].

m [111, 21]. M. [387]. mackerel [292, 525]. Macleods [860]. macrochims [889]. macrochir [818, 841]. macrochirus [648]. macrophyte [907]. maculatus [918, 729, 326]. magenta [264]. mainly [13]. maintained [70, 741]. maintenance [192]. majalis [104]. major [852, 154]. Malaŵi [544, 851, 401, 554]. Malapterurus [869]. Malaya [422]. male [571, 351, 645]. malinoptera [607]. Mallotus [882]. mammalian [405]. Man [181, 223]. management [638]. manganese [789]. mangrove [693]. Manitoba [920]. Manter [357]. Manx [429]. marginatum [686]. Marine [25, 459, 532, 764, 15, 847, 363, 771, 620, 552, 848, 757, 71, 653]. Marini [201]. marinus [842, 237, 309, 310, 311]. mark [862]. mark-recapture [862]. marked [636, 766]. marking [85, 824]. marmorata [376]. Mask [885]. mass [732]. Mastacembelus [251]. material [424, 407]. materials [189]. maternal [514]. maturation [844, 495, 716]. mature [538, 905]. maturity [571]. Maul [127]. maxima [358]. maximus [450, 820, 613, 199, 62]. may [495]. Mayen [661]. Mcllwaine [841]. Meads [752]. meal [820, 547, 856]. means [388]. measure [541]. measurement [103]. measuring [874]. Mechanical [832]. mechanisms [424, 657]. mediated [718]. Mediterranean [497, 53, 551]. Medway [785, 791, 906]. Meeting [48]. Melanogrammus [722, 461, 631, 263]. melanophore [760]. melanophores [755]. melas [688]. melastigma [800]. melastomus [651]. membras [452]. menhaden [423]. Mercury [234]. meridionalis [401]. Meristic [8, 446]. merlangi [100]. Merlangius [575, 628, 501, 429, 791, 834]. merlangus [100, 575, 628, 501, 111, 181, 223, 429, 791, 834]. Merluccius [931, 339]. Metabolic [816, 423, 541]. metabolism [630, 732, 713, 879, 405, 545, 106, 425, 382, 626, 90, 273]. metacercaria [357]. metacercariae [187, 371]. metacercarial [31]. metal [669]. metals [830, 741, 581]. metamorphosis [237]. Metazoan [733, 202, 920, 739]. Methallibure [6, 665]. methanesulphonate [561]. method [283, 492, 302, 51, 643, 862, 221, 212, 102]. methodology [233, 248]. methods [244, 570]. methylene [535]. methyltestosterone [365]. Mex [601]. Mexican [845]. microhaematocrit [561, 110]. Micromesistius [806, 588]. Micrometrus [352]. Micropharynx [459]. microps [886, 878]. Micropterm [832]. Micropterus [810, 686, 694, 105, 596, 584, 738]. micropyle [538]. microscope [899]. microscopic [770, 814]. microscopy [352, 281, 597, 316, 261]. microsomal [753, 884]. Microspectrophotometric [255]. microspheres [836]. Microsporidia [395]. migrating [540, 368]. migration [601, 704, 703, 705, 702]. migratory [1, 162, 324, 661]. milkfish [730]. minimus [352]. minnow [128, 690]. minnows [790, 809, 738, 863]. minutus [819]. mirabilis [77]. mirror [218, 270, 287, 288, 633, 200, 477]. Mitchill [110, 922, 729, 858, 920]. mitochondria [634]. mitochondria-rich [634]. mixed [789]. mixtures [27, 175]. model [738]. moderate [921]. modes [838]. Modification [241]. molitrix [656]. molluscicide [79, 626]. molluscs [374]. monitoring [233, 248, 212]. Monogenea [100, 487]. monstrosa [673]. months [884]. Moor [393]. Morass [860]. mordax [867]. morhua [275, 541]. mormyrid [622, 609]. Mormyriformes [294]. Mormyrus [294]. Morone [56, 904, 161, 493]. Morphological [804, 775, 589]. Morphology [846, 516, 490, 177, 815, 489, 422]. Morphometrics [332, 520, 823]. mortality [465, 369, 638, 554]. mosquitofish [224]. mossambica [65, 322, 379, 245, 354, 33, 253, 281, 285, 665]. mossambicus [732, 917, 717, 665, 776, 626, 508]. mother [125, 126, 170]. mountain [579, 875]. mouse [60]. movement [1, 860, 813]. Movements [196, 505, 236, 470, 668, 55, 297, 369, 485, 268, 706, 278]. MS [359, 561, 927, 928, 929, 529]. MS-222 [561]. MS2 [883]. mtilus [911]. mucous [197, 265, 567]. mucus [658, 518]. mud [573]. mudfish [163, 418]. mudminnow [854]. Mugil [205, 865, 359, 58, 209, 31, 549, 743, 487, 560, 44, 211, 210, 212, 330, 396]. Mugilidae [782]. Muller [413, 442, 556, 436, 79, 303, 317, 422, 62]. mullet [205, 877, 607, 865, 359, 724, 58, 209, 31, 549, 782, 743, 487, 560, 44, 211, 210, 212, 330, 396, 500]. Mullidae [479]. multi [520]. multi-species [520]. multifiliis [246, 200]. Multiple [458]. mummichog [830]. muscle [913, 932, 363, 770, 621, 620, 536, 651, 763, 382]. muscles [780, 188, 633, 673, 347]. musculature [904, 381, 382]. mustela [669]. Mycobacteriosis [569]. Mycological [16]. myotomal [188, 763]. myotomes [620]. Myxosoma [498, 534, 535]. myxosporidian [667].

n [321, 574, 578, 626]. N-tritylmorpholine [626]. NADP [625]. NADP-isocitrate [625]. naevus [436]. Naivasha [543]. National [393, 579]. natural [19, 511, 379, 871, 41, 120, 691, 927, 928, 929, 695, 554, 153, 301]. Nature [393, 677, 473]. Neagh [540]. near [934, 23, 64]. near-surface [934]. nebulosa [605]. nebulosus [688]. nec [91]. necrosis [632, 60, 675, 67, 300, 29, 46, 70, 99, 426, 124]. Nematocentris [872]. Nematoda [151, 445, 257, 649]. nematode [41]. Neochanna [72]. Neogobius [122]. neotype [114]. nerka [632, 121, 568]. nerve [207]. nerve-section [207]. Nesting [692]. net [390]. Netherlands [636, 766, 767, 911, 153]. neutralized [927, 928, 929, 529]. nevadensis [483]. newborn [876]. Newfoundland [384]. newly [213]. Nibea [91]. nickel [27]. Nicoll [357]. Nigeria [471, 609, 742, 150, 464, 531, 286]. night [593]. nigrodigitatus [531]. nigrofasciatum [725]. Nile [491]. nilotica [844, 866]. nine [900, 214, 109]. nine-spined [214, 109]. Nishio [31]. nitrogen [764, 596]. nitrophenol [564]. Nitzsch [600, 719]. nobilis [656]. Noemacheilus [397, 383]. non [434, 822, 901]. non-breeding [901]. non-diadromous [434]. non-stimulated [822]. normal [915, 179, 811, 29, 107, 192]. normoxic [792]. North [78, 133, 132, 139, 140, 408, 693, 315, 631, 642, 459, 454, 934, 749, 750, 751, 736, 661, 631, 292, 485, 488, 398, 688, 160, 62, 649]. north-east [631]. north-eastern [315]. North-west [642]. North-Western [398]. Northern [425, 696, 816, 540, 50, 266, 646, 934, 111, 10, 642, 98, 338]. norvegicus [407]. Norway [557, 661]. Norwegian [101]. Nosema [395]. note [98, 402, 766, 850, 819, 130, 768, 689, 235, 656, 193]. Notemigonus [545, 553, 527]. Notes [18, 111, 42, 622, 399, 588, 313, 3, 502, 36]. Notice [116]. Notopterus [90]. Notothenia [376]. Notropis [35, 454]. Nottinghamshire [130]. nov [398]. nuclear [220]. nucleus [645]. number [267]. numbers [93, 110]. numerical [221]. número [118, 135, 146, 158]. nutrient [194]. nutritional [839]. nutritive [34]. Nyasa [544].

O [756, 227]. O-group [227]. O. [772]. Obituary [116, 320]. Observations [4, 100, 775, 589, 37, 251, 152, 686, 303, 722, 161, 800, 240, 276, 435, 443, 841, 876, 568, 831, 446, 869, 252, 301, 486, 825, 376, 359, 820, 646, 687, 337, 78, 934, 110, 828, 338, 339, 705, 513]. observed [368]. Occult [518]. Occurrence [650, 857, 329, 100, 908, 427, 871, 61, 640, 588, 101, 304, 800, 549, 258, 649]. Ocean [459, 520, 578]. ocellata [876]. Odontogadus [111, 181, 223]. odour [385]. oesophagus [814]. oestradiol [545]. oestradiol-17 [545]. oestrogen [184]. off [181]. offshore [181]. offspring [400]. oil [724, 549]. olfaction [309]. olfactory [304, 389]. Oncorhynchm [824]. Oncorhynchus [632, 569, 772, 894, 915, 121, 568, 756]. Onji [31]. onlthe [755]. Ontario [137, 323]. Ontogenetic [298, 508]. oocytes [848]. oogenesis [858]. open [341]. operated [86]. operating [546]. operations [492]. opercular [742]. Ophicephalus [536]. Ophiocephalus [449]. oral [131]. Oregon [737]. organ [468]. organism [419, 498]. organization [389]. organophosphate [917, 776]. organs [815]. orientation [385]. osculus [273]. Osmerus [867]. osmolarity [65]. Osmoregulation [331]. Osmotic [465]. Osteoglossidae [422]. Osteoglossomorpha [294, 422]. other [622, 931, 82, 670, 343, 33]. otolith [757]. otoliths [180]. outbreaks [271]. oval [823, 836]. ovarian [835, 866, 463, 436, 716, 212, 284]. ovary [126, 749]. Overwinter [190]. ovulation [866, 308, 801, 895]. ovum [538]. oxidase [562]. oxide [574]. Oxygen [499, 528, 584, 935, 864, 682, 109, 416, 327, 396, 477].

Pacific [264]. paddlefish [777]. paganellus [427]. pages [627]. paired [22]. Pallas [37, 882, 901, 90]. pancreatectomy [816]. pancreatic [60, 675]. papers [48]. papillomas [378]. parameters [894, 810, 694, 506, 159, 353, 573]. parasite [880, 748, 761, 3, 181, 223, 774, 45]. Parasites [214, 847, 96, 323, 493, 128, 667, 202, 313, 678, 579, 552, 487, 560, 383, 111, 429, 920]. parasitic [160]. parasitica [459]. parasitical [258]. parasitism [413, 563]. parasitocoenoses [739]. Parasitology [737]. parathion [385]. Parental [773]. Paris [204]. parks [579]. Parophrys [737]. parr [444, 213]. Partial [163, 816]. particle [833]. particular [53]. Particulate [781, 424]. partitioning [904]. parvipinnis [410]. parvum [882]. PAS-GL [650]. PAS-positive [775]. passage [89, 207]. Pasteurella [419]. Pasteurella-like [419]. pathogenesis [124]. pathogenicity [303, 445]. Pathology [145, 115, 730, 674, 257, 46, 524, 270]. pattern [766]. patterns [852, 253]. pCO [894]. Pear [910, 926]. Pear-shaped [910, 926]. pearl [700]. pectoralis [442, 556]. peculiar [343]. pejerrey [280]. pelagic [592, 491, 82, 263]. pelamis [932]. Pellonula [446]. Peninsula [266, 692]. pennantii [78]. Peñuelas [280]. Peracute [613]. Perca [707, 880, 735, 208, 593, 850, 926, 612, 52, 582]. Percafluviatilis [910]. perception [385]. perch [707, 880, 872, 735, 208, 491, 593, 850, 910, 926, 639, 612, 52, 582, 16, 56]. Percidae [215]. Perciformes [352]. peregra [79]. performance [188]. perfused [772]. periodic [650]. Periodicity [702]. peripheral [798, 849]. permeability [76]. pesticide [634]. petenense [781, 827]. Petenia [542]. Peters [917, 65, 872, 322, 379, 245, 354, 732, 870, 33, 253, 281, 285, 626, 508, 192, 242, 241, 341, 476]. Petromyzon [237, 842]. pH [782, 417, 778]. phagocytic [424]. phenol [27, 714]. pheromone [718]. philander [387]. Philypnodon [334, 335, 336]. pholis [682, 731, 848, 477]. phosphatase [886]. Photokinetic [842]. photoperiod [666, 619]. Phoxinus [128, 809, 863]. physical [652]. physical-chemical [652]. Physiological [858, 808, 509, 287]. physiology [152]. physoclist [823, 836]. Pices [215]. picture [645]. pictus [622, 578]. pike [19, 87, 360, 88, 425, 696, 816, 435, 26, 909]. Pilcomayo [162]. Pimephales [790, 690, 738]. pinealectomy [546, 666, 527]. pink [772]. piperatus [338]. piscatorius [174, 108]. Pisces [127, 162, 322, 379, 299, 151, 399, 312, 72, 542, 384, 400, 114, 551, 36]. Piscesa [573]. piscivorous [787]. pit [515]. pituitary [405, 163, 422, 210, 154, 75, 666, 553, 527]. plaice [913, 225, 580, 89, 207, 424, 433, 432, 504, 512, 728, 856, 297, 3, 187, 395, 496, 699, 196, 107, 169, 328, 73, 160]. planarian [459]. planeri [22, 7]. plankton [234]. plant [220]. planting [166]. plantings [137]. plants [351, 34]. Plasma [174, 108, 184, 255, 481, 386, 293, 696, 892, 839, 753, 666]. platensis [162]. platessa [913, 225, 580, 89, 207, 424, 433, 432, 504, 728, 856, 297, 3, 187, 395, 496, 699, 196, 107, 169, 328, 73]. Platichthys [484, 328, 834]. Plecoglossus [308]. Plectorhinchus [578]. Plectropomus [326]. plerocercoid [4, 461, 367, 413]. plerocercus [631]. Pleuronectes [913, 225, 580, 89, 207, 424, 433, 432, 504, 728, 856, 297, 3, 187, 395, 496, 699, 196, 107, 169, 328, 73]. Pleuronectesplatessa [160]. Pleuronectiformes [177, 36]. Plover [566, 577]. Poecilia [870, 6, 716, 760, 416, 400, 192, 242, 241, 341]. poeciliid [125, 126, 170]. Poeclia [476]. poikilothermic [171]. poikilotherms [768]. Polish [82]. Pollachim [823]. Pollachius [836]. pollan [670, 670]. polluted [813]. Pollution [25, 204]. polyacrylamide [255]. polychlorinated [753]. polymorphism [625]. Polyodon [777]. polyploidy [388]. pomacentrid [266]. Pomacentridae [551]. Pomatomidae [315]. Pomatomus [315, 904]. Pomatoschistm [407]. Pomatoschistus [886, 878, 819]. Pomphorhynchus [303, 317, 523, 740]. pond [283, 852, 559, 549, 617]. pond-ecosystem [549]. ponds [81]. Ponto [122]. Ponto-Caspian [122]. pool [420]. populatioins [865]. Population [128, 43, 487, 560, 430, 295, 715, 871, 208, 870, 129, 683, 767, 911, 317, 277, 278, 372, 867, 523, 639, 803, 887, 566, 339, 186, 522, 483, 467, 909, 485, 221, 690, 515, 192, 149, 59, 680, 476, 791, 139, 335]. Populations [5, 283, 394, 98, 393, 557, 575, 588, 628, 23, 8, 881, 502, 752, 291, 785, 837]. porcine [892]. positive [775, 650]. Possible [718, 535, 215]. post [127]. post-larval [127]. potency [527]. potential [618]. Pout [557]. poutassou [806, 588]. powan [14]. power [220, 2, 143, 482]. practised [102]. Precipitating [504]. precision [874]. predation [914, 374]. Predator [738]. Predator-prey [738]. predators [827]. predatory [84, 851]. preferences [231]. Preliminary [463, 260, 646, 934, 605, 486, 473, 340, 209, 171, 658, 254]. premature [756]. preopticus [645]. prepared [691]. presence [861]. presented [48]. pressure [256, 455, 416]. pretreated [829]. prey [738, 804]. Priacanthus [216]. primary [526]. principal [583]. process [70]. Prochilodus [162]. produce [219]. produced [243]. producing [890]. Production [84, 17, 334, 335, 336, 69, 290, 641, 759, 767, 911, 923, 285, 64, 779, 400, 845, 167, 456, 336]. Productivity [483]. Progress [368]. Progressive [905]. prolactin [405, 510, 553]. promelas [790, 690, 738]. properties [652]. Prosopium [875]. Proteases [898]. Protein [642, 731, 606, 67]. proteins [255, 652, 282, 163, 293, 174]. proteinst [53]. Prototroctes [272]. Prototroctidae [272]. Protozoa [395]. proximate [237]. Pseudaphya [198]. Pseudobranch [771, 896]. Pseudochromi [264]. Pseudochromidae [264]. Pseudophyllidea [4, 685, 406, 413, 62]. publicados [118, 135, 146, 158]. Publications [796, 615]. published [118, 135, 146, 158]. pulse [143, 482]. punctatus [437, 798, 849, 840, 924, 892, 316, 536, 814, 617, 513]. Pungitius [214, 109]. Puntius [154]. pupfish [483]. quadracus [858]. quality [820]. quantitative [231]. Quarun [799]. Queensland [693, 872]. Quietula [664]. Quinaldine [532, 564, 682].

races [421, 325, 514]. raceway [617]. radiata [459]. radiated [384]. radiation [854]. radio [703]. radioactive [512, 728]. Radiocadmium [312]. radiography [820]. radiolabelled [836]. Rafinesque [35, 918, 889, 437, 161, 316, 440, 617, 688]. Rainbow [908, 231, 677, 890, 902, 392, 783, 260, 27, 565, 468, 528, 206, 92, 386, 39, 792, 826, 190, 194, 381, 382, 740, 811, 610, 828, 654, 720, 509, 159, 353, 558, 701, 829, 489, 414, 637, 839, 184, 813, 741, 754, 466, 657, 529, 474, 289, 714, 778, 745, 874, 709, 586]. rainbowfish [872]. Raitt [309, 310, 311]. Raja [459, 436]. rank [889]. Rao [91]. Raphidascaris [41]. rapid [51, 455]. Rare [113, 350]. rate [89, 207, 687, 767, 850, 911, 697, 423, 514, 887, 541, 836, 690, 289, 773]. rates [793, 554]. ratio [280, 540]. ration [596, 474, 582, 167, 586, 585]. ray [436]. rays [512]. readily [85]. Reading [64]. reared [342, 923, 209, 704, 410, 813, 773, 824]. rearing [86, 595]. reba [337]. recapture [283, 862, 447]. received [796]. receiving [365]. recognizable [85]. record [315, 185, 398]. recording [470]. records [784, 371]. recovery [188]. Recruitment [881, 540]. Red [932, 266, 486, 784, 87, 363, 694, 770, 633]. red-disease [87]. red-sore [694]. redescription [407]. reef [165, 479, 831, 660]. reference [880, 609, 681, 50, 685, 761, 776, 153]. regimes [666]. regulation [676]. reinhardtii [860]. Reinhardtius [36]. related [433, 400]. relation [830, 886, 188, 253, 490, 201]. relations [125, 126, 170, 577]. Relationship [894, 694, 540, 497, 748, 280, 591, 299, 376, 22, 110, 325, 602, 638, 855, 543, 441, 97]. Relationships [720, 623, 282]. relative [514, 543, 254, 242]. released [766]. releasing [527]. renal [450, 564]. rendalli [630, 591, 723, 713, 574, 480]. rendatti [387]. Renicola [371]. Renicolidae [371]. repair [171]. report [364, 473, 209, 94, 176, 658]. reproduction [844, 866, 646, 147, 722, 464, 240, 276, 435, 443, 855]. Reproductive [872, 345, 543, 622, 718, 481, 251, 454, 749, 750, 751, 292, 294, 422, 933, 709, 239, 112]. requirements [584]. rerio [533, 891, 721]. Research [204, 346, 533]. Reserve [393]. Reservoir [733, 445, 566, 577, 80]. resident [761, 661]. Resistance [829, 455]. resorption [463, 823, 836]. Resources [25, 233, 248]. respect [308]. respiration [413]. Respiratory [205, 273, 470]. Response [11, 4, 517, 391, 92, 218, 40, 526, 808, 161, 187, 883, 154, 330, 804, 241, 341, 553]. responses [921, 205, 189, 811, 842]. Restoration [913]. results [463]. Resumenes [118, 135, 146, 158]. retention [596]. reticulata [870, 6, 760, 192, 242, 241, 341, 476]. retina [106]. return [636]. returning [229]. returns [139]. Review [564, 274, 519, 936, 681, 755, 13, 155, 366, 391, 624, 495, 570, 685, 533, 614, 148, 551, 112]. Reviews [12, 24, 38, 49, 57, 74, 83, 95, 117, 134, 144, 156, 172, 191, 203, 232, 247, 262, 305, 319, 344, 361, 377, 404, 415, 428, 451, 460, 469, 494, 530, 576, 616, 644, 684, 727, 795, 925].

Revised [129, 482]. Rhabdosargus [226, 331, 295]. Rhabdospora [825]. rhabdovirus [360]. Rhinichthys [273]. Rhodesia [841]. rhomboidalis [342]. rhombus [62]. rhythm [405, 708]. Rhythmic [704, 711]. rhythms [710, 701, 548, 334]. rich [634, 206, 656]. Rich. [558, 414]. Richardson [677, 890, 902, 392, 908, 260, 574, 386, 39, 826, 194, 905, 828, 654, 509, 159, 353, 701, 489, 637, 839, 184, 741, 168, 474, 289, 714, 586, 585]. riisei [846]. ringer [225]. Rive [430]. River [1, 599, 394, 137, 434, 7, 813, 291, 84, 162, 393, 66, 147, 324, 317, 277, 278, 372, 23, 523, 443, 64, 881, 41, 383, 408]. rivers [233, 390, 5, 240, 276, 435, 706, 453, 762, 663, 248]. roach [511, 780, 681, 715, 871, 636, 687, 911, 147, 859, 898, 857, 803, 887, 240, 903, 855, 298, 515, 502, 663, 59]. Roccus [56]. rock [427]. rodlet [825, 896]. rohita [852, 802]. Rohu [802]. Role [904, 833, 309]. roles [620]. rossii [376]. rostrata [770]. rot [549]. rotenone [178]. route [873]. Routine [244, 630, 732, 713, 423]. rubescens [486]. rudd [511, 715, 859, 898, 903, 298, 502]. Rudolphi [79]. Rüppell [301]. Rutilu [887]. Rutilus [911, 511, 780, 681, 871, 636, 687, 147, 859, 898, 857, 803, 240, 903, 855, 298, 502, 663, 59, 887]. Rye [752].

S. [792, 661, 740, 53, 62]. S.E [427]. saccus [33]. salar [935, 864, 642, 590, 421, 625, 182, 302, 55, 495, 759, 419, 786, 32, 923, 661, 704, 703, 672, 368, 595, 702, 773, 706, 804, 833, 821, 236, 148, 149]. Salinity [608, 410, 396, 935, 226, 423, 654, 678, 307, 510, 839, 480]. Salmo [507, 677, 935, 864, 642, 890, 902, 392, 908, 783, 76, 590, 260, 735, 50, 599, 797, 565, 421, 574, 625, 468, 528, 182, 20, 475, 206, 640, 726, 914, 302, 55, 495, 386, 759, 697, 419, 197, 786, 39, 512, 728, 32, 826, 194, 526, 606, 822, 923, 633, 445, 739, 661, 740, 905, 811, 610, 828, 654, 509, 159, 353, 704, 703, 558, 137, 779, 881, 638, 522, 67, 919, 883, 510, 701, 793, 467, 672, 265, 572, 689, 489, 541, 548, 414, 426, 637, 839, 184, 741, 754, 466, 657, 168, 529, 505, 695]. Salmo [474, 368, 289, 714, 80, 142, 267, 268, 595, 702, 773, 441, 706, 778, 763, 804, 833, 821, 745, 874, 411, 236, 885, 709, 148, 149, 888, 586, 585]. salmoides [810, 686, 694, 105, 596, 584, 738, 832]. Salmon [786, 935, 864, 642, 632, 569, 772, 421, 625, 17, 182, 641, 365, 915, 302, 55, 495, 759, 419, 121, 32, 40, 923, 340, 348, 704, 703, 86, 444, 568, 120, 672, 29, 213, 229, 230, 210, 695, 705, 368, 873, 595, 702, 773, 706, 804, 833, 821, 236, 447, 148, 149, 888, 756, 824]. salmonicida [647]. salmonid [812, 98, 23, 679, 327]. Salmonidae [403]. salmonids [1, 546, 498, 324, 202, 300, 29, 46, 70, 99, 426]. Salmoniformes [72, 854, 403]. salmositica [828]. salt [805, 8, 789]. saltatrix [315, 904]. salts [175]. saltwater [465]. salvelini [661]. Salvelinm [797]. Salvelinus [652, 629, 600, 719, 110, 761, 661, 635, 259, 369, 345, 922, 265, 567, 417]. sample [390]. samples [327, 160]. sandeel [309, 310, 311]. Sandoz [359]. sapidissima [394, 710]. Saprolegnia [32, 40]. saprolegniasis [568]. Sardinella [332]. Sarotherodon [917, 732, 818, 717, 665, 841, 776, 626, 508]. Satiation [697]. saxatilis [904, 493]. Scagerac [462]. Scale [653, 886, 878]. Scale-eating [653]. scales [222, 281, 136, 643, 711]. Scanning [352, 281, 899, 261]. scarcer [164]. Scardinim [511]. Scardinius [859, 898, 903, 298, 502]. schelly [18]. Schiff [650]. Schistocephalus [413]. Schlegel [308, 422]. Schmaltz [918]. sciaenid [91, 97]. Sciaenidae [114]. Scleropages [422]. Scomberomorus [729]. Scophthalmus [450, 820, 613, 199, 62]. scorpii [62]. Scotland [61, 599, 17, 481, 55, 918, 557, 575, 588, 628, 403, 267, 268, 582, 239]. Scottish [50, 797, 182, 641, 683, 275, 602, 5, 186, 52, 706]. sculpin [356]. Scyliorhinus [933]. Sea [266, 485, 488, 222, 599, 640, 726, 495, 275, 342, 923, 602, 237, 842, 671, 185, 930, 769, 436, 462, 934, 749, 750, 751, 882, 631, 292, 674, 485, 488, 111, 160, 62, 649]. sea-age [495]. sea-bream [930]. seas [113, 350]. season [694, 859, 898, 903, 223, 201]. Seasonal [593, 852, 749, 750, 840, 699, 701, 378, 572, 284, 709, 609, 481, 744]. seasons [165, 831]. seawater [634, 30, 312, 756]. seawater-adapted [634]. second [524]. Secondary [230]. Secretary [94]. section [207]. sections [467]. segregation [340]. seines [655]. selected [155, 366, 894, 386, 874]. selection [473, 559, 484, 501, 77]. Selective [914, 374]. selectivity [351, 804]. self [137]. self-sustaining [137]. SEM [815, 550]. seminal [11, 154]. senegalensis [742]. sensitizing [219]. separated [255]. separating [446]. separation [163]. September [204]. septicaemic [419]. Sequence [566]. sequential [46]. sera [504, 99]. serological [328]. Serranidae [660, 458, 326]. Serum [606, 67, 873, 642, 65, 822, 307, 637, 154, 617, 745, 874]. settlement [297]. Seven [679]. Seventh [204]. several [507]. severe [382, 444]. Severn [277, 278, 372, 668, 484]. sewage [741, 752]. sex [280, 540, 223]. sexual [571, 473]. sexually [905]. shad [394, 668, 150, 781, 710, 465, 827]. shallow [794, 602]. shallow-water [794]. shanny [384]. shaped [910, 926]. shark [338]. sharks [651]. Shasta [353]. Shatt [746]. Shatt-Al-Arab [746]. shelters [476]. shiner [35]. shoals [809]. shock [388]. shore [42]. shown [196]. sialic [902]. ‘Siamese [717]. ‘Siamese-twins’ [717]. Sibaya [322, 379]. Sierra [655]. Sierrathrissa [446]. sign [351]. signal [843]. significance [650, 508, 149]. Siluriformes [299]. silver [540, 656]. Simcoe [35]. simple [166, 86, 123]. Simulium [776]. Sinai [266]. since [818]. single [296, 606, 789, 21]. Sinha [91]. site [126]. six [884]. Size [467, 69, 818, 820, 325, 547, 856, 523, 474, 804, 833, 167, 586, 201]. Skadar [374]. Skagerrak [506]. skate [459]. skeletal [659, 721, 536]. skin [886, 219, 187, 597, 658, 29, 107, 169, 518, 888]. skin-sensitizing [219]. skipjack [932]. Slapton [916]. sludge [741]. small [607, 227, 638, 334, 335, 336, 103]. smallmouth [176]. Smallwood [733]. smelt [867]. Smith [387, 163, 296, 293, 537, 217]. smolt [923, 705, 702, 123]. smoltification [756]. smolting [773]. smolts [935, 864, 704, 706]. snail [79]. snappers [254]. social [889]. Society [115, 47, 94]. sockeye [632, 121, 568, 873]. sodium [65, 805, 517, 720, 440]. soft [790]. solar [695, 888]. Sole [737]. Solea [799]. solidus [413]. soluble [606]. solution [225]. solutions [792, 789, 440, 714]. Some [486, 376, 879, 359, 343, 820, 337, 78, 58, 110, 542, 547, 141, 811, 109, 828, 357, 367, 525, 338, 339, 159, 353, 754, 884, 705, 2, 524, 714, 544, 453, 160, 513, 124, 36, 650, 589, 15, 82, 365, 726, 588, 179, 506, 509, 178, 501, 853, 429, 34, 554, 845, 242]. sonar [54]. sonic [103]. sore [694]. Sound [285, 259, 369, 345]. South [162, 322, 379, 355, 467, 376, 152, 339, 909, 813, 291, 500]. southern [646, 240, 276, 435, 272, 401, 160, 62, 835, 920]. sp [321, 151, 578, 406, 398]. space [610]. Spams [659]. spangled [872]. Spanish [118, 135, 146, 158]. Sparidae [226, 331, 295, 342]. sparrmanii [387]. Sparus [930]. spathaceum [715, 79]. spathula [777]. Spatial [821]. Spawning [599, 867, 930, 69, 394, 636, 473, 7, 340, 801, 165, 831, 44, 210, 75, 201, 301]. Special [157, 249, 880, 609, 681, 50, 308, 761, 153]. Specialization [272]. species [471, 622, 718, 748, 715, 916, 43, 491, 520, 82, 399, 63, 104, 22, 179, 662, 639, 685, 304, 389, 215, 178, 198, 21, 347, 927, 928, 929, 327, 254, 358, 153, 858, 845, 164]. species-flock [399]. specific [225, 832]. specificity [880, 442, 303, 685]. specimen [403]. specimens [97]. speckled [273]. speed [700]. speeds [763]. spermatogenesis [6]. spermatozoa [58]. spermatozoon [758]. spermiation [211]. Spicara [53]. spinax [651]. spined [214, 736, 109, 413, 563, 598, 167, 238, 456, 680]. Spitsbergen [661]. splendida [872]. splenectomized [811]. Spontaneous [721, 895]. spores [535]. sport [220]. spot [427, 14]. spotted [918, 186, 933]. spp [604]. sprat [501, 239]. sprattm [906]. Sprattus [501, 239, 906]. spring [220, 80, 48]. Springer [338]. squali [367]. Squalius [443]. Sri [865]. St [868, 787, 394]. Stäbchendrüscnzellen’ [343]. stability [874]. stage [307]. stages [805, 877, 635, 263]. stained [513]. staining [535, 9]. stappersii [491]. starvation [425, 654, 658]. starved [105]. statistics [402]. status [412, 491, 126, 740, 28, 472, 699, 701, 53, 839, 91]. steelhead [490]. Steindachner [226, 331, 295, 542, 176, 388]. Stellantchasmus [31]. stephani [395]. Stephanostomum [357]. sterile [671]. steroid [866, 753, 884]. Stichaeidae [384]. stickleback [214, 736, 109, 228, 555, 413, 501, 563, 598, 167, 238, 456]. sticklebacks [499, 847, 858, 680]. Still [627, 130]. stimulants [677]. stimulated [822]. Stizostedion [708]. stock [520]. stocking [342, 514, 893, 80]. stocks [151, 679]. Stokell [434, 334, 335, 336]. stomach [676, 814]. stomachs [329, 326]. Stomiatoidei [127]. stone [397, 383]. Storage [764]. Storer [384, 897]. Stour [443]. strains [417]. strategies [646, 782]. Strathbeg [441]. stream [17, 182, 641, 166, 483, 467, 505, 809]. streams [340, 638]. strength [687]. Streptococcus [908]. Stress [862, 902, 565, 537, 465, 90, 518, 271, 457]. striatus [449, 769]. striped [205, 904, 493, 487, 560, 44, 396]. Strontium [824, 222]. structural [594]. Structure [567, 295, 878, 758, 139, 803, 107, 823, 466]. structures [281]. Studies [677, 844, 866, 162, 607, 397, 250, 878, 381, 382, 661, 431, 566, 577, 187, 131, 598, 743, 853, 29, 46, 70, 107, 99, 169, 426, 515, 524, 752, 408, 769, 825, 463, 574, 314, 351, 227, 462, 53, 605, 884, 400, 148, 149, 309, 310, 311, 355]. Study [815, 21, 414, 231, 880, 430, 742, 517, 634, 424, 583, 277, 278, 372, 188, 633, 23, 340, 611, 506, 899, 171, 620, 793, 347, 550, 328, 327, 254, 358, 476, 744]. studying [703]. stunted [915, 803, 887]. stunting [871]. Sub [917, 713]. sub-adult [713]. Sub-lethal [917]. subbifurcata [384]. subjected [634]. Sublethal [362, 648, 889, 472, 817]. subsequent [370]. substances [462]. subviridis [877]. sucker [825]. Sueur [861]. sugar [537]. Sulphamerazine [168]. sulphate [564, 532, 517, 397]. Summaries [118, 135, 146, 158]. summer [582]. sunfish [129, 440]. superficial [662]. supplemental [137]. supplemented [365]. Surface [662, 418, 934, 281, 550]. surgical [546]. survey [233, 248, 363, 153, 649]. survival [864, 641, 166, 126, 342, 209, 678, 510, 912, 235, 773, 170]. susceptibility [632]. Suspected [72, 364]. Sustained [763]. sustaining [137]. Sweden [600, 719]. Swedish [467, 909]. swim [416]. Swimbladder [659, 819, 823, 836, 774]. Swimming [707, 913, 932, 904, 121, 188, 381, 382, 838, 700, 105, 621, 449, 763]. Swiss [60]. Swiss/ICR [60]. Sydenham [137]. sympatric [104]. Symposium [233, 248]. synchronization [793]. syndrome [450]. synthesis [731]. synthetic [279]. syntype [407]. syntypes [91]. system [231, 393, 424, 55, 671, 284, 45, 845]. systematic [53]. Systematics [670].

T [246, 798]. tag [85, 882, 631, 103]. tagged [213]. tagging [444, 196, 213, 229, 230]. Tags [96, 685]. taken [655]. Tanganyika [43, 491, 618]. tanks [603]. tapetum [216]. taxonomic [412]. Tay [55]. Tay-Tummel-Garry [55]. Technical [25]. technique [470, 592, 166, 180, 757]. techniques [366]. Tees [393]. Teesdale [676]. Tegid [880, 346, 78, 408]. Teify [23]. telemetry [703, 793, 548]. teleost [835, 349, 746, 405, 682, 352, 106, 175, 758, 662, 755, 90, 347, 541, 848, 457, 817, 653, 112]. teleostean [405]. Teleostei [226, 331, 295, 264, 198, 186, 458, 154, 407, 398]. teleosts [868, 15, 256, 570, 509, 771, 620]. television [346]. Temephos [917, 776]. Temminck [308]. temperate [349]. Temperature [226, 224, 56, 65, 349, 260, 921, 354, 630, 732, 806, 89, 820, 423, 859, 898, 179, 547, 856, 893, 187, 678, 903, 912, 637, 193, 388, 666, 745, 480, 311, 585, 681]. temperatures [723, 179, 444, 235, 330]. Temporal [888, 716]. tensions [109]. Terapon [653]. term [523, 649]. terrestrial [72]. test [341]. tested [34]. testicular [746]. testis [750]. tests [283, 503, 725, 721]. testudineus [788]. tetra [243]. tetra-acetate [243]. Thames [269, 9, 64]. their [222, 511, 43, 313, 685, 53, 414, 502, 149, 477]. thelohani [825]. them [502]. theoretical [453]. Therapon [872]. therapy [510]. Thermal [330, 77, 694, 875, 358]. these [874]. thin [255]. Thompson [670]. thorny [459]. Thorogobius [186]. those [302]. threadfin [781, 827]. Three [680, 471, 916, 921, 234, 662, 736, 109, 674, 413, 563, 598, 927, 928, 929, 254, 358, 362, 160, 167, 238, 456]. Three-spined [680, 736, 109, 413, 563, 598, 167, 238, 456]. throat [168]. Thunberg [578]. Thymallus [66, 901, 408]. thymidine [619]. thyroid [56]. thyroidal [699, 701]. thyroxine [700]. Thys [446]. tidal [269, 420]. Ties [173]. tilapia [782, 65, 844, 866, 322, 379, 907, 245, 387, 354, 630, 591, 732, 723, 713, 574, 634, 250, 63, 33, 253, 281, 285, 665, 81, 543, 388, 480]. time [820, 514, 547, 330]. Tissue [713, 171, 4, 282, 905]. tissues [343, 915, 386, 496, 429]. Titicaca [206]. Tjeukemeer [402, 636, 766, 767, 911]. together [402]. tolerance [608, 226, 30, 256, 510, 224, 410, 417, 56]. tolerances [396]. topotypes [114]. Torpedo [876]. total [405, 64]. tow [390]. Toxic [507, 790, 657]. Toxicity [279, 539, 178, 935, 27, 574, 397, 635, 789, 563, 598, 725, 721, 414, 853, 168]. toxicology [175]. trabajos [118, 135, 146, 158]. Trachinus [431]. trachurs [503, 292]. Trachurus [503, 292, 525]. Tracking [275, 368, 706]. tract [177, 88, 910, 926, 263, 439, 594]. transfer [465, 756]. Transmission [360, 300]. transplantation [170]. transport [380]. Transvaal [355]. traps [2]. trawling [544]. treated [801, 327, 645]. treatment [517, 792, 211, 388]. Tregaron [23]. Trematoda [100, 31]. Trewavas [543]. tri [574]. tri-n-butyltin [574]. trial [79]. tributaries [394, 98, 23, 679, 779]. tricaine [561]. trifluoromethyl [564]. trifluoromethyl-4-nitrophenol [564]. Trigger [765]. Trinidad [729]. triploids [400]. tripterygiid [398]. Tripterygiidae [398]. Tripterygion [398]. Trisopterus [557]. tritylmorpholine [626]. trophic [866]. tropical [349, 532, 723, 375, 559, 626]. Trout [908, 139, 140, 507, 231, 677, 890, 60, 902, 392, 222, 783, 76, 260, 27, 735, 50, 599, 797, 565, 652, 879, 468, 528, 17, 641, 20, 475, 206, 640, 726, 92, 914, 302, 386, 419, 197, 39, 512, 728, 110, 792, 826, 133, 190, 194, 526, 606, 822, 923, 381, 382, 633, 445, 739, 740, 905, 811, 635, 610, 828, 654, 720, 5, 509, 159, 353, 558, 137, 779, 881, 922, 638, 522, 67, 919, 883, 510, 701, 467, 829, 120, 672, 485, 488, 265, 572, 689, 489, 548, 414, 426, 417, 637, 839, 184, 813]. trout [647, 741, 754, 466, 657, 168, 529, 505, 695, 474, 289, 714, 80, 142, 267, 268, 441, 778, 763, 745, 874, 490, 411, 885, 709, 586]. trutta [507, 76, 735, 50, 599, 797, 20, 475, 640, 726, 914, 302, 419, 197, 512, 728, 526, 606, 822, 923, 445, 739, 661, 740, 137, 779, 881, 638, 522, 67, 919, 883, 510, 793, 467, 672, 265, 572, 689, 541, 548, 426, 505, 695, 80, 142, 267, 268, 441, 885]. Trypanorhyncha [461, 367]. trypanosome [580]. tschawytscha [756]. tshawytscua [569]. Tubifex [605]. Tummel [55]. tuna [932]. turbot [450, 820, 613, 199, 62]. turnover [881]. TV [784]. twaite [668]. Tweed [599]. twins’ [717]. two [427, 872, 718, 748, 780, 715, 421, 687, 639, 8, 304, 389, 5, 240, 276, 435, 444, 467, 851, 52, 656, 706, 858, 71]. two-spot [427]. Tylodelphys [715, 639]. type [771]. types [381, 673, 849, 620].

U.D.N. [40]. U.D.N.-infected [40]. U.S.A. [129, 737]. UDN [32, 67, 300]. UDN-infected [32]. Uganda [399, 294]. ulcerative [67, 300, 29, 46, 70, 99, 426]. Ullswater [18]. Ultrasonic [706, 548, 368]. Ultrastructural [440, 634, 389]. Ultrastructure [437, 538, 432, 760, 107]. Ulvaria [384]. Umbra [854]. umbratus [163, 296, 293, 217]. Umbridae [854]. unbaited [2]. unicolor [872]. unicopula [3]. unisexual [716]. unrestrained [470]. unusual [588]. upland [334, 335, 336, 680, 863]. upon [676, 402, 110, 472, 242]. Upper [676, 799, 881, 193, 408, 759, 779]. upriver [703]. upstream [1, 324]. Uptake [555, 626, 783, 528]. urine [289]. Use [509, 703, 300, 68, 784, 742, 244, 359, 546, 535, 611, 685, 882, 533, 827, 86, 793, 725, 414]. used [390]. using [820, 583, 836, 757]. Utah [273]. utilisation [713]. utilization [596, 336].

V [288, 426, 400]. Val [30, 227, 893, 656, 594]. Val. [280]. Valencia [542]. Valenciennes [877, 607, 78, 332, 53, 516, 235]. validate [784]. Validation [212]. Valley [59]. Valparaíso [280]. value [243, 289, 34]. values [590, 110, 26, 327, 617, 745]. var. [452]. Variability [874]. Variation [632, 726, 536, 609, 730]. Variations [438, 394, 799, 840, 888, 553]. variety [159, 353]. various [877, 562, 694, 21, 666]. Vascular [772]. vasculosus [33]. vendace [82]. Venezuela [542]. Ventilation [921, 826, 793]. Vertebral [318]. vertebrate [171]. Vertical [82]. vesicles [11]. vetulus [737]. Vibrio [448, 101, 583, 521]. vibriosis [613, 521]. vicinity [223]. Victoria [860, 299, 294, 900]. villosus [882]. vipera [431]. virens [823, 836]. Virgin [193, 254]. virus [632, 60, 675, 316]. Vistula [84]. Vitamin [924, 689]. vitellogenesis [211]. vitreum [708, 708]. vitro [545]. viviparous [352, 758, 125, 126, 170]. vivo [463, 545, 212]. volumetric [51]. Volvi [855]. vulgaris [434, 799].

Wagner [338]. Walb [36]. Walbaum [569, 290, 894, 904, 121, 538, 568, 493, 824]. Wales [880, 346, 759, 78, 133, 132, 139, 140, 199, 779, 881, 672, 813, 291, 408]. walkeri [464, 531]. walleye [708]. warm [409, 483]. Washington [356]. Water [204, 502, 790, 812, 245, 387, 354, 234, 397, 539, 409, 512, 794, 308, 138, 923, 130, 602, 905, 472, 307, 637, 671, 929]. water-borne [539]. water-electrolyte [472]. waters [900, 151, 557, 575, 588, 628, 141, 525, 328, 429, 236, 447, 765, 239]. Weber [387]. weeds [34]. weever [431]. weight [497, 280, 376, 820, 325, 547, 855, 543, 441]. weights [468]. well [819]. well-developed [819]. Welsh [863]. West [102, 918, 557, 588, 628, 239, 150, 575, 642]. western [186, 234, 520, 398]. Westmorland [393]. which [495]. whirling [498, 534, 535]. white [825, 932, 363, 770, 633, 161, 56]. whitefish [875, 774, 920]. whiting [100, 806, 575, 588, 628, 501, 111, 181, 223, 429]. wigwag [919]. wild [101, 583, 548, 328, 744]. williamsoni [875]. willughbii [629]. Wilson [394]. Windermere [13, 629, 612, 16]. Winter [220, 905]. Winter-spring [220]. within [4, 43, 161]. without [327]. work [68]. works [118, 135, 146, 158]. wound [349, 597]. wrasse [587]. Wu [91]. Wye [759, 779, 881]. Wyland [129].

X [854, 73]. X-irradiation [73]. X-radiation [854]. Xiphophorus [125, 126, 170].

year [687, 803, 201]. years [376, 679]. yellowfish [176]. yellowtail [897]. yield [618, 402, 342, 453]. yolk [731]. Yorkshire [297, 80]. young [591, 314, 865, 302, 423, 559, 199, 3, 64, 522, 743, 181, 16]. Yugoslavia [374].

Zambia [818]. Zealand [434, 334, 335, 336]. Zebra [721, 891]. zebrafish [533]. zilli [250]. zillii [907, 81]. Zinc [792, 790, 499, 27, 397, 175, 648, 889, 789, 228, 555, 747]. Zululand [868, 787].

References

Alabaster:1970:RFU

[1] John S. Alabaster. River flow and upstream movement and catch of migratory salmonids. Journal of Fish Biology, 2(1):1–13, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1970:SFA

[2] Brian Stott. Some factors affecting the catching power of unbaited fish traps. Journal of Fish Biology, 2(1):15–22, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackenzie:1970:GUG

[3] K. Mackenzie. Gyrodactylus unicopula Glukhova, 1955, from young plaice Pleuronectes platessa L. with notes on the ecology of the parasite. Journal of Fish Biology, 2(1):23–34, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arme:1970:OTR

[4] C. Arme and R. Wynne Owen. Observations on a tissue response within the body cavity of fish infected with the plerocercoid larvae of Ligula intestinalis (L.) (Cestoda: Pseudophyllidea). Journal of Fish Biology, 2 (1):35–37, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackay:1970:PTG

[5] David W. Mackay. Populations of trout and grayling in two Scottish rivers. Journal of Fish Biology, 2(1):39–45, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:1970:EMS

[6] P. Martin and N. R. Bromage. The effect of methallibure on spermatogenesis in Poecilia reticulata. Journal of Fish Biology, 2(1):47–51, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huggins:1970:CSB

[7] R. J. Huggins and Anne Thompson. Communal spawning of brook and river lampreys, Lampetra planeri Bloch and Lampetra fluviatilis L. Journal of Fish Biology, 2(1):53–54, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1970:MDB

[8] C. L. Lee and W. D. Williams. Meristic differences between two conspecific fish populations in Australian salt lakes. Journal of Fish Biology, 2(1):55– 56, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mathews:1970:ISC

[9] C. P. Mathews. Immersion staining of coarse fish in the Thames. Journal of Fish Biology, 2(1):57–58, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1970:GCF

[10] Alwyne Wheeler. Gobius cmentatus — a fish new to the northern European fauna. Journal of Fish Biology, 2(1):59–67, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nayyar:1970:RSV

[11] S. K. Nayyar and Bangalore I. Sundararaj. Response of the seminal vesicles of the hypophysectomized catfish Heteropneustes fossilis (Bloch) to exogenous androgens, estrogen and corticosteroid. Journal of Fish Biology, 2(1):69–78, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1970:BRa

[12] Anonymous. Book reviews. Journal of Fish Biology, 2(1):79–82, January 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagenal:1970:HRF

[13] T. B. Bagenal. An historical review of the fish and fisheries investigations of the Freshwater Biological Association, mainly at the Windermere Laboratory. Journal of Fish Biology, 2(2):83–101, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1970:BSD

[14] R. J. Roberts, J. Leckie, and H. D. Slack. ‘Bald spot’ disease in powan Coregonus lavaretus (L.). Journal of Fish Biology, 2(2):103–105, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Culkin:1970:FAS

[15] F. Culkin and R. J. Morris. The fatty acids of some marine teleosts. Journal of Fish Biology, 2(2):107–112, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wllloughby:1970:MAD

[16] L. G. Wllloughby. Mycological aspects of a disease of young perch in Windermere. Journal of Fish Biology, 2(2):113–116, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egglishaw:1970:PST

[17] Henry J. Egglishaw. Production of salmon and trout in a stream in Scotland. Journal of Fish Biology, 2(2):117–136, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagenal:1970:NBS

[18] T. B. Bagenal. Notes on the biology of the schelly Coregonus lavaretus (L.) in Haweswater and Ullswater. Journal of Fish Biology, 2(2):137– 154, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beukemaj:1970:AHA

[19] J. J. Beukemaj. Acquired hook-avoidance in the pike Esox lucius L. fished with artificial and natural baits. Journal of Fish Biology, 2(2):155–160, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:1970:DCI

[20] J. M. Elliott. Diel changes in invertebrate drift and the food of trout Salmo trutta L. Journal of Fish Biology, 2(2):161–165, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Molnar:1970:SHC

[21] Gy. Molnár and E. Tamássy. Study of the haemoglobin content of a single erythrocyte (M index) in various cultured fish species. Journal of Fish Biology, 2(2):167–171, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardisty:1970:RGD

[22] M. W. Hardisty. The relationship of gonadal development to the life cycles of the paired species of lamprey, Lampetra fluviatilis (L.) and Lampetru planeri (Bloch). Journal of Fish Biology, 2(2):173–181, April 1970. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1970:SSP

[23] A. Neville Jones. A study of salmonid populations of the River Teify and tributaries near tregaron. Journal of Fish Biology, 2(2):183–197, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1970:BRb

[24] Anonymous. Book reviews. Journal of Fish Biology, 2(2):199–201, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1970:FTC

[25] Anonymous. FAO Technical Conference on Marine Pollution and Its Effects on Living Resources and Fishing. Journal of Fish Biology, 2(2):202, April 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulcahy:1970:BVP

[26] M. F. Mulcahy. Blood values in the pike Esox lucius L. Journal of Fish Biology, 2(3):203–209, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1970:ALT

[27] V. M. Brown and R. A. Dalton. The acute lethal toxicity to rainbow trout of mixtures of copper, phenol, zinc and nickel. Journal of Fish Biology, 2 (3):211–216, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marlborough:1970:SBL

[28] D. Marlborough. The status of the burbot Lota lota (L.) (Gadidae) in Britain. Journal of Fish Biology, 2(3):217–222, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1970:SUDa

[29] Ronald J. Roberts, W. M. Shearer, K. G. R. Elson, and A. L. S. Munro. Studies on ulcerative dermal necrosis of salmonids: 1. The skin of the normal salmon head. Journal of Fish Biology, 2(3):223–229, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1970:TGC

[30] D. G. Cross. The tolerance of grass carp Ctenopharyngodon idella (Val.) to seawater. Journal of Fish Biology, 2(3):231–233, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1970:HSF

[31] Fah Onn Lee and Thomas C. Cheng. The histochemistry of Stellantchasmus falcatus Onji and Nishio, 1915 (Trematoda: Heterophyidae) metacercarial cyst in the mullet Mugil cephalus L. and histopathological alterations in the host. Journal of Fish Biology, 2(3):235–243, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hodkinson:1970:GCS

[32] M. Hodkinson and A. Hunter. Growth control of Saprolegnia from UDNinfected Atlantic salmon Salmo salar L. Journal of Fish Biology, 2(3): 245–248, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1970:DSV

[33] W. J. R. Lanzing. The development of a saccus vasculosus in Tilapia mossambica Peters and other cichlids. Journal of Fish Biology, 2(3):249– 252, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tan:1970:CNV

[34] Yee Thong Tan. Composition and nutritive value of some grasses, plants and aquatic weeds tested as diets. Journal of Fish Biology, 2(3):253– 257, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:1970:BES

[35] J. S. Campbell and H. R. MacCrimmon. Biology of the emerald shiner Notropis atherinoides Rafinesque in Lake Simcoe, Canada. Journal of Fish Biology, 2(3):259–273, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deGroot:1970:SNA

[36] S. J. de Groot. Some notes on an ambivalent behaviour of the Greenland halibut Reinhardtius hippoglossoides (Walb.) Pisces: Pleuronectiformes. Journal of Fish Biology, 2(3):275–279, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1970:OBG

[37] R. N. Gibson. Observations on the biology of the giant goby Gobius cobitis Pallas. Journal of Fish Biology, 2(3):281–288, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1970:BRc

[38] Anonymous. Book reviews. Journal of Fish Biology, 2(3):289–292, July 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herman:1970:EGR

[39] Roger Lee Herman. Effects of gossypol on rainbow trout Salmo guirdneri Richardson. Journal of Fish Biology, 2(4):293–303, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hodkinson:1970:IRU

[40] Madeline Hodkinson and Alan Hunter. Immune response of U.D.N.- infected salmon to Saprolegnia. Journal of Fish Biology, 2(4):305–311, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moravec:1970:LHN

[41] F. Moravec. On the life history of the nematode Raphidascaris acus (Bloch, 1779) in the natural environment of the River Bystřice, Czechoslovakia. Journal of Fish Biology, 2(4):313–322, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1970:NCS

[42] Alwyne Wheeler. Notes on a collection of shore fishes from Guernsey, Channel Islands. Journal of Fish Biology, 2(4):323–328, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulter:1970:PCW

[43] G. W. Coulter. Population changes within a group of fish species in Lake Tanganyika following their exploitation. Journal of Fish Biology, 2(4): 329–353, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shehadeh:1970:ISS

[44] Ziad H. Shehadeh and James N. Ellis. Induced spawning of the striped mullet Mugil cephatus L. Journal of Fish Biology, 2(4):355–360, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walkey:1970:ABE

[45] M. Walkey and R. H. Meakins. An attempt to balance the energy budget of a host-parasite system. Journal of Fish Biology, 2(4):361–372, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1970:SUDb

[46] Ronald J. Roberts, W. M. Shearer, A. L. S. Munro, and K. G. R. Elson. Studies on ulcerative dermal necrosis of salmonids: II. The sequential pathology of the lesions. Journal of Fish Biology, 2(4):373–378, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mawdesley-Thomas:1970:FSB

[47] L. E. Mawdesley-Thomas. The Fisheries Society of the British Isles. Journal of Fish Biology, 2(4):379–381, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1970:APP

[48] Anonymous. Abstracts of papers presented at the Spring Meeting 24 April 1970. Journal of Fish Biology, 2(4):382–384, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1970:BRd

[49] Anonymous. Book reviews. Journal of Fish Biology, 2(4):385–386, October 1970. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:1971:GBT

[50] R. N. Campbell. The growth of brown trout Salmo trutta L. in northern Scottish lochs with special reference to the improvement of fisheries. Journal of Fish Biology, 3(1):1–28, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellawell:1971:RVM

[51] J. M. Hellawell and R. Abel. A rapid volumetric method for the analysis of the food of fishes. Journal of Fish Biology, 3(1):29–37, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shafi:1971:AGP

[52] Muhammed Shafi and Peter S. Maitland. The age and growth of perch (Perca fluviatilis L.) in two Scottish lochs. Journal of Fish Biology, 3(1): 39–57, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollard:1971:SSM

[53] D. A. Pollard and P. Plchot. The systematic status of the Mediterranean centracanthid fishes of the genus Spicara, and in particular S. chryselis (Valenciennes), as indicated by electrophoretic studies of their eye-lens proteinst. Journal of Fish Biology, 3(1):59–72, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braithwaite:1971:SFC

[54] H. Braithwaite. A sonar fish counter. Journal of Fish Biology, 3(1):73– 82, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardner:1971:RCM

[55] Michael L. G. Gardner. Recent changes in the movements of adult salmon Salmo salar L. in the tay-tummel-garry system, Scotland. Journal of Fish Biology, 3(1):83–96, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

erlean:1971:TTT

[56] Andrew J. Mc erlean and H. J. Brinkley. Temperature tolerance and thyroid activity of the white perch Roccus (= Morone) americanus. Journal of Fish Biology, 3(1):97–114, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1971:BRa

[57] Anonymous. Book reviews. Journal of Fish Biology, 3(1):115–121, January 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hines:1971:SEF

[58] R. Hines and A. Yashouv. Some environmental factors ifluencing the activity of spermatozoa of Mugil capito Cuvier, a grey mullet. Journal of Fish Biology, 3(2):123–127, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:1971:DRR

[59] R. S. Wilson. The decline of a roach Rutilus rutilus (L.) population in Chew Valley Lake. Journal of Fish Biology, 3(2):129–137, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Angiolelli:1971:SIA

[60] R. F. Angiolelli and G. J. Rio. The Swiss/ICR (Ha) albino mouse as an experimental host for infectious pancreatic necrosis virus of trout. Journal of Fish Biology, 3(2):139–144, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:1971:OAC

[61] A. D. Campbell. The occurrence of Argulus (Crustacea: Branchiura) in Scotland. Journal of Fish Biology, 3(2):145–146, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deGroot:1971:BSM

[62] S. J. de Groot. Bothriocephalus scorpii (Müller) (Cestoda: Pseudophyllidea) in turbot Scophthalmus maximus (L.) and brill S. rhombus (L.) from the southern North Sea. Journal of Fish Biology, 3(2):147–149, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Facade:1971:FFH

[63] S. O. Facade. The food and feeding habits of Tilapia species in the Lagos Lagoon. Journal of Fish Biology, 3(2):151–156, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mathews:1971:CYF

[64] C. P. Mathews. Contribution of young fish to total production of fish in the River Thames near Reading. Journal of Fish Biology, 3(2):157– 180, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allanson:1971:IEL

[65] B. R. Allanson, A. Bok, and N. I. Van Wyk. The influence of exposure to low temperature on Tilapia mossambica Peters (Cichlidae). II. Changes in serum osmolarity, sodium and chloride ionconcentrations. Journal of Fish Biology, 3(2):181–185, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellawell:1971:FGT

[66] J. M. Hellawell. The food of the grayling Thymallus thymallus (L.) of the River Lugg, Herefordshire. Journal of Fish Biology, 3(2):187–197, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulcahy:1971:SPC

[67] Maire F. Mulcahy. Serum protein changes associated with ulcerative dermal necrosis (UDN) in the trout Salmo trutta L. Journal of Fish Biology, 3(2):199–201, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abram:1971:AEA

[68] F. S. H. Abram. An air-economizer apparatus for use in fish work. Journal of Fish Biology, 3(2):203–205, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagenal:1971:ISF

[69] T. B. Bagenal. The interrelation of the size of fish eggs, the date of spawning and the production cycle. Journal of Fish Biology, 3(2):207– 219, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1971:SUD

[70] Ronald J. Roberts, H. J. Ball, A. L. S. Munro, and W. M. Shearer. Studies on ulcerative dermal necrosis of salmonids: III. The healing process in fish maintained under experimental conditions. Journal of Fish Biology, 3(2): 221–224, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Western:1971:FDT

[71] J. R. H. Western. Feeding and digestion in two cottid fishes, the freshwater Cottus gobio L. and the marine Enophrys bubalis (Euphrasen). Journal of Fish Biology, 3(2):225–246, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eldon:1971:STD

[72] G. A. Eldon. Suspected terrestrial deposition of eggs by Neochanna apoda Günther (Pisces: Salmoniformes: Galaxiidae). Journal of Fish Biology, 3 (2):247–249, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:1971:EAX

[73] Eileen Ward, S. A. Beach, and E. D. Dyson. The effect of acute Xirradiation on the development of the plaice Pleuronectes platessa L. Journal of Fish Biology, 3(2):251–259, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1971:BRb

[74] Anonymous. Book reviews. Journal of Fish Biology, 3(2):261–262, April 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slnha:1971:ISC

[75] V. R. P. Slnha. Induced spawning in carp with fractionated fish pituitary extract. Journal of Fish Biology, 3(3):263–272, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolis:1971:DGP

[76] L. Bolis, P. Luly, and V. Baroncelli. D(+)-Glucose permeability in brown trout Salmo trutta L. erythrocytes. Journal of Fish Biology, 3(3):273– 275, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deVlaming:1971:TSB

[77] Victor L. de Vlaming. Thermal selection behaviour in the estuarine goby Gillichthys mirabilis Cooper. Journal of Fish Biology, 3(3):277–286, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haram:1971:SOF

[78] O. J. Haram and J. W. Jones. Some observations on the food of the gwyniad Coregonus clupeoides pennantii Valenciennes of Llyn Tegid (Lake Bala), North Wales. Journal of Fish Biology, 3(3):287–295, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crossland:1971:FTM

[79] N. O. Crossland, A. J. A. Pearson, and M. S. Bennet. A field trial with the molluscicide Frescon for control of Lymnaea peregra Müller, snail host of Diplostomum spathaceum (Rudolphi). Journal of Fish Biology, 3(3): 297–302, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Templeton:1971:IAA

[80] R. G. Templeton. An investigation of the advantages of autumn and spring stocking with brown trout Salmo trutta L. in a Yorkshire reservoir.

Journal of Fish Biology, 3(3):303–324, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payne:1971:ECT

[81] A. I. Payne. An experiment on the culture of Tilapia esculenta (Graham) and Tilapia zillii (Gervais) (Cichlidae) in fish ponds. Journal of Fish Biology, 3(3):325–340, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dembinski:1971:VDV

[82] W. Dembiński. Vertical distribution of vendace Coregonus albula L. and other pelagic fish species in some Polish lakes. Journal of Fish Biology, 3 (3):341–357, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1971:BRc

[83] Anonymous. Book reviews. Journal of Fish Biology, 3(3):359–368, July 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Backiel:1971:PFC

[84] Tadeusz Backiel. Production and food consumption of predatory fish in the Vistula River. Journal of Fish Biology, 3(4):369–405, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1971:RRT

[85] C. R. Goldspink and J. W. Banks. A readily recognizable tag for marking bream Abramis brama (L.). Journal of Fish Biology, 3(4):407–411, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minaur:1971:SEO

[86] J. Minaur. A simple electrically-operated feeder for use in salmon rearing. Journal of Fish Biology, 3(4):413–415, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bootsma:1971:HRD

[87] R. Bootsma. Hydrocephalus and red-disease in pike fry Esox lucius L. Journal of Fish Biology, 3(4):417–419, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucke:1971:AHA

[88] David Bucke. The anatomy and histology of the alimentary tract of the carnivorous fish the pike Esox lucius L. Journal of Fish Biology, 3(4):421– 431, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1971:ETR

[89] D. J. Edwards. Effect of temperature on rate of passage of food through the alimentary canal of the plaice Pleuronectes platessa L. Journal of Fish Biology, 3(4):433–439, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Narasimhan:1971:ESC

[90] P. V. Narasimhan and Bangalore I. Sundararaj. Effects of stress on carbohydrate metabolism in the teleost Notopterus notopterus (Pallas). Journal of Fish Biology, 3(4):441–451, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trewavas:1971:SSC

[91] Ethelwynn Trewavas. The syntypes of the sciaenid Corvina albida Cuvier and the status of Dendrophysa hooghliensis Sinha and Rao and Nibea coibor (nec Hamilton) of Chu, Lo & Wu. Journal of Fish Biology, 3(4): 453–461, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finn:1971:IRR

[92] J. Peter Finn and N. O. Nielson. The inflammatory response of rainbow trout. Journal of Fish Biology, 3(4):463–478, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gyldenholm:1971:CNF

[93] A. O. Gyldenholm and J. J. Scheel. Chromosome numbers of fishes. I. Journal of Fish Biology, 3(4):479–486, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mawdesley-Thomas:1971:FSB

[94] Lionel E. Mawdesley-Thomas. Fisheries society of the British Isles: Secretary’s report, 1970–71. Journal of Fish Biology, 3(4):487–488, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1971:BRd

[95] Anonymous. Book reviews. Journal of Fish Biology, 3(4):489–493, October 1971. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1972:FPB

[96] David I. Gibson. Flounder parasites as biological tags. Journal of Fish Biology, 4(1):1–9, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trewavas:1972:GRS

[97] Ethelwynn Trewavas and P. K. Talwar. On the generic relationship of the sciaenid fish Bola chaptis Hamilton, with a description of specimens from Burma. Journal of Fish Biology, 4(1):11–16, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cragg-Hine:1972:FNS

[98] D. Cragg-Hine. Further note on the salmonid populations of the Lough Erne tributaries in Northern Ireland. Journal of Fish Biology, 4(1):17– 20, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1972:SUD

[99] Ronald J. Roberts, W. M. Shearer, and A. L. S. Munro. Studies on ulcerative dermal necrosis of salmonids: IV. Failure to detect epithelial auto-antibodies in sera from diseased fish. Journal of Fish Biology, 4(1): 21–25, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arme:1972:OOD

[100] C. Arme and D. W. Halton. Observations on the occurrence of Diclidophora merlangi (Trematoda: Monogenea) on the gills of whiting, Gadus merlangus. Journal of Fish Biology, 4(1):27–32, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haastein:1972:OVD

[101] T. Haastein and G. Holt. The occurrence of vibrio disease in wild Norwegian fish. Journal of Fish Biology, 4(1):33–37, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welcomme:1972:EAM

[102] R. L. Welcomme. An evaluation of the acadja method of fishing as practised in the coastal lagoons of Dahomey (West Africa). Journal of Fish Biology, 4(1):39–55, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:1972:SST

[103] A. H. Young, P. Tytler, F. G. T. Holliday, and A. MacFarlane. A small sonic tag for measurement of locomotor behaviour in fish. Journal of Fish Biology, 4(1):57–65, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fisher:1972:KAS

[104] Paul B. Fisher and Joseph W. Rachlin. Karyotypic analysis of 2 sympatric species of fish, Fundulus heteroclitus L. and F. majalis. Journal of Fish Biology, 4(1):67–71, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurence:1972:CSA

[105] Geoffrey C. Laurence. Comparative swimming abilities of fed and starved larval largemouth bass (Micropterus salmoides). Journal of Fish Biology, 4(1):73–78, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drujan:1972:AIC

[106] B. D. Drujan and J. M. Díz Borges. Adrenaline induced changes in the metabolism of glycogen in the teleost retina. Journal of Fish Biology, 4(1):79–85, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1972:SSP

[107] Ronald J. Roberts, Hamish Young, and J. A. Milne. Studies on the skin of plaice (Pleuronectes platessa L.) 1. The structure and ultrastructure of normal plaice skin. Journal of Fish Biology, 4(1):87–98, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palacios:1972:PCC

[108] L. Palacios, M. Rubio, and J. Planas. Plasma chemical composition in the goosefish (Lophius piscatorius, l). Journal of Fish Biology, 4(1):99– 102, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1972:SEL

[109] David B. Lewis, M. Walkey, and H. J. G. Dartnall. Some effects of low oxygen tensions on the distribution of the three-spined stickleback Gasterosteus aculeatus L. and the nine-spined stickleback Pungitius pungitius (L.). Journal of Fish Biology, 4(1):103–108, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Houston:1972:SOU

[110] Arthur H. Houston and Mary Anne DeWilde. Some observations upon the relationship of microhaematocrit values to haemoglobin concentrations and erythrocyte numbers in the carp Cyprinus carpio L. and brook trout Salvelinus fontinalis (Mitchill). Journal of Fish Biology, 4(1):109– 115, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shotter:1972:NHP

[111] R. A. Shotter. Notes on helminth parasites of the whiting Odontogadus m. merlangus (L.) from the northern Irish Sea. Journal of Fish Biology, 4(1):117–130, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deVlaming:1972:ECT

[112] Victor L. de Vlaming. Environmental control of teleost reproductive cycles: a brief review. Journal of Fish Biology, 4(1):131–140, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1972:RLK

[113] Alwyne Wheeler and R. W. Blacker. Rare and little-known fishes in British seas in 1968 and 1969. Journal of Fish Biology, 4(1):141–170, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Talwar:1972:NTB

[114] P. K. Talwar and A. K. Datta. A neotype and topotypes for Bola chaptis Hamilton, 1822 (Pisces, Sciaenidae). Journal of Fish Biology, 4(1):171– 172, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:FSB

[115] Anonymous. Fisheries society of the British Isles: Fish Pathology Group. Journal of Fish Biology, 4(1):173, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:ON

[116] Anonymous. Obituary notice. Journal of Fish Biology, 4(1):175, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:BRa

[117] Anonymous. Book reviews. Journal of Fish Biology, 4(1):177–186, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:RENa

[118] Anonymous. Resumenes en español de los trabajos publicados en este número. (Spanish) [Summaries in Spanish of the works published in this issue]. Journal of Fish Biology, 4(1):187–192, January 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hawkins:1972:FHB

[119] Rosemary I. Hawkins and Lionel E. Mawdesley-Thomas. Fish haematology — a bibliography. Journal of Fish Biology, 4(2):193–232, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payne:1972:ENH

[120] R. H. Payne, A. R. Child, and A. Forrest. The existence of natural hybrids between the European trout and the Atlantic salmon. Journal of Fish Biology, 4(2):233–236, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harvey:1972:CSK

[121] H. H. Harvey and C. R. Bothern. Compensatory swimming in the Kokanee and sockeye salmon Oncorhynchus nerka (Walbaum). Journal of Fish Biology, 4(2):237–247, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biro:1972:NFL

[122] Péter Biró. Neogobius fluviatilis in Lake Balaton — a Ponto-Caspian goby new to the fauna of central Europe. Journal of Fish Biology, 4(2):249– 255, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weiss:1972:SES

[123] D. M. Weiss. A simple electronic smolt counter. Journal of Fish Biology, 4 (2):257–260, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yasutake:1972:SAP

[124] William T. Yasutake and Donald F. Amend. Some aspects of pathogenesis of infectious hematopoietic necrosis (IHN). Journal of Fish Biology, 4(2):

261–264, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogarth:1972:IRBa

[125] P. J. Hogarth. Immune relations between mother and foetus in the viviparous poeciliid fish Xiphophorus hellerii Haeckel: I. Antigenicity of the foetus. Journal of Fish Biology, 4(2):265–269, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogarth:1972:IRBb

[126] P. J. Hogarth. Immune relations between mother and foetus in the viviparous poeciliid fish Xiphophorus hellerii Haeckel: II. Lack of status of the ovary as a favourable site for allograft survival. Journal of Fish Biology, 4(2):271–275, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Badcock:1972:AAM

[127] J. R. Badcock and N. R. Merrett. On Argyripnus atlanticus, Maul 1952 (Pisces, Stomiatoidei), with a description of post-larval forms. Journal of Fish Biology, 4(2):277–287, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bibby:1972:PBH

[128] M. C. Bibby. Population biology of the helminth parasites of Phoxinus phoxinus (L.), the minnow, in a Cardiganshire lake. Journal of Fish Biology, 4(2):289–300, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gerking:1972:RFC

[129] Shelby D. Gerking. Revised food consumption estimate of a bluegill sunfish population in Wyland Lake, Indiana, U.S.A. Journal of Fish Biology, 4(2):301–308, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1972:NIC

[130] J. W. Jones and R. S. J. Linfield. A note on the identification of a crucian carp Carassius camssim (L.) 3× common carp Cyprinus carpio (L.) hybrid from a still water in Nottinghamshire. Journal of Fish Biology, 4(2):309– 310, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palmer:1972:SOG

[131] T. N. Palmer and Brenda E. Ryman. Studies on oral glucose intolerance in fish. Journal of Fish Biology, 4(2):311–319, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1972:LFL

[132] P. C. Hunt and J. W. Jones. The littoral fauna of Llyn Celyn, North Wales. Journal of Fish Biology, 4(2):321–331, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1972:FBT

[133] P. C. Hunt and J. W. Jones. The food of brown trout in Llyn Alaw, Anglesey, North Wales. Journal of Fish Biology, 4(2):333–352, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:BRb

[134] Anonymous. Book reviews. Journal of Fish Biology, 4(2):353–355, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:RENb

[135] Anonymous. Resumenes en español de los trabajos publicados en este número. (Spanish) [Summaries in Spanish of the works published in this issue]. Journal of Fish Biology, 4(2):357–361, April 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mathews:1972:GAC

[136] C. P. Mathews and W. P. Williams. Growth and annual check formation in scales of dace, Leuciscus leuciscm (L.). Journal of Fish Biology, 4(3): 363–367, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Millard:1972:ECS

[137] Timothy J. Millard and Hugh R. MacCrimmon. Evaluation of the contribution of supplemental plantings of brown trout Salmo trutta (L.) to a self-sustaining fishery in the Sydenham River, Ontario, Canada. Journal of Fish Biology, 4(3):369–384, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1972:EWL

[138] P. C. Hunt and J. W. Jones. The effect of water level fluctuations on a littoral fauna. Journal of Fish Biology, 4(3):385–394, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1972:TLAa

[139] P. C. Hunt and J. W. Jones. Trout in Llyn Alaw, Anglesey, North Wales. I. Population structure and angling returns. Journal of Fish Biology, 4(3): 395–408, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1972:TLAb

[140] P. C. Hunt and J. W. Jones. Trout in Llyn Alaw, Anglesey, North Wales. II. Growth. Journal of Fish Biology, 4(3):409–424, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1972:SAB

[141] Michael Kennedy and Patrick Fitzmaurice. Some aspects of the biology of gudgeon Gobio gobio (L.) in Irish waters. Journal of Fish Biology, 4(3): 425–440, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1972:AEB

[142] J. E. Thorpe and R. J. Roberts. An aeromonad epidemic in the brown trout (Salmo trutta L.). Journal of Fish Biology, 4(3):441–451, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weiss:1972:HPP

[143] D. M. Weiss. A high power pulse generator for electric fishing. Journal of Fish Biology, 4(3):453–459, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:BRc

[144] Anonymous. Book reviews. Journal of Fish Biology, 4(3):461–465, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:FPG

[145] Anonymous. Fish Pathology Group. Journal of Fish Biology, 4(3):466, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:RENc

[146] Anonymous. Resumenes en español de los trabajos publicados en este número. (Spanish) [Summaries in Spanish of the works published in this issue]. Journal of Fish Biology, 4(3):467–468, July 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellawell:1972:GRF

[147] J. M. Hellawell. The growth, reproduction and food of the roach Rutilus rutilus (L.), of the River Lugg, Herefordshire. Journal of Fish Biology, 4(4):469–486, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilkins:1972:BGAa

[148] Noël P. Wilkins. Biochemical genetics of the Atlantic salmon Salmo salar L. I. A review of recent studies. Journal of Fish Biology, 4(4):487–504, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilkins:1972:BGAb

[149] Noël P. Wilkins. Biochemical genetics of the Atlantic salmon Salmo salar L. II. The significance of recent studies and their application in population identification. Journal of Fish Biology, 4(4):505–517, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Facade:1972:BWA

[150] S. O. Facade and C. I. O. Olaniyan. The biology of the West African shad Ethmalosa fimbriata (Bowdich) in the Lagos Lagoon, Nigeria. Journal of Fish Biology, 4(4):519–533, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davey:1972:IAS

[151] John T. Davey. The incidence of Anisakis sp. larvae (Nematoda: Ascaridata) in the commercially exploited stocks of herring (Clupea harengus L., 1758), (Pisces: Clupeidae) in British and adjacent waters. Journal of Fish Biology, 4(4):535–554, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1972:OBP

[152] J. Hattingh. Observations on the blood physiology of five South African freshwater fish. Journal of Fish Biology, 4(4):555–563, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vooren:1972:EAI

[153] C. M. Vooren. Ecological aspects of the introduction of fish species into natural habitats in Europe, with special reference to The Netherlands: a literature survey. Journal of Fish Biology, 4(4):565–583, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sinha:1972:GHR

[154] V. R. P. Sinha. Gonadal hydration response and major electrolyte levels in serum and seminal fluid of Puntius gonionotus (Teleostei) injected with the fractionated fish pituitary extract. Journal of Fish Biology, 4(4):585– 592, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1972:HAH

[155] Peter C. Blaxhall. The haematological assessment of the health of freshwater fish: a review of selected literature. Journal of Fish Biology, 4 (4):593–604, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:BRd

[156] Anonymous. Book reviews. Journal of Fish Biology, 4(4):605–609, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:SA

[157] Anonymous. Special announcement. Journal of Fish Biology, 4(4):610, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1972:RENd

[158] Anonymous. Resumenes en español de los trabajos publicados en este número. (Spanish) [Summaries in Spanish of the works published in this issue]. Journal of Fish Biology, 4(4):611–613, October 1972. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:1973:SBP

[159] D. H. McCarthy, J. P. Stevenson, and M. S. Roberts. Some blood parameters of the rainbow trout (Salmo gairdneri Richardson) I. The Kamloops variety. Journal of Fish Biology, 5(1):1–8, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wickins:1973:SDP

[160] J. F. Wickins and I. S. Macfarlane. Some differences in the parasitic fauna of three samples of plaice (Pleuronectesplatessa L.) from the southern North Sea. Journal of Fish Biology, 5(1):9–19, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joy:1973:OIR

[161] James E. Joy and Larry P. Jones. Observations on the inflammatory response within the dermis of a white bass, Morone chrysops (Rafinesque), infected with Lernaea cruciata (Copepoda: Caligidea). Journal of Fish Biology, 5(1):21–23, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bayley:1973:SMC

[162] P. B. Bayley. Studies on the migratory characin, Prochilodus platensis Holmberg 1889, (Pisces, Characoidei) in the River Pilcomayo, South America. Journal of Fish Biology, 5(1):25–40, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1973:PSP

[163] J. Hattingh and P. J. du Toit. Partial separation of the pituitary proteins of Labeo umbratus, Smith (mudfish). Journal of Fish Biology, 5(1):41– 47, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1973:SFF

[164] Alwyne Wheeler and Peter S. Maitland. The scarcer freshwater fishes of the British Isles: I. Introduced species. Journal of Fish Biology, 5(1):49– 68, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munro:1973:SSC

[165] J. L. Munro, V. C. Gaut, R. Thompson, and P. H. Reeson. The spawning seasons of Caribbean reef fishes. Journal of Fish Biology, 5(1):69–84, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1973:SEB

[166] G. S. Harris. A simple egg box planting technique for estimating the survival of eggs deposited in stream gravel. Journal of Fish Biology, 5(1): 85–88, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1973:ESF

[167] R. J. Wootton. The effect of size of food ration on egg production in the female three-spined stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 5(1):89–96, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1973:STC

[168] Charlie E. Smith, Jamieson E. Holway, and Glen L. Hammer. Sulphamerazine toxicity in cut-throat trout broodfish Salmo clarki (Richardson). Journal of Fish Biology, 5(1):97–101, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1973:SSP

[169] Ronald J. Roberts, Marcus Bell, and Hamish Young. Studies on the skin of plaice (Pleuronectes platessa L.) II. The development of larval plaice skin. Journal of Fish Biology, 5(1):103–108, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogarth:1973:IRB

[170] P. J. Hogarth. Immune relations between mother and foetus in the viviparous poeciliid fish Xiphophorus hellerii Haeckel: III. Survival of embryos after ectopic transplantation. Journal of Fish Biology, 5(1):109– 113, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mawdesley-Thomas:1973:TRP

[171] L. E. Mawdesley-Thomas and D. Bucke. Tissue repair in a poikilothermic vertebrate, Carassius auratus (L.): a preliminary study. Journal of Fish Biology, 5(1):115–119, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:BRa

[172] Anonymous. Book reviews. Journal of Fish Biology, 5(1):121–123, January 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:T

[173] Anonymous. Ties. Journal of Fish Biology, 5(1):124, January 1973. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jimenez:1973:PPG

[174] Maximo Jimenez and Jose Planas. Plasma proteins of the goosefish Lophius piscatorius (L.). Journal of Fish Biology, 5(2):125–130, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eisler:1973:ATE

[175] Ronald Eisler and George R. Gardner. Acute toxicology to an estuarine teleost of mixtures of cadmium, copper and zinc salts. Journal of Fish Biology, 5(2):131–142, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulder:1973:RAF

[176] P. F. S. Mulder and G. W. Franke. A report on the artificial fertilization of the smallmouth yellowfish, Barbus holubi (Steindachner, 1894). Journal of Fish Biology, 5(2):143–145, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braber:1973:MAT

[177] L. Braber and S. J. de Groot. On the morphology of the alimentary tract of flatfishes (Pleuronectiformes). Journal of Fish Biology, 5(2):147– 153, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meadows:1973:TRS

[178] B. S. Meadows. Toxicity of rotenone to some species of coarse fish and invertebrates. Journal of Fish Biology, 5(2):155–163, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horoszewicz:1973:LDT

[179] Lidia Horoszewicz. Lethal and ‘disturbing’ temperatures in some fish species from lakes with normal and artificially elevated temperature. Journal of Fish Biology, 5(2):165–181, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moriarty:1973:TEE

[180] C. Moriarty. A technique for examining eel otoliths. Journal of Fish Biology, 5(2):183–184, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shotter:1973:CPFa

[181] R. A. Shotter. A comparison of the parasite fauna of young whiting, Odontogadus merlangus (L.) (Gadidae) from an inshore and an offshore location off the Isle of Man. Journal of Fish Biology, 5(2):185–195, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egglishaw:1973:EFG

[182] Henry J. Egglishaw and P. E. Shackley. An experiment on faster growth of salmon Salmo salar (L.) in a Scottish stream. Journal of Fish Biology, 5(2):197–204, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Facade:1973:FFI

[183] S. O. Facade and C. I. O. Olaniyan. The food and feeding interrelationship of the fishes in the lagos lagoon. Journal of Fish Biology, 5(2):205–225, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schreck:1973:POL

[184] Carl B. Schreck, Robert T. Lackey, and M. Lloyd Hopwood. Plasma oestrogen levels in rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 5(2):227–230, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Todd:1973:FRF

[185] P. R. Todd. First record of the freshwater eel Anguilla dieffenbachii Gray to be caught at sea. Journal of Fish Biology, 5(2):231–232, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:1973:WSP

[186] P. J. Miller, A. L. Rice, and A. D. F. Johnstone. A western Scottish population of the leopard-spotted goby, Thorogobius ephippiatus (Lowe) (Teleostei: Gobioidea). Journal of Fish Biology, 5(2):233–239, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McQueen:1973:SSP

[187] A. McQueen, K. Mackenzie, Ronald J. Roberts, and Hamish Young. Studies on the skin of plaice (Pleuronectes platessa L.) III. The effect of temperature on the inflammatory response to the metacercariae of Cryptocotyle lingua (Creplin, 1825) (Digenea: Heterophyidae). Journal of Fish Biology, 5(2):241–247, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1973:SSP

[188] I. A. Johnston and G. Goldspink. A study of the swimming performance of the Crucian carp Carassius carassius (L.) in relation to the effects of exercise and recovery on biochemical changes in the myotomal muscles and liver. Journal of Fish Biology, 5(2):249–260, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1973:IRD

[189] J. E. Harris. The immune responses of dace Leuciscus leuciscm (L.) to injected antigenic materials. Journal of Fish Biology, 5(2):261–276, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1973:OFR

[190] P. C. Hunt and K. O’Hara. Overwinter feeding in rainbow trout. Journal of Fish Biology, 5(2):277–280, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:BRb

[191] Anonymous. Book reviews. Journal of Fish Biology, 5(2):281–284, March 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warren:1973:ENP

[192] E. W. Warren. The establishment of a ‘normal’ population and its behavioural maintenance in the guppy — Poecilia reticulata (Peters). Journal of Fish Biology, 5(3):285–304, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sylvester:1973:NUL

[193] J. R. Sylvester. A note on the upper lethal temperature of juvenile Haemulon flavolineatum from the Virgin Islands. Journal of Fish Biology, 5(3): 305–307, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingham:1973:IHR

[194] L. Ingham and C. Arme. Intestinal helminths in rainbow trout, Salmo gairdneri (Richardson): Absence of effect on nutrient absorption and fish growth. Journal of Fish Biology, 5(3):309–313, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rio:1973:CEG

[195] G. J. Rio, F. J. Magnavita, J. A. Rubin, and Wm H. Beckert. Characteristics of an established goldfish Carassim auratus (L.) cell line. Journal of Fish Biology, 5(3):315–321, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riley:1973:MGP

[196] J. D. Riley. Movements of 0-group plaice Pleuronectes platessa L. as shown by latex tagging. Journal of Fish Biology, 5(3):323–343, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1973:HAM

[197] J. E. Harris, Aileen Watson, and S. Hunt. Histochemical analysis of mucous cells in the epidermis of brown trout Salmo trutta L. Journal of Fish Biology, 5(3):345–351, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:1973:SPT

[198] P. J. Miller. The species of Pseudaphya (Teleostei: Gobiidae) and the evolution of aphyiine gobies. Journal of Fish Biology, 5(3):353–365, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1973:EYT

[199] A. Jones. The ecology of young turbot, Scophthalmus maximus (L.), at Borth, Cardiganshire, Wales. Journal of Fish Biology, 5(3):367–383, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mines:1973:IMF

[200] Ron S. Mines and Dan T. Spira. Ichthyophthirius multifiliis (Fouquet) in the mirror carp, Cyprinus carpio L. I. Course of infection. Journal of Fish Biology, 5(3):385–392, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deCiechomski:1973:SEA

[201] Janina D. de Ciechomski. The size of the eggs of the Argentine anchovy, Engraulis anchoita (Hubbs & Marini) in relation to the season of the year and to the area of spawning. Journal of Fish Biology, 5(3):393–398, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hicks:1973:MPS

[202] Franklin J. Hicks and William Threlfall. Metazoan parasites of salmonids and coregonids from coastal Labrador. Journal of Fish Biology, 5(3):399– 415, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:BRc

[203] Anonymous. Book reviews. Journal of Fish Biology, 5(3):417–418, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:SIC

[204] Anonymous. Seventh International Conference on Water Pollution Research, Paris 9–13 September 1974. Journal of Fish Biology, 5(3):419, May 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cech:1973:RRS

[205] Joseph J. Cech, Jr. and Donald E. Wohlschlag. Respiratory responses of the striped mullet, Mugil cephalus (L.) to hypoxic conditions. Journal of Fish Biology, 5(4):421–428, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Everett:1973:RTS

[206] G. V. Everett. The rainbow trout Salmo gairdneri (Rich.) fishery of Lake Titicaca. Journal of Fish Biology, 5(4):429–440, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1973:EDN

[207] D. J. Edwards. The effects of drugs and nerve-section on the rate of passage of food through the gut of the plaice, Pleuronectes platessa (L.). Journal of Fish Biology, 5(4):441–446, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cannon:1973:DIH

[208] L. R. G. Cannon. Diet and intestinal helminths in a population of perch, Perca flavescens. Journal of Fish Biology, 5(4):447–457, July 1973. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuo:1973:PRD

[209] Ching-Ming Kuo, Ziad H. Shehadeh, and Kathy K. Milken. A preliminary report on the development, growth and survival of laboratory reared larvae of the grey mullet, Mugil cephalus L. Journal of Fish Biology, 5(4):459– 470, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shehadeh:1973:ISG

[210] Ziad H. Shehadeh, Ching-Ming Kuo, and Kathy K. Milisen. Induced spawning of grey mullet Mugil cephalus L. with fractionated salmon pituitary extract. Journal of Fish Biology, 5(4):471–478, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shehadeh:1973:EEH

[211] Ziad H. Shehadeh, William D. Madden, and Thomas P. Dohl. The effect of exogenous hormone treatment on spermiation and vitellogenesis in the grey mullet, Mugil cephalus L. Journal of Fish Biology, 5(4):479–487, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shehadeh:1973:VVM

[212] Ziad H. Shehadeh, Ching-Ming Kuo, and Kathy K. Milisen. Validation of an in vivo method for monitoring ovarian development in the grey mullet (Mugil cephalus L.). Journal of Fish Biology, 5(4):489–496, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1973:HSTa

[213] Ronald J. Roberts, A. MacQueen, W. M. Shearer, and H. Young. The histopathology of salmon tagging: I. The tagging lesion in newly tagged parr. Journal of Fish Biology, 5(4):497–503, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dartnall:1973:PNS

[214] H. J. G. Dartnall. Parasites of the nine-spined stickleback Pungitius pungitius (L). Journal of Fish Biology, 5(4):505–509, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:1973:AFE

[215] F. Douglas Martin and Rollin C. Richmond. An analysis of five enzymegene loci in four etheostomid species (Percidae: Pices) in an area of possible introgression. Journal of Fish Biology, 5(4):511–517, July 1973. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nicol:1973:TLE

[216] J. A. C. Nicol and E. S. Zyznar. The tapetum lucidum in the eye of the big-eye Priacanthus arenatus Cuvier. Journal of Fish Biology, 5(4): 519–522, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Toit:1973:HLU

[217] P. J. Du Toit, J. Hattingh, and J. C. Schabort. The haemoglobins of Labeo umbratus (Smith) and Labeo capensis (Smith). Journal of Fish Biology, 5 (4):523–526, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hines:1973:IMC

[218] Ron S. Hines and Dan T. Spira. Ichthyophthiriasis in the mirror carp: II. Leukocyte response. Journal of Fish Biology, 5(4):527–534, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1973:AIC

[219] J. E. Harris. The apparent inability of cyprinid fish to produce skinsensitizing antibody. Journal of Fish Biology, 5(4):535–540, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marcy:1973:WSS

[220] B. C. Marcy, Jr. and R. C. Galvin. Winter-spring sport fishery in the heated discharge of a nuclear power plant. Journal of Fish Biology, 5(4): 541–547, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:1973:NIM

[221] T. J. Pitcher and P. D. M. Macdonald. A numerical integration method for fish population fecundity. Journal of Fish Biology, 5(4):549–553, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagenal:1973:DBB

[222] T. B. Bagenal, the Late F. J. H. Mackereth, and J. Heron. The distinction between brown trout and sea trout by the strontium content of their scales. Journal of Fish Biology, 5(4):555–557, July 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shotter:1973:CPFb

[223] R. A. Shotter. Changes in the parasite fauna of whiting Odontogadus merlangus L. with age and sex of host, season, and from different areas in the vicinity of the Isle of Man. Journal of Fish Biology, 5(5):559–573, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otto:1973:TTM

[224] Robert G. Otto. Temperature tolerance of the mosquitofish, Gambmia affinis (Baird and Girard). Journal of Fish Biology, 5(5):575–585, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cobb:1973:SRS

[225] J. L. S. Cobb, N. C. Fox, and R. M. Santer. A specific ringer solution for the plaice (Pleuronectes platessa L.). Journal of Fish Biology, 5(5): 587–591, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:1973:TST

[226] Stephen J. M. Blaber. Temperature and salinity tolerance of juvenile Rhabdosargus holubi [Steindachner (Teleostei: Sparidae)]. Journal of Fish Biology, 5(5):593–598, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1973:ASC

[227] D. J. Edwards. Aquarium studies on the consumption of small animals by O-group grass carp, Ctenopharyngodon idella (Val.). Journal of Fish Biology, 5(5):599–605, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthiessen:1973:EDZ

[228] P. Matthiessen and A. E. Brafield. The effects of dissolved zinc on the gills of the stickleback Gasterosteus aculeatus (L.). Journal of Fish Biology, 5 (5):607–613, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1973:HSTb

[229] Ronald J. Roberts, A. McQueen, W. M. Shearer, and H. Young. The histopathology of salmon tagging: II. The chronic tagging lesion in returning adult fish. Journal of Fish Biology, 5(5):615–619, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1973:HSTc

[230] Ronald J. Roberts, A. McQueen, W. M. Shearer, and H. Young. The histopathology of salmon tagging: III. Secondary infections associated with tagging. Journal of Fish Biology, 5(5):621–623, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adron:1973:SQS

[231] J. W. Adron, P. T. Grant, and C. B. Cowey. A system for the quantitative study of the learning capacity of rainbow trout and its application to the study of food preferences and behaviour. Journal of Fish Biology, 5(5): 625–636, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:BRd

[232] Anonymous. Book reviews. Journal of Fish Biology, 5(5):637–639, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:EIFa

[233] Anonymous. European Inland Fisheries Advisory Commission: Symposium on methodology for the survey, monitoring and appraisal of fishery resources in lakes and large rivers. Journal of Fish Biology, 5(5):640, September 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cocoros:1973:MCF

[234] Glenn Cocoros, Phyllis H. Cahn, and William Siler. Mercury concentrations in fish, plankton and water from three Western Atlantic estuaries. Journal of Fish Biology, 5(6):641–647, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1973:NEL

[235] B. Stott and D. G. Cross. A note on the effect of lowered temperatures on the survival of eggs and fry of the grass carp Ctenopharyngodon idella (Valenciennes). Journal of Fish Biology, 5(6):649–658, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Went:1973:MSS

[236] A. E. J. Went. Movements of salmon Salmo salar (L.) to and from Irish waters. Journal of Fish Biology, 5(6):659–671, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe:1973:CPB

[237] D. R. Lowe, F. W. H. Beamish, and I. C. Potter. Changes in the proximate body composition of the landlocked sea lamprey Petromyzon marinus (L.) during larval life and metamorphosis. Journal of Fish Biology, 5(6):673– 682, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1973:FTS

[238] R. J. Wootton. Fecundity of the three-spined stickleback, Gasterosteus aculeatus (L.). Journal of Fish Biology, 5(6):683–688, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deSilva:1973:ARB

[239] S. S. de Silva. Aspects of the reproductive biology of the sprat, Sprattus sprattus (L.) in inshore waters of the west coast of Scotland. Journal of Fish Biology, 5(6):689–705, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1973:OAG

[240] R. H. K. Mann. Observations on the age, growth, reproduction and food of the roach Rutilus rutilus (L.) in two rivers in southern England. Journal of Fish Biology, 5(6):707–736, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warren:1973:MRH

[241] E. W. Warren. Modification of the response to high density conditions in the guppy, Poecilia reticulata (Peters). Journal of Fish Biology, 5(6): 737–752, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warren:1973:ERD

[242] E. W. Warren. The effects of relative density upon some aspects of the behaviour of the guppy — Poecilia reticulata (Peters). Journal of Fish Biology, 5(6):753–765, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1973:EHV

[243] Peter C. Blaxhall. Error in haematocrit value produced by inadequate concentration of ethylenediamine tetra-acetate. Journal of Fish Biology, 5(6):767–769, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1973:RHM

[244] P. C. Blaxhall and K. W. Daisley. Routine haematological methods for use with fish blood. Journal of Fish Biology, 5(6):771–781, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1973:ATM

[245] M. S. Caulton and B. J. Hill. The ability of Tilapia mossambica (Peters) to enter deep water. Journal of Fish Biology, 5(6):783–788, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1973:CCI

[246] D. G. Cross and P. A. Hursey. Chloramine-T for the control of Ichthyophthirim multifiliis (Fouquet). Journal of Fish Biology, 5(6):789–798, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:BRe

[247] Anonymous. Book reviews. Journal of Fish Biology, 5(6):799–802, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:EIFb

[248] Anonymous. European Inland Fisheries Advisory Commission: Symposium on methodology for the survey, monitoring and appraisal of fishery resources in lakes and large rivers. Journal of Fish Biology, 5(6):803, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1973:SA

[249] Anonymous. Special announcement. Journal of Fish Biology, 5(6):804, November 1973. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezzat:1974:SBC

[250] A. A. Ezzat, M. B. Shabana, and A. M. Farghaly. Studies on the blood characteristics of Tilapia zilli (Gervais) I. Blood cells. Journal of Fish Biology, 6(1):1–12, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gupta:1974:ORB

[251] S. Gupta. Observations on the reproductive biology of Mastacembelus armatus (Lacèpede). Journal of Fish Biology, 6(1):13–21, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanderWaal:1974:OBH

[252] B. C. W. van der Waal. Observations on the breeding habits of Clarias gariepinus (Burchell). Journal of Fish Biology, 6(1):23–27, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1974:DCP

[253] W. J. R. Lanzing and C. C. Bower. Development of colour patterns in relation to behaviour in Tilapia mossambica (Peters). Journal of Fish Biology, 6(1):29–41, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sylvester:1974:PSL

[254] J. R. Sylvester. A preliminary study of the length composition, distribution and relative abundance of three species of deepwater snappers from the Virgin Islands. Journal of Fish Biology, 6(1):43–49, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balsano:1974:MEA

[255] J. S. Balsano and E. M. Rasch. Microspectrophotometric and enzymatic analyses of fish plasma proteins electrophoretically separated in thin polyacrylamide gels. Journal of Fish Biology, 6(1):51–59, January 1974. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hill:1974:PCD

[256] B. J. Hill and M. S. Caulton. A pressure chamber for determination of depth tolerance in teleosts. Journal of Fish Biology, 6(1):61–65, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paperna:1974:HDP

[257] I. Paperna. Hosts, distribution and pathology of infections with larvae of Eustrongylides (Dioctophymidae, Nematoda) in fishes from East African lakes. Journal of Fish Biology, 6(1):67–76, January 1974. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanderson:1974:NOP

[258] J. M. Sanderson. A new occurrence of the parasitical copepod Lernaea cyprinacea (L.). Journal of Fish Biology, 6(1):77–78, January 1974. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:1974:FGA

[259] J. W. Moore and I. A. Moore. Food and growth of Arctic char, Salvelinus alpinus (L.), in the Cumberland Sound area of Baffin Island. Journal of Fish Biology, 6(1):79–92, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brocksen:1974:PII

[260] Robert W. Brocksen and Jan P. Bugge. Preliminary investigations on the influence of temperature on food assimilation by rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 6(1):93–97, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Webb:1974:FCD

[261] C. J. Webb. Fish chromosomes: a display by scanning electron microscopy. Journal of Fish Biology, 6(1):99–100, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1974:BRa

[262] Anonymous. Book reviews. Journal of Fish Biology, 6(1):101–102, January 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacKenzie:1974:IDA

[263] K. MacKenzie. Immature digeneans from the alimentary tract of larval and juvenile pelagic stages of haddock, Melanogrammus aeglefinus (L.). Journal of Fish Biology, 6(2):103–106, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubbock:1974:NMP

[264] Roger Lubbock and Barry Goldman. A new magenta Pseudochromi (Teleostei: Pseudochromidae) from the Pacific. Journal of Fish Biology, 6(2):107–110, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1974:DMC

[265] Alan D. Pickering. The distribution of mucous cells in the epidermis of the brown trout Salmo trutta (L.) and the char Salvelinus alpinus (L.). Journal of Fish Biology, 6(2):111–118, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fishelson:1974:BEP

[266] L. Fishelson, D. Popper, and A. Avidor. Biosociology and ecology of pomacentrid fishes around the Sinai Peninsula (northern Red Sea). Journal of Fish Biology, 6(2):119–133, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1974:ENB

[267] J. E. Thorpe. Estimation of the number of brown trout Salmo trutta (L.) in Loch Leven, Kinross, Scotland. Journal of Fish Biology, 6(2):135– 152, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1974:MBT

[268] J. E. Thorpe. The movements of brown trout, Salmo trutta (L.) in Loch Leven, Kinross, Scotland. Journal of Fish Biology, 6(2):153–180, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1974:LIB

[269] Martyn T. Harris and Alwyne Wheeler. Ligula infestation of bleak Alburnus alburnus (L.) in the tidal Thames. Journal of Fish Biology, 6(2): 181–188, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hines:1974:IMCa

[270] Ron S. Hines and Dan T. Spira. Ichthyophthiriasis in the mirror carp Cyprinus carpio (L.) III. Pathology. Journal of Fish Biology, 6(2):189– 196, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Snieszko:1974:EES

[271] S. F. Snieszko. The effects of environmental stress on outbreaks of infectious diseases of fishes. Journal of Fish Biology, 6(2):197–208, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McDowall:1974:SDS

[272] R. M. McDowall. Specialization in the dentition of the southern graylings — genus Prototroctes (Galaxioidei: Prototroctidae). Journal of Fish Biology, 6(2):209–213, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rajagopal:1974:RMU

[273] Pokkavil K. Rajagopal and Robert H. Kramer. Respiratory metabolism of Utah chub, Gila atraria (Girard) and speckled dace, Rhinichthys osculus (Girard). Journal of Fish Biology, 6(2):215–222, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1974:BRb

[274] Anonymous. Book review. Journal of Fish Biology, 6(2):223, March 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hawkins:1974:TCG

[275] A. D. Hawkins, D. N. MacLennan, G. G. Urquhart, and C. Robb. Tracking cod Gadus morhua L. in a Scottish sea loch. Journal of Fish Biology, 6(3): 225–236, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1974:OAG

[276] R. H. K. Mann. Observations on the age, growth, reproduction and food of the dace, Leuciscus leuciscus (L.), in two rivers in southern England.

Journal of Fish Biology, 6(3):237–253, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1974:PSBa

[277] P. C. Hunt and J. W. Jones. A population study of Barbus barbus (L.) in the River Severn, England: I. Densities. Journal of Fish Biology, 6(3): 255–267, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1974:PSBb

[278] P. C. Hunt and J. W. Jones. A population study of Barbus barbus (L.) in the River Severn, England: II. Movements. Journal of Fish Biology, 6(3): 269–278, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abel:1974:TSD

[279] P. D. Abel. Toxicity of synthetic detergents to fish and aquatic invertebrates. Journal of Fish Biology, 6(3):279–298, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burbidge:1974:AGL

[280] Richard G. Burbidge, Mario C. Carrasco, and Peter A. Brown. Age, growth, length–weight relationship, sex ratio and food habits of the Argentine pejerrey, Basilichthys bonariensis (Cuv. and Val.), from Lake Peñuelas, Valparaíso, Chile. Journal of Fish Biology, 6(3):299–305, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1974:SMS

[281] W. J. R. Lanzing and D. R. Higginbotham. Scanning microscopy of surface structures of Tilapia mossambica (Peters) scales. Journal of Fish Biology, 6(3):307–310, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:1974:GRC

[282] A. Ferguson. The genetic relationships of the coregonid fishes of Britain and Ireland indicated by electrophoretic analysis of tissue proteins. Journal of Fish Biology, 6(3):311–315, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beukema:1974:ETB

[283] J. J. Beukema and G. J. de Vos. Experimental tests of a basic assumption of the capture-recapture method in pond populations of carp Cyprinus carpio L. Journal of Fish Biology, 6(3):317–329, May 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Viswanathan:1974:SCH

[284] N. Viswanathan and Bangalore I. Sundararaj. Seasonal changes in the hypothalamo-hypophyseal-ovarian system in the catfish, Heteropneustes fossilis (Bloch). Journal of Fish Biology, 6(3):331–340, May 1974. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1974:SPC

[285] W. J. R. Lanzing. Sound production in the cichlid Tilapia mossambica Peters. Journal of Fish Biology, 6(4):341–347, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1974:FFH

[286] D. S. C. Lewis. The food and feeding habits of Hydrocynus forskahlii Cuvier and Hydrocynus brevis Günther in Lake Kainji, Nigeria. Journal of Fish Biology, 6(4):349–363, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hines:1974:IMCb

[287] Ron S. Hines and Dan T. Spira. Ichthyophthiriasis in the mirror carp Cyprinus carpio (L.) IV. Physiological dysfunction. Journal of Fish Biology, 6(4):365–371, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hines:1974:IMCc

[288] Ron S. Hines and Dan T. Spira. Ichthyophthiriasis in the mirror carp Cyprinus carpio (L.) V. Acquired immunity. Journal of Fish Biology, 6 (4):373–378, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swift:1974:CUF

[289] D. J. Swift and R. Lloyd. Changes in urine flow rate and haematocrit value of rainbow trout Salmo gairdneri (Richardson) exposed to hypoxia. Journal of Fish Biology, 6(4):379–387, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boyd:1974:CLP

[290] John F. Boyd and Richard C. Simmonds. Continuous laboratory production of fertile fundulusheteroclitus (Walbaum) eggs lacking chorionic fibrils. Journal of Fish Biology, 6(4):389–394, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:1974:FPI

[291] R. Williams and M. F. Harcup. The fish populations of an industrial river in South Wales. Journal of Fish Biology, 6(4):395–414, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macer:1974:RBH

[292] C. T. Macer. The reproductive biology of the horse mackerel Trachurus trachurs (L.) in the North Sea and English Channel. Journal of Fish Biology, 6(4):415–438, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1974:PPL

[293] J. Hattingh. The plasma proteins of Labeo umbratus (Smith) and Labeo capensis (Smith). Journal of Fish Biology, 6(4):439–446, July 1974. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1974:RCM

[294] D. B. C. Scott. The reproductive cycle of Mormyrus kannume Forssk. (Osteoglossomorpha, Mormyriformes) in Lake Victoria, Uganda. Journal of Fish Biology, 6(4):447–454, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:1974:PSG

[295] Stephen J. M. Blaber. The population structure and growth of juvenile Rhabdosargus holubi (Steindachner) (Teleostei: Sparidae) in a closed estuary. Journal of Fish Biology, 6(4):455–460, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1974:EHS

[296] J. Hattingh and P. J. Du Toit. Electrophoresis of haemoglobin in single erythrocytes of Labeo umbratus (Smith) and Labeo capensis (Smith). Journal of Fish Biology, 6(4):461–463, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lockwood:1974:SDM

[297] Stephen J. Lockwood. The settlement, distribution and movements of 0group plaice Pleuronectes platessa (L.) in Filey Bay, Yorkshire. Journal of Fish Biology, 6(4):465–477, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:1974:OCH

[298] Julio E. Perez and Norman Maclean. Ontogenetic changes in haemoglobins in roach, Rutilus rutilus (L.) and rudd, Scardinius erythrophthalmus (L.). Journal of Fish Biology, 6(4):479–482, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chilvers:1974:FBD

[299] R. M. Chilvers and J. M. Gee. The food of Bagrus docmac (Forssk.) (Pisces: Siluriformes) and its relationship with Haplochromis Hilgendorf (Pisces: Cichlidae), in Lake Victoria, East Africa. Journal of Fish Biology, 6(4):483–505, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OBrien:1974:ULF

[300] D. J. O’Brien. Use of lesion filtrates for the transmission of UDN (ulcerative dermal necrosis) in salmonids. Journal of Fish Biology, 6(4):507–511, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vonWesternhagen:1974:ONS

[301] H. von Westernhagen. Observations on the natural spawning of Alectis indicus (Rüppell) and Caranx ignobilis (Forssk.) (Carangidae). Journal of Fish Biology, 6(4):513–516, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardiner:1974:EMD

[302] W. Ross Gardiner. An electrophoretic method for distinguishing the young fry of salmon Salmo salar (L.) from those of trout Salmo trutta (L.). Journal of Fish Biology, 6(4):517–519, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1974:ODS

[303] P. M. Hine and C. R. Kennedy. Observations on the distribution, specificity and pathogenicity of the acanthocephalan Pomphorhynchus laevis (Müller). Journal of Fish Biology, 6(4):521–535, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe:1974:OCA

[304] G. A. Lowe. The occurrence of ciliary aggregations on the olfactory epithelium of two species of gadoid fish. Journal of Fish Biology, 6(4):537–539, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1974:BRc

[305] Anonymous. Book reviews. Journal of Fish Biology, 6(4):541–542, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1974:Aa

[306] Anonymous. Announcments. Journal of Fish Biology, 6(4):543, July 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Natochin:1974:IWS

[307] Yu. V. Natochin and E. A. Lavrova. The influence of water salinity and stage in life history on ion concentration of fish blood serum. Journal of Fish Biology, 6(5):545–555, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirose:1974:ECH

[308] Keiji Hirose and Rikizo Ishida. Effects of cortisol and human chorionic gonadotrophin (HCG) on ovulation in ayu Plecoglossus altivelis (Temminck & Schlegel) with special respect to water and ion balance. Journal of Fish Biology, 6(5):557–564, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winslade:1974:BSLa

[309] P. Winslade. Behavioural studies on the lesser sandeel Ammodytes marinus (Raitt) I. The effect of food availability on activity and the role of olfaction in food detection. Journal of Fish Biology, 6(5):565–576, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winslade:1974:BSLb

[310] P. Winslade. Behavioural studies on the lesser sandeel Ammodytes marinus (Raitt) II. The effect of light intensity on activity. Journal of Fish Biology, 6(5):577–586, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winslade:1974:BSLc

[311] P. Winslade. Behavioural studies on the lesser sandeel Ammodytes marinus (Raitt) III. The effect of temperature on activity and the environmental control of the annual cycle of activity. Journal of Fish Biology, 6 (5):587–599, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eisler:1974:RES

[312] Ronald Eisler. Radiocadmium exchange with seawater by Fundulus heteroclitus (L.) (Pisces: Cyprinodontidae). Journal of Fish Biology, 6(5): 601–612, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1974:CBI

[313] C. R. Kennedy. A checklist of British and Irish freshwater fish parasites with notes on their distribution. Journal of Fish Biology, 6(5):613–644, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:1974:LSF

[314] S. S. De Silva and F. Balbontin. Laboratory studies on food intake, growth and food conversion of young herring, Clupea harengus (L.). Journal of Fish Biology, 6(5):645–658, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1974:FNE

[315] T. F. Cross and M. C. Lyes. A first north-eastern Atlantic record of the bluefish Pomatomus saltatrix (L. 1758), (fam. Pomatomidae). Journal of Fish Biology, 6(5):659–660, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plumb:1974:HEM

[316] J. A. Plumb, J. L. Gaines, E. C. Mora, and G. G. Bradley. Histophathology and electron microscopy of channel catfish virus in infected channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 6(5): 661–664, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1974:PBA

[317] P. M. Hine and C. R. Kennedy. The population biology of the acanthocephalan Pomphorhynchus laevis (Müller) in the River Avon. Journal of Fish Biology, 6(5):665–679, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucke:1974:VAC

[318] David Bucke. Vertebral abnormalities in common bream Abramis brama (L.). Journal of Fish Biology, 6(5):681–682, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1974:BRd

[319] Anonymous. Book reviews. Journal of Fish Biology, 6(5):683–685, September 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jolly:1974:O

[320] Donald Jolly. Obituary. Journal of Fish Biology, 6(6):687, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cervinka:1974:DGD

[321] S. Červinka, J. Vítovec, J. Lom, J. Hoška, and F. Kubů. Dermocystidiosis — a gill disease of the carp due to Dermocystidium cyprini n. sp. Journal of Fish Biology, 6(6):689–699, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruton:1974:GTM

[322] M. N. Bruton and B. R. Allanson. The growth of Tilapia mossambica Peters (Pisces: Cichlidae) in Lake Sibaya, South Africa. Journal of Fish Biology, 6(6):701–715, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Molnar:1974:PFL

[323] Kalman Molnar, George Hanek, and C. H. Fernando. Parasites of fishes from Laurel Creek, Ontario. Journal of Fish Biology, 6(6):717–728, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellawell:1974:UMB

[324] J. M. Hellawell, H. Leatham, and G. I. Williams. The upstream migratory behaviour of salmonids in the River Frome, Dorset. Journal of Fish Biology, 6(6):729–744, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulata:1974:RGE

[325] G. Hulata, R. Moav, and G. Wohlfarth. The relationship of gonad and egg size to weight and age in the European and Chinese races of the common carp Cyprinus carpio L. Journal of Fish Biology, 6(6):745–758, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walker:1974:EFH

[326] M. H. Walker and C. G. Alexander. Encapsulated fish hooks from the stomachs of Plectropomus maculatus (Bloch) (Serranidae). Journal of Fish Biology, 6(6):759–761, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soivio:1974:CHV

[327] Antti Soivio, Kai Westman, and Keijo Nyholm. Changes in haematocrit values in blood samples treated with and without oxygen: a comparative study with four salmonid species. Journal of Fish Biology, 6(6):763–769, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russell:1974:LWP

[328] P. H. Russell. Lymphocystis in wild plaice Pleuronectes platessa (L.), and flounder, Platichthys flesus (L.), in British coastal waters: a histopathological and serological study. Journal of Fish Biology, 6(6):771–778, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stickney:1974:OCA

[329] Robert R. Stickney and Sandra E. Shumway. Occurrence of cellulase activity in the stomachs of fishes. Journal of Fish Biology, 6(6):779–790, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sylvester:1974:TRJ

[330] J. R. Sylvester. Thermal response of juvenile Hawaiian mullet Mugil cephalus (L.) to acclimation time and fluctuating low temperatures. Journal of Fish Biology, 6(6):791–796, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:1974:OJR

[331] Stephen J. M. Blaber. Osmoregulation in juvenile Rhabdosargus holubi (Steindachner) (Teleostei: Sparidae). Journal of Fish Biology, 6(6):797– 800, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okera:1974:MCG

[332] W. Okera. Morphometrics, ‘condition’ and gonad development of the East African Sardinella gibbosa (Bleeker) and Sardinella albella (Valenciennes). Journal of Fish Biology, 6(6):801–812, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1974:Ab

[333] Anonymous. Announcement. Journal of Fish Biology, 6(6):813, November 1974. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staples:1975:PBUa

[334] D. J. Staples. Production biology of the upland bully Philypnodon breviceps Stokell in a small New Zealand lake: I. Life history, food, feeding and activity rhythms. Journal of Fish Biology, 7(1):1–24, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staples:1975:PBUb

[335] D. J. Staples. Production biology of the upland bully Philypnodon breviceps Stokell in a small New Zealand lake: II. Population dynamics. Journal of Fish Biology, 7(1):25–45, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staples:1975:PBUc

[336] D. J. Staples. Production biology of the upland bully Philypnodon breviceps Stokell in a small New Zealand lake: III. Production, food consumption and efficiency of food utilization. Journal of Fish Biology, 7(1):47–69, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gupta:1975:SOB

[337] S. Gupta. Some observations on the biology of Cirrhinus reba (Cuvier). Journal of Fish Biology, 7(1):71–76, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mathews:1975:SOEa

[338] C. P. Mathews. Some observations on the ecology of Galeus piperatus Springer and Wagner, a little known shark endemic to the Northern Gulf of California. Journal of Fish Biology, 7(1):77–82, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mathews:1975:SOEb

[339] C. P. Mathews. Some observations on the ecology and the population dynamics of Merluccius angustimanus in the South Gulf of California. Journal of Fish Biology, 7(1):83–94, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1975:PSF

[340] A. Neville Jones. A preliminary study of fish segregation in salmon spawning streams. Journal of Fish Biology, 7(1):95–104, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warren:1975:IDR

[341] E. W. Warren and Shirley Callaghan. Individual differences in response to an open field test by the guppy — Poecilia reticulata (Peters). Journal of Fish Biology, 7(1):105–113, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Houde:1975:ESD

[342] Edward D. Houde. Effects of stocking density and food density on survival, growth and yield of laboratory-reared larvae of sea bream Archosargus rhomboidalis (L.) (Sparidae). Journal of Fish Biology, 7(1):115–127, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flood:1975:SAI

[343] Mary T. Flood, Ross F. Nigrelli, and Joseph F. Gennaro, Jr. Some aspects of the infrastructure of the ‘Stäbchendrüscnzellen’, a peculiar cell associated with the endothelium of the bulbus arteriosus and with other fish tissues. Journal of Fish Biology, 7(1):129–138, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1975:BRa

[344] Anonymous. Book reviews. Journal of Fish Biology, 7(1):139–142, January 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:1975:RBA

[345] J. W. Moore. Reproductive biology of anadromous Arctic char, Salvelinus alpinus (L.)? in the Cumberland Sound area of Baffin Island. Journal of Fish Biology, 7(2):143–151, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chubb:1975:FTF

[346] James C. Chubb, J. W. Jones, and J. W. Banks. Film and television in fishery research at Llyn Tegid (Bala Lake), Wales. Journal of Fish Biology, 7(2):153–157, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Patterson:1975:HSL

[347] S. Patterson, I. A. Johnston, and G. Goldspink. A histochemical study of the lateral muscles of five teleost species. Journal of Fish Biology, 7(2): 159–166, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laird:1975:FBJ

[348] Lindsay M. Laird, Ronald J. Roberts, W. M. Shearer, and J. F. McArdle. Freeze branding of juvenile salmon. Journal of Fish Biology, 7(2):167– 171, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:1975:CET

[349] C. D. Anderson and R. J. Roberts. A comparison of the effects of temperature on wound healing in a tropical and a temperate teleost. Journal of Fish Biology, 7(2):173–182, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1975:RLK

[350] Alwyne Wheeler, R. W. Blacker, and S. F. Pirie. Rare and little-known fishes in British seas in 1970 and 1971. Journal of Fish Biology, 7(2):183– 201, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duthu:1975:ASS

[351] Gwen S. Duthu and Ronald H. Kilgen. Aquarium studies on the selectivity of 16 aquatic plants as food by fingerling hybrids of the cross between Ctenopharyngodon idella [male sign] and Cyprinus carpio [female sign]. Journal of Fish Biology, 7(2):203–208, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dobbs:1975:SEM

[352] Gary H. Dobbs. Scanning electron microscopy of intraovarian embryos of the viviparous teleost, Micrometrus minimus (Gibbons), (Perciformes: Embiotocidae). Journal of Fish Biology, 7(2):209–214, March 1975. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:1975:SBP

[353] D. H. McCarthy, J. P. Stevenson, and M. S. Roberts. Some blood parameters of the rainbow trout (Salmo gairdneri Richardson) II. The Shasta variety. Journal of Fish Biology, 7(2):215–219, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1975:ETA

[354] M. S. Caulton and B. J. Hill. The effect of temperature on the ability of Tilapia mossambica Peters to enter deep water. Journal of Fish Biology, 7(2):221–226, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanderWaal:1975:AGS

[355] B. C. W. van der Waal and H. J. Schoonbee. Age and growth studies of Clarias gariepinus (Burchell) (Clariidae) in the Transvaal, South Africa. Journal of Fish Biology, 7(2):227–233, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ikusemiju:1975:AEL

[356] Kola Ikusemiju. Aspects of the ecology and life history of the sculpin, coitus aleuticm(Gilbert), in Lake Washington. Journal of Fish Biology, 7(2):235–245, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacKenzie:1975:SABa

[357] K. MacKenzie and Janet M. Liversidge. Some aspects of the biology of the cercaria and metacercaria of Stephanostomum baccatum (Nicoll, 1907) Manter, 1934 (Digenea: Acanthocolpidae). Journal of Fish Biology, 7(2): 247–256, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sylvester:1975:CTM

[358] J. R. Sylvester. Critical thermal maxima of three species of Hawaiian estuarine fish: a comparative study. Journal of Fish Biology, 7(2):257– 262, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dick:1975:SOU

[359] Gillian L. Dick. Some observations on the use of MS 222 sandoz with grey mullet (Mugil chelo Cuvier). Journal of Fish Biology, 7(2):263–268, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bootsma:1975:TEP

[360] R. Bootsma, P. de Kinkelin, and Monique Le Berre. Transmission experiments with pike fry (Esox indus L.) rhabdovirus. Journal of Fish Biology, 7(2):269–276, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1975:BRb

[361] Anonymous. Book reviews. Journal of Fish Biology, 7(2):277–281, March 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaughan:1975:SET

[362] Gene E. Vaughan and Daniel W. Coble. Sublethal effects of three ectoparasites on fish. Journal of Fish Biology, 7(3):283–294, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greek-Walker:1975:SRW

[363] M. Greek-Walker and G. A. Pull. A survey of red and white muscle in marine fish. Journal of Fish Biology, 7(3):295–300, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucke:1975:RSE

[364] D. Bucke, D. McCcarthy, and B. Hill. A report of suspected erythrodermatitis in carp in Great Britain. Journal of Fish Biology, 7(3):301–303, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fagerlund:1975:GIS

[365] U. H. M. Fagerlund and J. R. McBride. Growth increments and some flesh and gonad characteristics of juvenile Coho salmon receiving diets supplemented with 17α-methyltestosterone. Journal of Fish Biology, 7 (3):305–314, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1975:FCT

[366] P. C. Blaxhall. Fish chromosome techniques — a review of selected literature. Journal of Fish Biology, 7(3):315–320, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacKenzie:1975:SABb

[367] K. MacKenzie. Some aspects of the biology of the plerocercoid of Gilquinia squali Fabricius 1794 (Cestoda: Trypanorhyncha). Journal of Fish Biology, 7(3):321–327, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stasko:1975:PMA

[368] Aivars B. Stasko. Progress of migrating Atlantic salmon (Salmo salar) along an estuary, observed by ultrasonic tracking. Journal of Fish Biology, 7(3):329–338, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:1975:DMM

[369] J. W. Moore. Distribution, movements, and mortality of anadromous Arctic char, Salvelinus alpinus L., in the Cumberland Sound area of Baffin Island. Journal of Fish Biology, 7(3):339–348, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1975:EEF

[370] D. G. Cross and B. Stott. The effect of electric fishing on the subsequent capture of fish. Journal of Fish Biology, 7(3):349–357, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackenzie:1975:RMD

[371] K. Mackenzie. Renicola metacercariae (Digenea: Renicolidae) in clupeoid fish: new host records. Journal of Fish Biology, 7(3):359–360, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1975:PSB

[372] P. C. Hunt and J. W. Jones. A population study of Barbus barbus L. in the River Severn, England: III. Growth. Journal of Fish Biology, 7(3): 361–376, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Timms:1975:ICG

[373] A. M. Timms. Intraspecific communication in goldfish. Journal of Fish Biology, 7(3):377–389, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stein:1975:SPC

[374] R. A. Stein, J. F. Kitchell, and Borivoj Knežzevic. Selective predation by carp (Cyprinus carpio L.) on benthic molluscs in Skadar Lake, Yugoslavia. Journal of Fish Biology, 7(3):391–399, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eccles:1975:FAG

[375] D. H. Eccles. Fishes of the African Great Lakes as candidates for introduction into large tropical impoundments. Journal of Fish Biology, 7(3): 401–405, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crisp:1975:SOG

[376] D. T. Crisp and S. M. Carrick. Some observations on the growth and length: weight relationship of the South Georgia cod Notothenia rossii marmorata Fischer during the first four years of life. Journal of Fish Biology, 7(3):407–409, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1975:BRc

[377] Anonymous. Book reviews. Journal of Fish Biology, 7(3):411–414, May 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peters:1975:SFI

[378] Gabriele Peters. Seasonal fluctuations in the incidence of epidermal papillomas of the European eel Anguilla anguilla L. Journal of Fish Biology, 7 (4):415–422, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruton:1975:ABT

[379] M. N. Bruton and R. E. Boltt. Aspects of the biology of Tilapia mossambica Peters (Pisces: Cichlidae) in a natural freshwater lake (Lake Sibaya, South Africa). Journal of Fish Biology, 7(4):423–445, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1975:TFF

[380] J. Hattingh, F. Le Roux Fourie, and J. H. J. van Vuren. The transport of freshwater fish. Journal of Fish Biology, 7(4):447–449, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1975:SSMa

[381] I. A. Johnston, P. S. Ward, and G. Goldspink. Studies on the swimming musculature of the rainbow trout I. Fibre types. Journal of Fish Biology, 7(4):451–458, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1975:SSMb

[382] I. A. Johnston. Studies on the swimming musculature of the rainbow trout. II. Muscle metabolism during severe hypoxia. Journal of Fish Biology, 7 (4):459–467, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rumpus:1975:HPB

[383] A. E. Rumpus. The helminth parasites of the bullhead Cottus gobio (L.) and the stone loach Noemacheilus barbatulus (L.) from the River Avon, Hampshire. Journal of Fish Biology, 7(4):469–483, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

LeDrew:1975:BRS

[384] Bevin R. LeDrew and John M. Green. Biology of the radiated shanny Ulvaria subbifurcata Storer in Newfoundland (Pisces: Stichaeidae). Journal of Fish Biology, 7(4):485–495, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rand:1975:IOF

[385] G. Rand, H. Kleerekoper, and J. Matis. Interaction of odour and flow perception and the effects of parathion in the locomotor orientation of the goldfish Carassius auratus L. Journal of Fish Biology, 7(4):497–504, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaudet:1975:EAP

[386] M. Gaudet, J.-G. Racicot, and C. Leray. Enzyme activities of plasma and selected tissues in rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 7(4):505–512, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1975:ACF

[387] Mark S. Caulton. The ability of the cichlid fishes Tilapia rendatti Boulenger, Tilapia sparrmanii A. Smith and Hemihaplochromis (=Pseudocrenilabrus) philander (M. Weber) to enter deep water. Journal of Fish Biology, 7(4):513–517, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valenti:1975:IPT

[388] R. J. Valenti. Induced polyploidy in Tilapia aurea (Steindachner) by means of temperature shock treatment. Journal of Fish Biology, 7(4): 519–528, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe:1975:UOO

[389] G. A. Lowe and N. K. MacLeod. The ultrastructural organization of olfactory epithelium of two species of gadoid fish. Journal of Fish Biology, 7(4):529–532, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1975:ATN

[390] D. J. A. Brown and T. E. Langford. An assessment of a tow net used to sample coarse fish fry in rivers. Journal of Fish Biology, 7(4):533–538, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corbel:1975:IRF

[391] M. J. Corbel. The immune response in fish: a review. Journal of Fish Biology, 7(4):539–563, July 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atherton:1975:EDL

[392] W. D. Atherton. The effect of different levels of dietary fat on the growth of rainbow trout (Salmo gairdneri Richardson). Journal of Fish Biology, 7(5):565–571, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crisp:1975:PFR

[393] D. T. Crisp, R. H. K. Mann, and Jean C. McCormack. The populations of fish in the River Tees system on the Moor House National Nature Reserve, Westmorland. Journal of Fish Biology, 7(5):573–593, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carscadden:1975:LHV

[394] James E. Carscadden and William C. Leggett. Life history variations in populations of American shad, Alosa sapidissima (Wilson), spawning in tributaries of the St John River, New Brunswick. Journal of Fish Biology, 7(5):595–609, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McVicar:1975:IPP

[395] A. H. McVicar. Infection of plaice Pleuronectes platessa L. with Glugea (Nosema) stephani (Hagenmüller 1899) (Protozoa: Microsporidia) in a fish farm and under experimental conditions. Journal of Fish Biology, 7 (5):611–619, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sylvester:1975:SOT

[396] J. R. Sylvester, C. E. Nash, and C. R. Emberson. Salinity and oxygen tolerances of eggs and larvae of Hawaiian striped mullet, Mugil cephalus L. Journal of Fish Biology, 7(5):621–629, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeLGSolbe:1975:STZ

[397] J. F. De L.G.Solbé and Valerie A. Flook. Studies on the toxicity of zinc sulphate and of cadmium sulphate to stone loach Noemacheilus barbatulus (L.) in hard water. Journal of Fish Biology, 7(5):631–637, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1975:TAS

[398] Alwyne Wheeler and James Dunne. Tripterygion atlanticus sp. nov. (Teleostei–Tripterygiidae) the first record of a tripterygiid fish in North-Western Europe. Journal of Fish Biology, 7(5):639–649, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunn:1975:ENH

[399] I. G. Dunn. Ecological notes on the Haplochromis (Pisces: Cichlidae) species-flock of Lake George, Uganda (East Africa). Journal of Fish Biology, 7(5):651–666, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strommen:1975:CSP

[400] Carol A. Strommen, Ellen M. Rasch, and Joseph S. Balsano. Cytogenetic studies of Poecilia (Pisces) V. Cytophotometric evidence for the production of fertile offspring by triploids related to Poecilia formosa. Journal of Fish Biology, 7(5):667–676, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tweddle:1975:AGC

[401] D. Tweddle. Age and growth of the catfish Bagrus meridionalis Günther in southern Lake Malaŵi. Journal of Fish Biology, 7(5):677–685, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1975:DTF

[402] C. R. Goldspink and J. W. Banks. A description of the Tjeukemeer Fishery together with a note upon the yield statistics between 1964 and 1970. Journal of Fish Biology, 7(5):687–708, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1975:HSC

[403] D. B. C. Scott. A hermaphrodite specimen of Coregonus lavaretus (L.) (Salmoniformes, Salmonidae) from Loch Lomond, Scotland. Journal of Fish Biology, 7(5):709, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1975:BRd

[404] Anonymous. Book reviews. Journal of Fish Biology, 7(5):711–716, September 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeVlaming:1975:DRP

[405] Victor L. De Vlaming, Martin Sage, and Robert Tiegs. A diurnal rhythm of pituitary prolactin activity with diurnal effects of mammalian and teleostean prolactin on total body lipid deposition and liver lipid metabolism in teleost fishes. Journal of Fish Biology, 7(6):717–726, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korting:1975:LDB

[406] Wolfgang Körting. Larval development of Bothriocephalus sp. (Cestoda: Pseudophyllidea) from carp (Cyprinm carpio L.) in Germany. Journal of Fish Biology, 7(6):727–733, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Webb:1975:RPN

[407] C. J. Webb and P. J. Miller. A redescription of Pomatoschistm norvegicus (Collett, 1903) (Teleostei: Gobioidei) based on syntype material. Journal of Fish Biology, 7(6):735–747, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woolland:1975:SGT

[408] J. V. Woolland and J. W. Jones. Studies on grayling, Thymallus thymallus L., in Llyn Tegid and the upper River Dee, North Wales. I. Age and growth. Journal of Fish Biology, 7(6):749–773, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gupta:1975:DCG

[409] S. Gupta. The development of carp gonads in warm water aquaria. Journal of Fish Biology, 7(6):775–782, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rao:1975:STL

[410] Teegavarapu R. Rao. Salinity tolerance of laboratory-reared larvae of the California killifish, Fundulus parvipinnis Girard. Journal of Fish Biology, 7(6):783–790, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weintraub:1975:IAL

[411] Matthew J. Weintraub. The influence of activity level on bradycardia in trout (Salmo gairdneri) during hypoxia. Journal of Fish Biology, 7 (6):791–796, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balon:1975:ELK

[412] Eugene K. Balon. The eels of Lake Kariba: distribution, taxonomic status, age, growth and density. Journal of Fish Biology, 7(6):797–815, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meakins:1975:EPP

[413] R. H. Meakins and M. Walkey. The effects of parasitism by the plerocercoid of Schistocephalus solidus Muller 1776 (Pseudophyllidea) on the respiration of the three-spined stickleback Gasterosteus aculeatus L. Journal of Fish Biology, 7(6):817–824, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Racicot:1975:BLE

[414] J. G. Racicot, M. Gaudet, and C. Leray. Blood and liver enzymes in rainbow trout (Salmo gairdneri Rich.) with emphasis on their diagnostic use: Study of CCl_{4} toxicity and a case of Aeromonas infection. Journal of Fish Biology, 7(6):825–835, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1975:BRe

[415] Anonymous. Book reviews. Journal of Fish Biology, 7(6):837–841, November 1975. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:1976:EHP

[416] Sterhen Ross. Effects of hyperbaric pressure on oxygen concentration in the swim bladder of Poecilia latipinna. Journal of Fish Biology, 8(1):1– 4, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:1976:DLP

[417] G. D. Robinson, W. A. Dunson, J. E. Wright, and G. E. Mamolito. Differences in low pH tolerance among strains of brook trout (Salvelinus fontinalis). Journal of Fish Biology, 8(1):5–17, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1976:SAM

[418] J. Hattingh. The surface area of mudfish. Journal of Fish Biology, 8(1): 19–22, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haarstein:1976:ASD

[419] Tore Hårstein and GrahamL. Bullock. An acute septicaemic disease of brown trout (Salmo trutta) and Atlantic salmon (Salmo salar) caused by a Pasteurella-like organism. Journal of Fish Biology, 8(1):23–26, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Michael:1976:CGA

[420] E. Michael and P. Chadwick. A comparison of growth and abundance for tidal pool fishes in California and British Columbia. Journal of Fish Biology, 8(1):27–34, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Child:1976:ETR

[421] A. R. Child, Ann M. Burnell, and N. P. Wilkins. The existence of two races of Atlantic salmon (Salmo salar L.) in the British Isles. Journal of Fish Biology, 8(1):35–43, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1976:RBS

[422] D. B. C. Scott and J. D. Fuller. The reproductive biology of Scleropages formosus (Müller & Schlegel) (Osteoglossomorpha, Osteoglossidae) in Malaya, and the morphology of its pituitary gland. Journal of Fish Biology, 8(1):45–53, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hettler:1976:ITS

[423] W. F. Hettler. Influence of temperature and salinity on routine metabolic rate and growth of young Atlantic menhaden. Journal of Fish Biology, 8(1):55–65, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:1976:DMF

[424] A. E. Ellis, A. L. S. Munroe, and R. J. Roberts. Defence mechanisms in fish: 1. A study of the phagocytic system and the fate of intraperitoneally injected particulate material in the plaice (Pleuronectes platessa L.). Journal of Fish Biology, 8(1):67–78, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ince:1976:ESF

[425] Bernard W. Ince and Alan Thorpe. The effects of starvation and forcefeeding on the metabolism of the northern pike, Esox lucius L. Journal of Fish Biology, 8(1):79–88, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1976:SUD

[426] Ronald J. Roberts and Barry J. Hill. Studies on ulcerative dermal necrosis of salmonids V. The histopathology of the condition in brown trout (Salmo trutta L.). Journal of Fish Biology, 8(1):89–92, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barron:1976:ORG

[427] R. J. C. Barron. The occurrence of the rock goby, Gobius paganellus L. 1758 and the two-spot goby, Chaparrudo flavescens (Fabricius, 1779) in the Blackwater Estuary, Essex (S.E. England). Journal of Fish Biology, 8(1):93–95, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1976:BRa

[428] Anonymous. Book reviews. Journal of Fish Biology, 8(1):97–100, January 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shotter:1976:DSH

[429] R. A. Shotter. The distribution of some helminth and copepod parasites in tissues of whiting, Merlangius merlangus L., from Manx waters. Journal of Fish Biology, 8(2):101–117, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arcement:1976:SKP

[430] Reese J. Arcement and Joseph W. Rachlin. A study of the karyotype of a population of banded killifish (Fundulus diaphanus) from the Hudson rive. Journal of Fish Biology, 8(2):119–125, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1976:SBL

[431] David B. Lewis. Studies of the biology of the lesser weever fish Trachinus vipera Cuvier: I. Adaptations to a benthic habit. Journal of Fish Biology, 8(2):127–138, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:1976:UPP

[432] H. W. Ferguson. The ultrastructure of plaice (Pleuronectes platessa) leucocytes. Journal of Fish Biology, 8(2):139–142, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:1976:LRC

[433] A. E. Ellis. Leucocytes and related cells in the plaice Pleuronectes platessa. Journal of Fish Biology, 8(2):143–156, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cadwallader:1976:BBN

[434] P. L. Cadwallader. Breeding biology of a non-diadromous galaxiid, Galaxias vulgaris Stokell, in a New Zealand river. Journal of Fish Biology, 8 (2):157–177, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1976:OAGa

[435] R. H. K. Mann. Observations on the age, growth, reproduction and food of the pike Esox lucius (L.) in two rivers in southern England. Journal of Fish Biology, 8(2):179–197, February 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buit:1976:OCC

[436] Marie Henriette Du Buit. The ovarian cycle of the cuckoo ray, Raja naevus (Müller and Henle), in the Celtic Sea. Journal of Fish Biology, 8(3):199– 207, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hinton:1976:ULC

[437] David E. Hinton and Charlotte R. Pool. Ultrastructure of the liver in channel catfish Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 8(3):209–219, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meakins:1976:VEC

[438] Robin H. Meakins. Variations in the energy content of freshwater fish. Journal of Fish Biology, 8(3):221–224, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Page:1976:HIC

[439] Jimmy W. Page, James W. Andrews, Takeshi Murai, and Morgan W. Murray. Hydrogen ion concentration in the gastrointestinal tract of channel catfish. Journal of Fish Biology, 8(3):225–228, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sorensen:1976:UCH

[440] Elsie M. B. Sorensen. Ultrastructural changes in the hepatocytes of green sunfish, Lepomis cyanellus Rafinesque, exposed to solutions of sodium arsenate. Journal of Fish Biology, 8(3):229–240, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasurer:1976:AGL

[441] James W. Treasurer. Age, growth and length–weight relationship of brown trout Salmo trutta (L.) in the Loch of Strathbeg, Aberdeenshire. Journal of Fish Biology, 8(3):241–253, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boxshall:1976:HSL

[442] G. A. Boxshall. The host specificity of Lepeophtheirus pectoralis (Müller, 1776) (Copepoda: Caligidae). Journal of Fish Biology, 8(3):255–264, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1976:OAGb

[443] R. H. K. Mann. Observations on the age, growth, reproduction and food of the chub Squalius cephalus (L.) in the River Stour, Dorset. Journal of Fish Biology, 8(3):265–288, March 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:1976:HST

[444] R. I. G. Morgan and R. J. Roberts. The histopathology of salmon tagging: IV. The effect of severe exercise on the induced tagging lesion in salmon parr at two temperatures. Journal of Fish Biology, 8(4):289–292, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1976:DPL

[445] C. R. Kennedy and S. F. Lie. The distribution and pathogenicity of larvae of Eustrongylides (Nematoda) in brown trout Salmo trutta L. in Fernworthy Reservoir, Devon. Journal of Fish Biology, 8(4):293–302, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otobo:1976:OMC

[446] Fidelis O. Otobo. Observations on meristic characters separating Pellonula afzeliusi Johnels, 1954 from Sierrathrissa leonensis Thys Van den Audernaerde, 1969. Journal of Fish Biology, 8(4):303–310, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Went:1976:RBS

[447] A. E. J. Went. The recapture of British salmon in Irish waters. Journal of Fish Biology, 8(4):311–316, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:1976:VDE

[448] D. H. McCarthy. Vibrio disease in eels. Journal of Fish Biology, 8(4): 317–320, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vivekanandan:1976:EFS

[449] E. Vivekanandan. Effects of feeding on the swimming activity and growth of Ophiocephalus striatus. Journal of Fish Biology, 8(4):321–330, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:1976:HRS

[450] C. D. Anderson, R. J. Roberts, K. MacKenzie, and A. H. McVicar. The hepato-renal syndrome in cultured turbot (Scophthalmus maximus L.). Journal of Fish Biology, 8(4):331–341, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1976:BRb

[451] Anonymous. Book reviews. Journal of Fish Biology, 8(4):343–344, April 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aneer:1976:LDB

[452] Gunnar Aneer and Olle Ljungberg. Lymphocystis disease in Baltic herring (Clupea harengus var. membras L.). Journal of Fish Biology, 8(5):345– 350, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welcomme:1976:SGT

[453] Robin L. Welcomme. Some general and theoretical considerations on the fish yield of African rivers. Journal of Fish Biology, 8(5):351–364, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heins:1976:RBA

[454] David C. Heins and Glenn H. Clemmer. The reproductive biology, age and growth of the North American cyprinid, Notropis longirostris (Hay).

Journal of Fish Biology, 8(5):365–379, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lampert:1976:ERF

[455] Winfried Lampert. Experiments on the resistance of fish to rapid increase in hydrostatic pressure. Journal of Fish Biology, 8(5):381–383, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1976:CEP

[456] R. J. Wootton and G. W. Evans. Cost of egg production in the threespined stickleback (Gasterosteus aculeatus L.). Journal of Fish Biology, 8 (5):385–395, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soivio:1976:HES

[457] Antti Soivio and Aimo Oikari. Haematological effects of stress on a teleost, Esox lucius L. Journal of Fish Biology, 8(5):397–411, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:1976:MGH

[458] Julio E. Perez and Norman Maclean. Multiple globins and haemoglobins in the bass, Dicentrarchus labrax (L.) (Serranidae: Teleostei). Journal of Fish Biology, 8(5):413–417, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ball:1976:MPJ

[459] Ian R. Ball and R. A. Khan. On Micropharynx parasitica Jägerskiöld, a marine planarian ectoparasitic on thorny skate, Raja radiata Donovan, from the North Atlantic Ocean. Journal of Fish Biology, 8(5):419–426, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1976:BRc

[460] Anonymous. Book reviews. Journal of Fish Biology, 8(5):427–430, May 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubieniecki:1976:ABP

[461] B. Lubieniecki. Aspects of the biology of the plerocercoid of Grillotia erinaceus (van Beneden, 1858) (Cestoda: Trypanorhyncha) in haddock Melanogrammus aeglefinus (L.). Journal of Fish Biology, 8(6):431–439, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faunge:1976:CSI

[462] Ragnar Fäunge, Ulf Lidman, and Åke Larsson. Comparative studies of inorganic substances in the blood of fishes from the Scagerac Sea. Journal of Fish Biology, 8(6):441–448, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bieniarz:1976:PRV

[463] K. Bieniarz and P. Epler. Preliminary results of the in vivo studies on ovarian resorption in carp (Cyprinus carpio L.). Journal of Fish Biology, 8(6):449–451, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ikusemiju:1976:DRG

[464] Kola Ikusemiju. Distribution, reproduction and growth of the catfish, Chrysichthys walkeri (Günther) in the lekki lagoon, Nigeria. Journal of Fish Biology, 8(6):453–458, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leggett:1976:OSM

[465] William C. Leggett and Robert N. O’Boyule. Osmotic stress and mortality in adult American shad during transfer from saltwater to freshwater. Journal of Fish Biology, 8(6):459–469, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smart:1976:EAE

[466] Geoffrey Smart. The effect of ammonia exposure on gill structure of the rainbow trout (Salmo gairdneri). Journal of Fish Biology, 8(6):471–475, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otto:1976:SGP

[467] Christian Otto. Size, growth, population density and food of brown trout Salmo trutta L. in two sections of a south Swedish stream. Journal of Fish Biology, 8(6):477–488, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Denton:1976:BCO

[468] Joan E. Denton and M. K. Yousef. Body composition and organ weights of rainbow trout, Salmo gairdneri. Journal of Fish Biology, 8(6):489– 499, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1976:BRd

[469] Anonymous. Book reviews. Journal of Fish Biology, 8(6):501–502, June 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Capra:1976:TRR

[470] M. F. Capra. A technique for recording the respiratory movements of unrestrained elasmobranch fish. Journal of Fish Biology, 9(1):1–2, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arawomo:1976:FFT

[471] G. A. O. Arawomo. Food and feeding of three Citharinus species in Lake Kainji, Nigeria. Journal of Fish Biology, 9(1):3–10, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarty:1976:EES

[472] L. S. McCarty and A. H. Houston. Effects of exposure to sublethal levels of cadmium upon water-electrolyte status in the goldfish (Carassius auratus). Journal of Fish Biology, 9(1):11–19, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hancock:1976:PRS

[473] R. S. Hancock, J. W. Jones, and R. Shaw. A preliminary report on the spawning behaviour and nature of sexual selection in the barbel, Barbus barbus (L.). Journal of Fish Biology, 9(1):21–28, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staples:1976:IBS

[474] D. J. Staples and M. Nomura. Influence of body size and food ration on the energy budget of rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 9(1):29–43, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:1976:DDE

[475] J. M. Elliott. The downstream drifting of eggs of brown trout, Salmo trutta L. Journal of Fish Biology, 9(1):45–50, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamagishi:1976:ESP

[476] Hiroshi Yamagishi. Experimental study on population dynamics in the guppy, Poeclia reticulata (Peters). Effect of shelters on the increase of population density. Journal of Fish Biology, 9(1):51–65, July 1976. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wirtz:1976:IOC

[477] P. Wirtz and J. Davenport. Increased oxygen consumption in blennies (Blennim pholis L.) exposed to their mirror images. Journal of Fish Biology, 9(1):67–74, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gupta:1976:AGC

[478] S. Gupta and Ch. Meske. Abnormalities of the gonads of carp. Journal of Fish Biology, 9(1):75–77, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munro:1976:ABE

[479] J. L. Munro. Aspects of the biology and ecology of Caribbean reef fishes: Mullidae (goat–fishes). Journal of Fish Biology, 9(1):79–97, July 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:1976:ETS

[480] A. K. Whitfield and S. J. M. Blaber. The effects of temperature and salinity on Tilapia rendalli Boulenger 1896. Journal of Fish Biology, 9(2): 99–104, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fuller:1976:RCC

[481] J. D. Fuller, D. B. C. Scott, and R. Fraser. The reproductive cycle of Coregonus lavaretus (L) in Loch Lomond, Scotland, inrelation to seasonal changes in plasma cortisol concentration. Journal of Fish Biology, 9(2): 105–117, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weiss:1976:HPP

[482] D. M. Weiss. A high power pulse generator for electric fishing — revised circuit. Journal of Fish Biology, 9(2):119–123, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naiman:1976:PHP

[483] Robert J. Naiman. Productivity of a herbivorous pupfish population (Cyprinodon nevadensis) in a warm desert stream. Journal of Fish Biology, 9(2):125–137, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:1976:BFSa

[484] J. W. Moore and I. A. Moore. The basis of food selection in flounders, Platichthys flesus (L.), in the Severn Estuary. Journal of Fish Biology, 9(2):139–156, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pemberton:1976:STNa

[485] R. Pemberton. Sea trout in North Argyll Sea lochs, population, distribution and movements. Journal of Fish Biology, 9(2):157–179, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atkinson:1976:SPF

[486] R. J. A. Atkinson. Some preliminary field observations of the burrows of the red band fish, Cepola rubescens L. Journal of Fish Biology, 9(2):181– 183, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rawson:1976:PBP

[487] Mac V. Rawson, Jr. Population biology of parasites of striped mullet, Mugil cephalus L. I. Monogenea. Journal of Fish Biology, 9(2):185–194, August 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pemberton:1976:STNb

[488] R. Pemberton. Sea trout in North Argyll Sea lochs: II. Diet. Journal of Fish Biology, 9(3):195–208, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:1976:FMB

[489] Imants G. Priede. Functional morphology of the bulbus arteriosus of rainbow trout (Salmo gairdneri Richardson). Journal of Fish Biology, 9 (3):209–216, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:1976:MFL

[490] Douglas D. Weber and Michael H. Schiewe. Morphology and function of the lateral line of juvenile steelhead trout in relation to gas-bubble disease. Journal of Fish Biology, 9(3):217–233, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulter:1976:BLS

[491] G. W. Coulter. The biology of Lates species (Nile perch) in Lake Tanganyika, and the status of the pelagic fishery for Lates species and Luciolates stappersii (Blgr.). Journal of Fish Biology, 9(3):235–259, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1976:MCE

[492] D. G. Cross. A method of comparing the efficiencies of electric fishing operations. Journal of Fish Biology, 9(3):261–265, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paperna:1976:PDS

[493] I. Paperna and D. E. Zwerner. Parasites and diseases of striped bass, Morone saxatilis (Walbaum), from the lower Chesapeake Bay. Journal of Fish Biology, 9(3):267–282, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1976:BRe

[494] Anonymous. Book reviews. Journal of Fish Biology, 9(3):283–287, September 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardner:1976:RFW

[495] Michael L. G. Gardner. A review of factors which may influence the seaage and maturation of Atlantic salmon Salmo salar L. Journal of Fish Biology, 9(4):289–327, October 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:1976:ILL

[496] C. K. Murray and T. C. Fletcher. The immunohistochemical localization of lysozyme in plaice (Pleuronectes platessa L.) tissues. Journal of Fish Biology, 9(4):329–334, October 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bannister:1976:LWR

[497] J. V. Bannister. The length–weight relationship, condition factor and gut contents of the dolphin-fish Coryphaena hippurus (L.) in the Mediterranean. Journal of Fish Biology, 9(4):335–338, October 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halliday:1976:BMC

[498] M. M. Halliday. The biology of Myxosoma cerebralis: the causative organism of whirling disease of salmonids. Journal of Fish Biology, 9(4): 339–357, October 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brafield:1976:OCS

[499] A. E. Brafield and P. Matthiessen. Oxygen consumption by sticklebacks (Gasterosteus aculeatus L.) exposed to zinc. Journal of Fish Biology, 9 (4):359–370, October 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanderElst:1976:IJM

[500] R. P. van der Elst and J. H. Wallace. Identification of the juvenile mullet of the east coast of South Africa. Journal of Fish Biology, 9(4):371– 374, October 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:1976:BFSb

[501] J. W. Moore and I. A. Moore. The basis of food selection in some estuarine fishes. eels, Anguilla anguilla (L.), whiting, Merlangius merlangus (L.), sprat, Sprattus sprattus (L.) and stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 9(5):375–390, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1976:PRR

[502] Alwyne Wheeler. On the populations of roach (Rutilus rutilus), rudd (Scardinius erythrophthalmus), and their hybriin esthwaite water, with notes on the distinctions between them. Journal of Fish Biology, 9(5): 391–400, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeCristini:1976:EAD

[503] F. De Cristini and G. Graziosi. Electrophoretic analyses and denaturation tests of the haemoglobins of Trachurus trachurs (L.). Journal of Fish Biology, 9(5):401–403, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1976:PAS

[504] J. E. Harris and B. J. Cottrell. Precipitating activity in the sera of plaice, Pleuronectes platessa L. to a helminth antigen. Journal of Fish Biology, 9 (5):405–410, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:1976:MBT

[505] D. J. Solomon and R. G. Templeton. Movements of brown trout Salmo trutta L. in a chalk stream. Journal of Fish Biology, 9(5):411–423, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larsson:1976:CSS

[506] Åke Larsson, Maj-Lis Johansson-Sjöbeck, and Ragnar Fänge. Comparative study of some haematological and biochemical blood parameters in fishes from the Skagerrak. Journal of Fish Biology, 9(5):425–440, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abel:1976:TAS

[507] P. D. Abel. Toxic action of several lethal concentrations of an anionic detergent on the gills of the brown trout (Salmo trutta L.). Journal of Fish Biology, 9(5):441–446, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:1976:HFS

[508] Julio E. Pérez and Norman Maclean. The haemoglobins of the fish Sarotherodon mossambicus (Peters): functional significance and ontogenetic changes. Journal of Fish Biology, 9(5):447–455, November 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madden:1976:UEF

[509] Jane A. Madden and Arthur H. Houston. Use of electroanaesthesia with freshwater teleosts: some physiological consequences in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 9(6):457–462, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oduleye:1976:EHP

[510] S. O. Oduleye. The effects of hypophysectomy, prolactin therapy and environmental calcium on freshwater survival and salinity tolerance in the brown trout Salmo trutta L. Journal of Fish Biology, 9(6):463–470, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brassington:1976:EIR

[511] R. A. Brassington and A. Ferguson. Electrophoretic identification of roach (Rutilus rutilus L.), rudd (scardinim erythrophthalmus L.), bream (Abramis brama L.) and their natural hybrids. Journal of Fish Biology, 9 (6):471–477, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hewett:1976:ARC

[512] C. J. Hewett and D. F. Jefferies. The accumulation of radioactive caesium from water by the brown trout (Salmo trutta) and its comparison with plaice and rays. Journal of Fish Biology, 9(6):479–489, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:1976:SOS

[513] R. W. Williams and M. C. Warner. Some observations on the stained blood cellular elements of channel catfish, Ictalurus punctatus. Journal of Fish Biology, 9(6):491–497, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulata:1976:EMA

[514] G. Hulata, R. Moav, and G. Wohlfarth. The effects of maternal age, relative hatching time and density of stocking on growth rate of fry in the European and Chinese races of the common carp. Journal of Fish Biology, 9(6):499–513, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sweeting:1976:SLI

[515] R. A. Sweeting. Studies on Ligula intestinalis (L.) effects on a roach population in a gravel pit. Journal of Fish Biology, 9(6):515–522, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stanley:1976:MAG

[516] Jon G. Stanley and Jack B. Jones. Morphology of androgenetic and gynogenetic grass carp, Ctenopharyngodon idella (Valenciennes). Journal of Fish Biology, 9(6):523–528, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromage:1976:HSR

[517] N. R. Bromage and A. Fuchs. A histological study of the response of the interrenal cells of the goldfish (Carassius auratus) to treatment with sodium lauryl sulphate. Journal of Fish Biology, 9(6):529–535, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1976:OHF

[518] Albert C. Smith and Franklin Ramos. Occult haemoglobin in fish skin mucus as an indicator of early stress. Journal of Fish Biology, 9(6):537– 541, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1976:BRf

[519] Anonymous. Book review. Journal of Fish Biology, 9(6):543–544, December 1976. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Darracott:1977:AMF

[520] Ann Darracott. Availability, morphometrics, feeding and breeding activity in a multi-species, demersal fish stock of the Western Indian Ocean. Journal of Fish Biology, 10(1):1–16, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodsaether:1977:CIF

[521] Marianne C. Rødsæther, Jan Olafsen, Jan Raa, Kjell Myhre, and Johan B. Steen. Copper as an initiating factor of vibriosis (Vibrio anguillarum) in eel (Anguilla anguilla). Journal of Fish Biology, 10(1):17–21, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mortensen:1977:PDY

[522] Erik Mortensen. The population dynamics of young trout (Salmo trutta L.) in a Danish brook. Journal of Fish Biology, 10(1):23–33, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1977:LTC

[523] C. R. Kennedy and A. Rumpus. Long-term changes in the size of the Pomphorhynchus laevis (Acanthocephala) population in the River Avon. Journal of Fish Biology, 10(1):35–42, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sweeting:1977:SLI

[524] R. A. Sweeting. Studies on Ligula intestinalis some aspects of the pathology in the second intermediate host. Journal of Fish Biology, 10(1):43–50, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macer:1977:SAB

[525] C. T. Macer. Some aspects of the biology of the horse mackerel [ Trachurus trachurus (L.)] in waters around Britain. Journal of Fish Biology, 10(1): 51–62, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingram:1977:PIR

[526] George A. Ingram and John B. Alexander. The primary immune response of brown trout (Salmo trutta) to injection with cellular antigens. Journal of Fish Biology, 10(1):63–72, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deVlaming:1977:EPP

[527] Victor L. de Vlaming and Mary Jo Vodicnik. Effects of pinealectomy on pituitary gonadotrophs, pituitary gonadotropin potency and hypothalamic gonadotropin releasing activity in Notemigonus crysoleucas. Journal of Fish Biology, 10(1):73–86, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eddy:1977:OUR

[528] F. B. Eddy. Oxygen uptake by rainbow trout blood, Salmo gairdneri. Journal of Fish Biology, 10(1):87–90, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soivio:1977:EAM

[529] Antti Soivio, Keijo Nyholm, and Marja Huhti. Effects of anaesthesia with MS 222, neutralized MS 222 and benzocaine on the blood constituents of rainbow trout, Salmo gairdneri. Journal of Fish Biology, 10(1):91– 101, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1977:BRa

[530] Anonymous. Book reviews. Journal of Fish Biology, 10(1):103–104, January 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ikusemiju:1977:FFH

[531] Kola Ikusemiju and C. I. O. Olaniyan. The food and feeding habits of the catfishes, Chrysichthys walkeri (Günther), Chrysichthys filamentosus (Boulenger) and Chrysichthys nigrodigitatus (Lacépède) in the Lekki Lagoon, Nigeria. Journal of Fish Biology, 10(2):105–112, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blasiola:1977:QSN

[532] George C. Blasiola, Jr. Quinaldine sulphate, a new anaesthetic formulation for tropical marine fishes. Journal of Fish Biology, 10(2):113–119, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laale:1977:BUZ

[533] Hans W. Laale. The biology and use of zebrafish, Brachydanio rerio in fisheries research. A literature review. Journal of Fish Biology, 10(2):121– 173, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffman:1977:CWDa

[534] Glenn L. Hoffman and Joseph J. O’Grodnick. Control of whirling disease (Myxosoma cerebralis): effects of drying, and disinfection with hydrated lime or chlorine. Journal of Fish Biology, 10(2):175–179, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffman:1977:CWDb

[535] Glenn L. Hoffman and Maria E. Markiw. Control of whirling disease (Myxosoma cerebralis): use of methylene blue staining as a possible indicator of effect of heat on spores. Journal of Fish Biology, 10(2):181–183, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siddiqui:1977:VFA

[536] Abdul Q. Siddiqui and Anwar H. Siddiqui. Variation in the free amino acid content of skeletal muscle of Ophicephalus punctatus (Bloch). Journal of Fish Biology, 10(2):185–189, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1977:BSI

[537] J. Hattingh. Blood sugar as an indicator of stress in the freshwater fish, Labeo capensis (Smith). Journal of Fish Biology, 10(2):191–195, February 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuchnow:1977:UCM

[538] Karl P. Kuchnow and James R. Scott. Ultrastructure of the chorion and its micropyle apparatus in the mature Fundulus heteroclitus (Walbaum) ovum. Journal of Fish Biology, 10(3):197–201, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferraro:1977:ATW

[539] Linda A. Ferraro, Richard E. Wolke, and Paul P. Yevich. Acute toxicity of water-borne dimethylnitrosamine (DMN) to Fundulus heteroclitus (L.). Journal of Fish Biology, 10(3):203–209, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parsons:1977:RBE

[540] J. Parsons, K. U. Vickers, and Y. Warden. Relationship between elver recruitment and changes in the sex ratio of silver eels Anguilla anguilla L. migrating from Lough Neagh, Northern Ireland. Journal of Fish Biology, 10(3):211–229, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:1977:HRM

[541] I. G. Priede and P. Tytler. Heart rate as a measure of metabolic rate in teleost fishes; Salmo gairdneri, Salmo trutta and Gadus morhua. Journal of Fish Biology, 10(3):231–242, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Infante:1977:SAB

[542] O. Infante and G. W. LaBar. Some aspects of the biology of Petenia krausii Steindachner (Pisces: Cichlidae) in Lake Valencia, Venezuela. Journal of Fish Biology, 10(3):243–249, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siddiqui:1977:RBL

[543] Abdul Q. Siddiqui. Reproductive biology, length–weight relationship and relative condition of Tilapia leucosticta (Trewavas) in Lake Naivasha, Kenya. Journal of Fish Biology, 10(3):251–260, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:1977:SED

[544] J. L. Turner. Some effects of demersal trawling in Lake Malaŵi (Lake Nyasa) from 1968 to 1974. Journal of Fish Biology, 10(3):261–271, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeVlaming:1977:VVE

[545] V. L. De Vlaming, J. Shing, G. Paquette, and R. Vuchs. In vivo and in vitro effects of oestradiol-17β on lipid metabolism in Notemigonus crysoleucas. Journal of Fish Biology, 10(3):273–285, March 1977. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goetz:1977:SOA

[546] F. W. Goetz, Jr., R. A. Hoffman, and W. L. Pancoej. A surgical operating apparatus for fish and its use in the pinealectomy of salmonids. Journal of Fish Biology, 10(3):287–290, March 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1977:SET

[547] M. Jobling, D. Gwyther, and D. J. Grove. Some effects of temperature, meal size and body weight on gastric evacuation time in the dab Limanda limanda (L). Journal of Fish Biology, 10(3):291–298, March 1977. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:1977:UTC

[548] I. G. Priede and A. H. Young. The ultrasonic telemetry of cardiac rhythms of wild brown trout (Salmo trutta L.) as an indicator of bio-energetics and behaviour. Journal of Fish Biology, 10(4):299–318, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minchew:1977:OFR

[549] C. Douglas Minchew and James D. Yarbrough. The occurrence of fin rot in mullet (Mugil cephalus) associated with crude oil contamination of an estuarine pond-ecosystem. Journal of Fish Biology, 10(4):319–323, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rajbanshi:1977:AGS

[550] V. K. Rajbanshi. The architecture of the gill surface of the catfish, Heteropneutes fossilis (Bloch): SEM study. Journal of Fish Biology, 10(4): 325–329, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:1977:RDF

[551] Elizabeth M. Wood. A review of damsel fishes (Pisces: Pomacentridae) of the genus Chromis from the central and eastern Atlantic and the Mediterranean. Journal of Fish Biology, 10(4):331–345, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Natarajan:1977:BPD

[552] P. Natarajan and P. S. B. R. James. A bibliography of parasites and diseases of marine and freshwater fishes of India. Journal of Fish Biology, 10(4):347–369, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deVlaming:1977:DVP

[553] Victor L. de Vlaming and Mary Jo Vodicnik. Diurnal variations in pituitary gonadotropin content and in gonadal response to exogenous gonadotropin and prolactin in Notemigonus crysoleucas. Journal of Fish Biology, 10(4):371–383, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tweddle:1977:AGN

[554] D. Tweddle and J. L. Turner. Age, growth and natural mortality rates of some cichlid fishes of Lake Malaŵi. Journal of Fish Biology, 10(4):385– 398, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthiessen:1977:ULD

[555] Peter Matthiessen and A. E. Brafield. Uptake and loss of dissolved zinc by the stickleback Gasterosteus aculeatus L. Journal of Fish Biology, 10(4): 399–410, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boxshall:1977:HIL

[556] G. A. Boxshall. The histopathology of infection by Lepeophtheirus pectoralis (Müller) (Copepoda: Caligidae). Journal of Fish Biology, 10(4): 411–415, April 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordon:1977:FPIa

[557] J. D. M. Gordon. The fish populations in inshore waters of the west of Scotland. The biology of the Norway pout (Trisopterus esmarkii). Journal of Fish Biology, 10(5):417–430, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meyer:1977:DCE

[558] Wilfried Meyer and Ingrid Sauerbier. The development of catecholamines in embryos and larvae of the rainbow trout (Salmo gairdneri Rich.). Journal of Fish Biology, 10(5):431–435, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jafri:1977:FSY

[559] A. K. Jafri and S. Mustafa. Food selection by young Catla catla (Hamilton-Buchanan) in a tropical fish pond. Journal of Fish Biology, 10(5):437– 440, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rawson:1977:PBP

[560] Mac V. Rawson, Jr. Population biology of parasites of striped mullet, Mugil cephalus L. Crustacea. Journal of Fish Biology, 10(5):441–451, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hattingh:1977:ETM

[561] J. Hattingh. The effect of tricaine methanesulphonate (MS-222) on the microhaematocrit of fish blood. Journal of Fish Biology, 10(5):453–455, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fickeisen:1977:DAA

[562] D. H. Fickeisen and G. W. Brown, Jr. D-amino acid oxidase in various fishes. Journal of Fish Biology, 10(5):457–465, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pascoe:1977:EPT

[563] D. Pascoe and P. Cram. The effect of parasitism on the toxicity of cadmium to the three-spined stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 10(5):467–472, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:1977:BRR

[564] J. L. Allen and J. B. Hunn. Book review: Renal excretion in channel catfish following injection of quinaldine sulphate or 3- trifluoromethyl-4-nitrophenol. Journal of Fish Biology, 10(5):473–479, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casillas:1977:ESB

[565] Edmundo Casillas and Lynwood S. Smith. Effect of stress on blood coagulation and haematology in rainbow trout (Salmo gairdneri). Journal of Fish Biology, 10(5):481–491, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Man:1977:SIFa

[566] Hanson S. H. Man and I. J. Hodgkiss. Studies on the ichthyo-fauna in Plover Cove Reservoir, Hong Kong: I. Sequence of fish population changes. Journal of Fish Biology, 10(5):493–503, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1977:SHE

[567] Alan D. Pickering and D. J. Macey. Structure, histochemistry and the effect of handling on the mucous cells of the epidermis of the char Salvelinus alpinus (L.). Journal of Fish Biology, 10(5):505–512, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neish:1977:OSA

[568] Gordon A. Neish. Observations on saprolegniasis of adult sockeye salmon, Oncorhynchus nerka (Walbaum). Journal of Fish Biology, 10(5):513–522, May 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ashburner:1977:MHC

[569] L. D. Ashburner. Mycobacteriosis in hatchery-confined Chinook salmon (Oncorhynchus tshawytscua Walbaum) in Australia. Journal of Fish Biology, 10(6):523–528, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horsley:1977:RBF

[570] R. W. Horsley. A review of the bacterial flora of teleosts and elasmobranchs, including methods for its analysis. Journal of Fish Biology, 10 (6):529–553, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bleniarz:1977:IIS

[571] K. Bleniarz and P. Epler. Investigations on inducing sexual maturity in the male eel Anguilla anguilla L. Journal of Fish Biology, 10(6):555–559, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1977:SCE

[572] Alan D. Pickering. Seasonal changes in the epidermis of the brown trout Salmo trutta (L.). Journal of Fish Biology, 10(6):561–566, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mishra:1977:HPA

[573] N. Mishra, P. K. Pandey, J. S. Datta Munshi, and B. R. Singh. Haematological parameters of an air-breathing mud eel, Amphipnous cuchia (Ham.) (Amphipnoidae; Piscesa). Journal of Fish Biology, 10(6):567– 573, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chliamovitch:1977:BHH

[574] Y.-P. Chliamovitch and C. Kuhn. Behavioural, haematological and histological studies on acute toxicity of bis(tri-n-butyltin) oxide on Salmo gairdneri Richardson and Tilapia rendalli Boulenger. Journal of Fish Biology, 10(6):575–585, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordon:1977:FPIb

[575] J. D. M. Gordon. The fish populations in inshore waters of the West Coast of Scotland. The distribution, abundance and growth of the whiting (Merlangius merlangus L.). Journal of Fish Biology, 10(6):587–596, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1977:BRb

[576] Anonymous. Book reviews. Journal of Fish Biology, 10(6):597–600, June 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Man:1977:SIFb

[577] Hanson S. H. Man and I. J. Hodgkiss. Studies on the ichthyo-fauna in Plover Cove Reservoir, Hong Kong: feeding and food relations. Journal of Fish Biology, 11(1):1–13, July 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khalil:1977:GKS

[578] L. F. Khalil. Gyrometra kunduchi n. sp., a cestodarian from Plectorhinchus pictus (Thunberg, 1792) from the Indian Ocean. Journal of Fish Biology, 11(1):15–19, July 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mudry:1977:HAP

[579] Dwight R. Mudry and R. Stewart Anderson. Helminth and arthropod parasites of fishes in the mountain national parks of Canada. Journal of Fish Biology, 11(1):21–33, July 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cottrell:1977:TPP

[580] B. J. Cottrell. A trypanosome from the plaice, Pleuronectes platessa L. Journal of Fish Biology, 11(1):35–47, July 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weis:1977:EHM

[581] Judith S. Weis and Peddrick Weis. Effects of heavy metals on development of the killifish, Fundulus heteroclitus. Journal of Fish Biology, 11(1):49– 54, July 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1977:DRA

[582] J. E. Thorpe. Daily ration of adult perch, Perca fluviatilis L. during summer in Loch Leven, Scotland. Journal of Fish Biology, 11(1):55–68, July 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haastein:1977:SVA

[583] T. Håstein and J. E. Smith. A study of Vibrio anguillarum from farmed and wild fish using principal components analysis. Journal of Fish Biology, 11(2):69–75, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spoor:1977:ORE

[584] W. A. Spoor. Oxygen requirements of embryos and larvae of the largemouth bass, Micropterus salmoides (Lacépède). Journal of Fish Biology, 11(2):77–86, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wurtsbaugh:1977:ETR

[585] Wayne A. Wurtsbaugh and Gerald E. Davis. Effects of temperature and ration level on the growth and food conversion efficiency of Salmo gairdneri, Richardson. Journal of Fish Biology, 11(2):87–98, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wurtsbaugh:1977:EFS

[586] Wayne A. Wurtsbaugh and Gerald E. Davis. Effects of fish size and ration level on the growth and food conversion efficiency of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 11(2):99–104, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dipper:1977:AGF

[587] F. A. Dipper, C. R. Bridges, and A. Menz. Age, growth and feeding in the ballan wrasse Labrus bergylta Ascanius 1767. Journal of Fish Biology, 11(2):105–120, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordon:1977:FPIc

[588] J. D. M. Gordon. The fish populations of inshore waters of the west coast of Scotland. The unusual occurrence of the blue whiting (Micromesistius poutassou) and some notes on its biology. Journal of Fish Biology, 11 (2):121–124, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Child:1977:OMB

[589] A. R. Child and D. J. Solomon. Observations on morphological and biochemical features of some cyprinid hybrids. Journal of Fish Biology, 11 (2):125–131, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borjeson:1977:EHB

[590] Hans Börjeson. Effects of hypercapnia on the buffer capacity and haematological values in Salmo salar (L.). Journal of Fish Biology, 11(2):133– 142, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1977:RBC

[591] M. S. Caulton and E. Bursell. The relationship between changes in condition and body composition in young Tilapia rendalli Boulenger. Journal of Fish Biology, 11(2):143–150, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coles:1977:TDD

[592] T. F. Coles, G. N. Swinney, and J. W. Jones. A technique for determining the distribution of pelagic fish larvae. Journal of Fish Biology, 11(2):151– 159, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1977:SCD

[593] John F. Craig. Seasonal changes in the day and night activity of adult perch, Perca fluviatilis L. Journal of Fish Biology, 11(2):161–166, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stroband:1977:GDD

[594] H. W. J. Stroband. Growth and diet dependant structural adaptations of the digestive tract in juvenile grass carp (Ctenopharyngodon idella, Val.). Journal of Fish Biology, 11(2):167–174, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1977:BDL

[595] J. E. Thorpe. Bimodal distribution of length of juvenile Atlantic salmon (Salmo salar L.) under artificial rearing conditions. Journal of Fish Biology, 11(2):175–184, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Savitz:1977:ERL

[596] J. Savitz, Elizabeth Albanese, Mary Jean Evinger, and Pamela Kolasinski. Effect of ration level on nitrogen excretion, nitrogen retention and efficiency of nitrogen utilization for growth in largemouth bass (Micropterus salmoides). Journal of Fish Biology, 11(2):185–192, August 1977. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phromsuthirak:1977:EMW

[597] Pramarn Phromsuthirak. Electron microscopy of wound healing in the skin of Gasterosteus aculeatus. Journal of Fish Biology, 11(2):193–206, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pascoe:1977:STC

[598] D. Pascoe and D. L. Mattey. Studies on the toxicity of cadmium to the three-spined stickleback Gasterosteus aculeatus L. Journal of Fish Biology, 11(2):207–215, August 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:1977:SCB

[599] J. S. Campbell. Spawning characteristics of brown trout and sea trout Salmo trutta L. in Kirk Burn, River Tweed, Scotland. Journal of Fish Biology, 11(3):217–229, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henricson:1977:ADD

[600] Jan Henricson. The abundance and distribution of Diphyllobothrium dendriticum (Nitzsch) and D. ditremum (Creplin) in the char Salvelinus alpinus (L.) in Sweden. Journal of Fish Biology, 11(3):231–248, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezzat:1977:MEA

[601] A. Ezzat and S. El-Serafy. The migration of elvers of Anguilla anguilla L. in the Mex canal, Alexandria, Egypt. Journal of Fish Biology, 11(3): 249–256, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kislalioglu:1977:FRS

[602] M. Kislalioglu and R. N. Gibson. The feeding relationship of shallow water fishes in a Scottish sea loch. Journal of Fish Biology, 11(3):257– 266, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shireman:1977:ICG

[603] J. V. Shireman, D. E. Colle, and R. W. Rottmann. Intensive culture of grass carp, Ctenopharyngodon idella, in circular tanks. Journal of Fish Biology, 11(3):267–272, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanDyke:1977:DDL

[604] J. M. Van Dyke and D. L. Sutton. Digestion of duckweed (Lemna spp.) by the grass carp (Ctenopharyngodon idella). Journal of Fish Biology, 11 (3):273–278, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reddy:1977:PSC

[605] S. Ravichandra Reddy, Katre Shakuntala, and K. V. Rajagopal. Preliminary studies on the conversion of Tubifex tubifex as food by elvers of Anguilla nebulosa (Gray and Hardwicke). Journal of Fish Biology, 11 (3):279–281, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingram:1977:SPC

[606] George A. Ingram and John B. Alexander. Serum protein changes in brown trout (Salmo trutta L.) after single injections of soluble and cellular antigens. Journal of Fish Biology, 11(3):283–291, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ching:1977:SSG

[607] Chong Vino Ching. Studies on the small grey mullet Liza malinoptera (Valenciennes). Journal of Fish Biology, 11(3):293–308, September 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Daham:1977:STG

[608] Najim K. Al-Daham and M. Nazir Bhatti. Salinity tolerance of Gambusia affinis (Baird & Girard) and Heteropneustes fossilis (Bloch). Journal of Fish Biology, 11(4):309–313, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:1977:FFM

[609] B. F. Blake. Food and feeding of the mormyrid fishes of Lake Kainji, Nigeria, with special reference to seasonal variation and interspecific differences. Journal of Fish Biology, 11(4):315–328, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:1977:AAE

[610] Hiram W. Li and Robert W. Brocksen. Approaches to the analysis of energetic costs of intraspecific competition for space by rainbow trout (Salmo gairdneri). Journal of Fish Biology, 11(4):329–341, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kallner:1977:SCC

[611] Anders Kallner. A study of the conversion of cholesterol to bile acids by the use of a fistulated bile duct. Journal of Fish Biology, 11(4):343– 347, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1977:ELF

[612] Alan D. Pickering and L. G. Willoughby. Epidermal lesions and fungal infection on the perch, Perca fluviatilis L., in Windermere. Journal of Fish Biology, 11(4):349–354, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horne:1977:PVJ

[613] M. T. Horne, R. H. Richards, R. J. Roberts, and P. C. Smith. Peracute vibriosis in juvenile turbot Scophthalmus maximus. Journal of Fish Biology, 11(4):355–361, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:1977:RCC

[614] D. J. Solomon. A review of chemical communication in freshwater fish. Journal of Fish Biology, 11(4):363–376, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1977:CRP

[615] Anonymous. Comment on recent publications. Journal of Fish Biology, 11(4):377–378, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1977:BRc

[616] Anonymous. Book reviews. Journal of Fish Biology, 11(4):379–383, October 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warner:1977:CBS

[617] Mark C. Warner and R. Whitney Williams. Comparison between serum values of pond and intensive raceway cultured channel catfish Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 11(5):385–391, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulter:1977:AEF

[618] G. W. Coulter. Approaches to estimating fish biomass and potential yield in Lake Tanganyika. Journal of Fish Biology, 11(5):393–408, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yadav:1977:EPI

[619] M. Yadav and S. Diana Ooi. Effect of photoperiod on the incorporation of ^{3}H-thymidine into the gonads of Carassius auratus. Journal of Fish Biology, 11(5):409–416, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mosse:1977:FRD

[620] Peter R. L. Mosse and Richard C. L. Hudson. The functional roles of different muscle fibre types identified in the myotomes of marine teleosts: a behavioural, anatomical and histochemical study. Journal of Fish Biology, 11(5):417–430, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Love:1977:ADM

[621] R. Malcolm Love, Lesley J. Munro, and Ian Robertson. Adaptation of the dark muscle of cod to swimming activity. Journal of Fish Biology, 11 (5):431–436, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:1977:ARB

[622] B. F. Blake. Aspects of the reproductive biology of Hippopotamyms pictus from Lake Kainji, with notes on four other mormyrid species. Journal of Fish Biology, 11(5):437–445, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prejs:1977:RBF

[623] Andrzej Prejs and Mieczyslaw Blaszczyk. Relationships between food and cellulase activity in freshwater fishes. Journal of Fish Biology, 11(5):447– 452, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:1977:LFR

[624] A. E. Ellis. The leucocytes of fish: a review. Journal of Fish Biology, 11 (5):453–491, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1977:NID

[625] T. F. Cross and R. H. Payne. NADP-isocitrate dehydrogenase polymorphism in the Atlantic salmon Salmo salar. Journal of Fish Biology, 11 (5):493–496, November 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthiessen:1977:UME

[626] P. Matthiessen. Uptake, metabolism and excretion of the molluscicide N-tritylmorpholine by the tropical food fish Sarotherodon mossambicus (Peters). Journal of Fish Biology, 11(5):497–509, November 1977. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jolly:1977:SMP

[627] Donald W. Jolly. Still more pages in 1978. Journal of Fish Biology, 11(6): 511, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordon:1977:FPId

[628] J. D. M. Gordon. The fish populations in inshore waters of the west coast of Scotland. The food and feeding of the whiting (Merlangius merlangus L.). Journal of Fish Biology, 11(6):513–529, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frost:1977:FCS

[629] Winifred E. Frost. The food of charr, Salvelinus willughbii (Günther), in Windermere. Journal of Fish Biology, 11(6):531–547, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1977:ETR

[630] M. S. Caulton. The effect of temperature on routine metabolism in Tilapia rendalli Boulenger. Journal of Fish Biology, 11(6):549–553, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubieniecki:1977:PGE

[631] B. Lubieniecki. The plerocercus of Grillotia erinaceus as a biological tag for haddock Melanogrammus aeglefinus in the North Sea and north-east Atlantic. Journal of Fish Biology, 11(6):555–565, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amend:1977:VSS

[632] Donald F. Amend and John R. Nelson. Variation in the susceptibility of sockeye salmon Oncorhynchus nerka to infectious haemopoietic necrosis virus. Journal of Fish Biology, 11(6):567–573, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1977:CSG

[633] Ian A. Johnston. A comparative study of glycolysis in red and white muscles of the trout (Salmo gairdneri) and mirror carp (Cyprinus carpio). Journal of Fish Biology, 11(6):575–588, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coleman:1977:USM

[634] R. Coleman, Z. Yaron, and Z. Ilan. An ultrastructural study of the mitochondria-rich ‘chloride ‘cells from the gills of freshwater and seawateradapted Tilapia aurea subjected to a pesticide. Journal of Fish Biology, 11(6):589–594, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larson:1977:ATI

[635] G. L. Larson, F. E. Hutchins, and D. A. Schlesinger. Acute toxicity of inorganic chloramines to early life stages of brook trout (Salvelinus fontinalis). Journal of Fish Biology, 11(6):595–598, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1977:RMR

[636] C. R. Goldspink. The return of marked roach (Rutilus rutilus L.) to spawning grounds in Tjeukemeer, The Netherlands. Journal of Fish Biology, 11 (6):599–603, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sauer:1977:EAS

[637] D. M. Sauer and G. Haider. Enzyme activities in the serum of rainbow trout, Salmo gairdneri Richardson; the effects of water temperature. Journal of Fish Biology, 11(6):605–612, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mortensen:1977:DDM

[638] Erik Mortensen. Density-dependent mortality of trout fry (Salmo trutta L.) and its relationship to the management of small streams. Journal of Fish Biology, 11(6):613–617, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1977:PBT

[639] C. R. Kennedy and R. Burrough. The population biology of two species of eyefluke, Diplostomum gasterostei and Tylodelphys clavata, in perch. Journal of Fish Biology, 11(6):619–633, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahy:1977:CFO

[640] Edward Fahy. Characteristics of the freshwater occurrence of sea trout Salmo trutta in Ireland. Journal of Fish Biology, 11(6):635–646, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egglishaw:1977:GSP

[641] Henry J. Egglishaw and P. E. Shackley. Growth, survival and production of juvenile salmon and trout in a Scottish stream, 1966–75. Journal of Fish Biology, 11(6):647–672, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alexander:1977:PCS

[642] J. B. Alexander. Protein concentration in the serum of the Atlantic salmon (Salmo salar) in North-west England and Northern Ireland. Journal of Fish Biology, 11(6):673–679, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottaway:1977:MAF

[643] Elisabeth M. Ottaway and K. Simkiss. A method for assessing factors influencing ‘false check’formation in fish scales. Journal of Fish Biology, 11(6):681–687, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1977:BRd

[644] Anonymous. Book reviews. Journal of Fish Biology, 11(6):689–692, December 1977. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sokolowska:1978:HPH

[645] M. Sokolowska, P. Epler, and K. Bieniarz. The histological picture of the hypothalamus (the nucleus preopticus) and hypophysis in male Anguilla anguilla L. treated with hormones. Journal of Fish Biology, 12(1):1– 4, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fox:1978:POD

[646] P. J. Fox. Preliminary observations on different reproduction strategies in the bullhead (Cottus gobio L.) in northern and southern England. Journal of Fish Biology, 12(1):5–11, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shieh:1978:CBE

[647] H. S. Shieh. Changes of blood enzymes in brook trout induced by infection with Aeromonas salmonicida. Journal of Fish Biology, 12(1):13–18, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellgaard:1978:LAB

[648] Erik G. Ellgaard, Joseph E. Tusa, and Anthony A. Malizia, Jr. Locomotor activity of the bluegill Lepomis macrochirus: hyperactivity induced by sublethal concentrations of cadmium, chromium and zinc. Journal of Fish Biology, 12(1):19–23, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanBanning:1978:LTS

[649] Paul van Banning and Herman B. Becker. Long-term survey data (1965– 1972) on the occurrence of Anisakis larvae (Nematoda: Ascaridida) in herring, Clupea harengus L., from the North Sea. Journal of Fish Biology, 12(1):25–33, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1978:OPA

[650] D. Louise Barber and Jean E. Mills Westermann. Occurrence of the periodic acid-Schiff positive granular leucocyte (PAS-GL) in some fishes and its significance. Journal of Fish Biology, 12(1):35–43, January 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Totland:1978:LIA

[651] Geir K. Totland, Harald Kryvi, and Erik Slinde. LDH isoenzymes in the axial muscle of the sharks Galeus melastomus and Etmopterus spinax. Journal of Fish Biology, 12(1):45–50, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christensen:1978:CPC

[652] G. M. Christensen, James T. Fiandt, and Barbara A. Poeschl. Cells, proteins, and certain physical-chemical properties of brook trout (Salvelinus fontinalis) blood. Journal of Fish Biology, 12(1):51–60, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:1978:SEH

[653] A. K. Whitfield and S. J. M. Blaber. Scale-eating habits of the marine teleost Terapon jarbua (Forsskål). Journal of Fish Biology, 12(1):61–70, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacLeod:1978:ESS

[654] M. G. MacLeod. Effects of salinity and starvation on the alimentary canal anatomy of the rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 12(1):71–79, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okera:1978:FTB

[655] W. Okera. Fishes taken by the beach seines fishing at Lumley, Freetown (Sierra Leone). Journal of Fish Biology, 12(1):81–88, January 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1978:NGT

[656] B. Stott and B. R. Buckley. A note on the growth of two exotics in England; the silver carp (Hypophthalmichthys molitrix Val.) and the bighead (Aristichthys nobilis Rich.). Journal of Fish Biology, 12(1):89–91, January 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smart:1978:ITM

[657] G. R. Smart. Investigations of the toxic mechanisms of ammonia to fish– gas exchange in rainbow trout (Salmo gairdneri) exposed to acutely lethal concentrations. Journal of Fish Biology, 12(1):93–104, January 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramos:1978:KBF

[658] Franklin Ramos and Albert C. Smith. Ketone bodies in fish skin mucus as an indicator of starvation: a preliminary report. Journal of Fish Biology, 12(2):105–108, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paperna:1978:SSD

[659] I. Paperna. Swimbladder and skeletal deformations in hatchery bred Spams aurata. Journal of Fish Biology, 12(2):109–114, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:1978:ABE

[660] R. Thompson and J. L. Munro. Aspects of the biology and ecology of Caribbean reef fishes: Serranidae (hinds and groupers). Journal of Fish Biology, 12(2):115–146, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1978:SBE

[661] C. R. Kennedy. Studies on the biology of Eubothrium salvelini and E. crassum in resident and migratory Salvelinus alpinus and Salmo trutta and in S. salar in North Norway and the islands of Spitsbergen and Jan Mayen. Journal of Fish Biology, 12(2):147–162, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunter:1978:SFS

[662] C. R. Hunter and P. L. Nayudu. Surface folds in superficial epidermal cells of three species of teleost fish. Journal of Fish Biology, 12(2):163– 166, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1978:HCR

[663] Alwyne Wheeler and Keith Easton. Hybrids of chub and roach (Leuciscus cephalus and Rutilus rutilus) in English rivers. Journal of Fish Biology, 12(2):167–171, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cook:1978:KAG

[664] Peter C. Cook. Karyotypic analysis of the gobiid fish genus Quietula Jordan and Evermann. Journal of Fish Biology, 12(2):173–179, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1978:EMG

[665] W. J. R. Lanzing. Effect of methallibure on gonad development and carotenoid content of the fins of Sarotherodon mossambicus (Tilapia mossambica). Journal of Fish Biology, 12(2):181–185, February 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vodicnik:1978:EPP

[666] Mary Jo Vodicnik, Roman E. Kral, Victor L. de Vlaming, and Larry W. Crim. The effects of pinealectomy on pituitary and plasma gonadotropin levels in Carassius auratus exposed to various photoperiod–temperature regimes. Journal of Fish Biology, 12(3):187–196, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dykova:1978:HCF

[667] I. Dykova and J. Lom. Histopathological changes in fish gills infected with myxosporidian parasites of the genus Henneguya. Journal of Fish Biology, 12(3):197–202, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claridge:1978:GMT

[668] P. N. Claridge and D. C. Gardner. Growth and movements of the twaite shad, Alosa fallax (Lacépède), in the Severn Estuary. Journal of Fish Biology, 12(3):203–211, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Badsha:1978:ABH

[669] K. S. Badsha and M. Sainsbury. Aspects of the biology and heavy metal accumulation of Ciliata mustela. Journal of Fish Biology, 12(3):213–220, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:1978:SIP

[670] A. Ferguson, K-J. M. Himberg, and G. Svärdson. Systematics of the Irish pollan (Coregonus pollan Thompson): an electrophoretic comparison with other Holarctic Coregoninae. Journal of Fish Biology, 12(3):221– 233, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shackley:1978:SSW

[671] Susan E. Shackley and P. E. King. A sterile sea water circulating system for the jar culture of developing fish eggs. Journal of Fish Biology, 12(3):

235–237, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedley:1978:CFB

[672] R. B. Pedley and J. W. Jones. The comparative feeding behaviour of brown trout, Salmo trutta L. and Atlantic salmon, Salmo salar L. in Llyn Dwythwch, Wales. Journal of Fish Biology, 12(3):239–256, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kryvi:1978:FTL

[673] Harald Kryvi and Geir K. Totland. Fibre types in locomotory muscles of the cartilaginous fish Chimaera monstrosa. Journal of Fish Biology, 12(3):257–265, March 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCain:1978:FDP

[674] Bruce B. McCain, Mark S. Myers, William D. Gronlund, S. R. Wellings, and Charles E. Alpers. The frequency, distribution, and pathology of three diseases of demersal fishes in the Bering Sea. Journal of Fish Biology, 12 (4):267–276, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:1978:ECD

[675] Diane G. Elliott and Donald F. Amend. Efficacy of certain disinfectants against infectious pancreatic necrosis virus. Journal of Fish Biology, 12 (4):277–286, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crisp:1978:EIR

[676] D. T. Crisp, R. H. K. Mann, and Jean C. McCormack. The effects of impoundment and regulation upon the stomach contents of fish at Cow Green, Upper Teesdale. Journal of Fish Biology, 12(4):287–301, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adron:1978:SCN

[677] J. W. Adron and A. M. Mackie. Studies on the chemical nature of feeding stimulants for rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 12(4):303–310, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moller:1978:EST

[678] H. Möller. The effects of salinity and temperature on the development and survival of fish parasites. Journal of Fish Biology, 12(4):311–323, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1978:SYC

[679] G. J. A. Kennedy and C. D. Strange. Seven years on — a continuing investigation of salmonid stocks in Lough Erne tributaries. Journal of Fish Biology, 12(4):325–330, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1978:ACF

[680] R. J. Wootton, G. W. Evans, and L. Mills. Annual cycle in female threespined sticklebacks (Gasterosteus aculeatus L.) from an upland and lowland population. Journal of Fish Biology, 12(4):331–343, April 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broughton:1978:BRI

[681] Nigel M. Broughton and Neville V. Jones. Book review: An investigation into the growth of 0-group roach, (Rutilus rutilus L.) with special reference to temperature. Journal of Fish Biology, 12(4):345–357, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dixon:1978:EAQ

[682] R. N. Dixon and P. Milton. The effects of the anaesthetic quinaldine on oxygen consumption in an intertidal teleost Blennius pholis (L.). Journal of Fish Biology, 12(4):359–369, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1978:BSP

[683] R. N. Gibson and I. A. Ezzi. The biology of a Scottish population of Fries’ goby, Lesueurigobius friesii. Journal of Fish Biology, 12(4):371–389, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1978:BRa

[684] Anonymous. Book reviews. Journal of Fish Biology, 12(4):391–392, April 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1978:BSH

[685] C. R. Kennedy. The biology, specificity and habitat of the species of Eubothrium (Cestoda: Pseudophyllidea), with reference to their use as biological tags: a review. Journal of Fish Biology, 12(5):393–410, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hazen:1978:OEC

[686] Terry C. Hazen and Gerald W. Esch. Observations on the ecology of Clinostomum marginatum in largemouth bass (Micropterus salmoides). Journal of Fish Biology, 12(5):411–420, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1978:COG

[687] C. R. Goldspink. Comparative observations on the growth rate and year class strength of roach Rutilus rutilus L. in two Cheshire lakes, England. Journal of Fish Biology, 12(5):421–433, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1978:IMR

[688] Alwyne Wheeler. Ictalurus melas (Rafinesque, 1820) and I. nebulosus (Lesueur, 1819): the North American catfishes in Europe. Journal of Fish Biology, 12(5):435–439, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1978:NFV

[689] Alan D. Pickering. A note on the failure of vitamin a to influence the epidermis of the brown trout, Salmo trutta L. Journal of Fish Biology, 12(5):441–447, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1978:EPD

[690] Helen T. Smith, Carl B. Schreck, and O. Eugene Maughan. Effect of population density and feeding rate on the fathead minnow (Pimephales promelas). Journal of Fish Biology, 12(5):449–455, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shireman:1978:GGC

[691] Jerome V. Shireman, Douglas E. Colle, and Roger W. Rottmann. Growth of grass carp fed natural and prepared diets under intensive culture. Journal of Fish Biology, 12(5):457–463, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daniels:1978:NBH

[692] Robert A. Daniels. Nesting behaviour of Harpagifer bispinis in Arthur Harbour, Antarctic Peninsula. Journal of Fish Biology, 12(5):465–474, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beumer:1978:FEF

[693] J. P. Beumer. Feeding ecology of four fishes from a mangrove creek in north Queensland, Australia. Journal of Fish Biology, 12(5):475–490, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hazen:1978:RST

[694] Terry C. Hazen, Gerald W. Esch, Armand B. Glassman, and J. Whitfield Gibbons. Relationship of season, thermal loading and red-sore disease with various haematological parameters in Micropterus salmoides. Journal of Fish Biology, 12(5):491–498, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:1978:IJN

[695] D. J. Solomon and A. R. Child. Identification of juvenile natural hybrids between Atlantic salmon (Salmo solar L.) and trout(Salmo trutta L.). Journal of Fish Biology, 12(5):499–501, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ince:1978:EIP

[696] Bernard W. Ince and Alan Thorpe. The effects of insulin on plasma amino acid levels in the northern pike, Esox lucius L. Journal of Fish Biology, 12(5):503–506, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grove:1978:SAF

[697] D. J. Grove, L. G. Loizides, and J. Nott. Satiation amount, frequency of feeding and gastric emptying rate in Salmo gairdneri. Journal of Fish Biology, 12(5):507–516, May 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jolly:1978:E

[698] Donald W. Jolly. Editorial. Journal of Fish Biology, 12(6):517–518, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osborn:1978:SCT

[699] R. H. Osborn and T. H. Simpson. Seasonal changes in thyroidal status in the plaice, Pleuronectes platessa L. Journal of Fish Biology, 12(6):519– 526, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katz:1978:EDT

[700] Anne Harris Katz and Harvey M. Katz. Effects of DL-thyroxine on swimming speed in pearl danio Brachydanio albolineatus (Blyth). Journal of Fish Biology, 12(6):527–530, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osborn:1978:SDR

[701] R. H. Osborn, T. H. Simpson, and A. F. Youngson. Seasonal and diurnal rhythms of thyroidal status in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 12(6):531–540, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1978:PAS

[702] J. E. Thorpe and R. I. G. Morgan. Periodicity in Atlantic salmon Salmo salar L. smolt migration. Journal of Fish Biology, 12(6):541–548, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCleave:1978:URT

[703] J. D. McCleave, J. H. Power, and S. A. Rommel, Jr. Use of radio telemetry for studying upriver migration of adult Atlantic salmon (Salmo salar). Journal of Fish Biology, 12(6):549–558, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCleave:1978:RAE

[704] J. D. McCleave. Rhythmic aspects of estuarine migration of hatcheryreared Atlantic salmon (Salmo salar) smolts. Journal of Fish Biology, 12(6):559–570, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:1978:SOS

[705] D. J. Solomon. Some observations on salmon smolt migration in a chalkstream. Journal of Fish Biology, 12(6):571–574, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tytler:1978:UTM

[706] P. Tytler, J. E. Thorpe, and W. M. Shearer. Ultrasonic tracking of the movements of Atlantic salmon smolts (Salmo salar L.) in the estuaries of two Scottish rivers. Journal of Fish Biology, 12(6):575–586, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1978:SAP

[707] J. S. Alabaster and B. Stott. Swimming activity of perch, Perca fluviatilis L. Journal of Fish Biology, 12(6):587–591, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelso:1978:DRA

[708] J. R. M. Kelso. Diel rhythm in activity of walleye, Stizostedion vitreum vitreum. Journal of Fish Biology, 12(6):593–599, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:1978:SCR

[709] C. Whitehead, N. R. Bromage, and J. R. M. Forster. Seasonal changes in reproductive function of the rainbow trout (Salmo gairdneri). Journal of Fish Biology, 12(6):601–608, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katz:1978:CRJ

[710] H. M. Katz. Circadian rhythms in juvenile American shad, Alosa sapidissima. Journal of Fish Biology, 12(6):609–614, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottaway:1978:RGA

[711] E. M. Ottaway. Rhythmic growth activity in fish scales. Journal of Fish Biology, 12(6):615–623, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1978:A

[712] Anonymous. Amouncements. Journal of Fish Biology, 12(6):624, June 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1978:TDE

[713] M. S. Caulton. Tissue depletion and energy utilisation during routine metabolism by sub-adult Tilapia rendalli Boulenger. Journal of Fish Biology, 13(1):1–6, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swift:1978:SEE

[714] D. J. Swift. Some effects of exposing rainbow trout (Salmo gairdneri Richardson) to phenol solutions. Journal of Fish Biology, 13(1):7–17, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burrough:1978:PBT

[715] R. J. Burrough. The population biology of two species of eyefluke, Diplostomum spathaceum and Tylodelphys clavata, in roach and rudd. Journal of Fish Biology, 13(1):19–32, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monaco:1978:CET

[716] Paul J. Monaco, Ellen M. Rasch, and Joseph S. Balsano. Cytological evidence for temporal differences during the asynchronous ovarian maturation of bisexual and unisexual fishes of the genus Poecilia. Journal of Fish Biology, 13(1):33–44, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulata:1978:STS

[717] G. Hulata and S. Rothbard. ‘Siamese-twins’ in Sarotherodon mossambicus. Journal of Fish Biology, 13(1):45, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bloom:1978:PPM

[718] Herbert D. Bloom and Alfred Perlmutter. Possible pheromone mediated reproductive isolation in two species of cyprinid fishes of the genus, brachydanio. Journal of Fish Biology, 13(1):47–50, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henricson:1978:DID

[719] Jan Henricson. The dynamics of infection of Diphyllobothrium dendriticum (Nitzsch) and D. ditremum (Creplin) in the char Salvelinus alpinus (L.) in Sweden. Journal of Fish Biology, 13(1):51–71, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacLeod:1978:RBD

[720] M. G. MacLeod. Relationships between dietary sodium chloride, food intake and food conversion in the rainbow trout. Journal of Fish Biology, 13(1):73–78, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piron:1978:SSD

[721] R. D. Piron. Spontaneous skeletal deformities in the zebra danio (Brachydanio rerio) bred for fish toxicity tests. Journal of Fish Biology, 13(1): 79–83, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hislop:1978:OEF

[722] J. R. G. Hislop, A. P. Robb, and J. A. Gauld. Observations on effects of feeding level on growth and reproduction in haddock, Melanogrammus aeglefinus (L.) in captivity. Journal of Fish Biology, 13(1):85–98, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1978:IHT

[723] M. S. Caulton. The importance of habitat temperatures for growth in the tropical cichlid Tilapia rendalli Boulenger. Journal of Fish Biology, 13(1): 99–112, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giles:1978:BAF

[724] Robert C. Giles, Lewis R. Brown, and C. Douglas Minchew. Bacteriological aspects of fin erosion in mullet exposed to crude oil. Journal of Fish Biology, 13(1):113–117, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piron:1978:BCC

[725] R. D. Piron. Breeding the convict cichlid (Cichlasoma nigrofasciatum) for use in laboratory fish toxicity tests. Journal of Fish Biology, 13(1): 119–122, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahy:1978:VSB

[726] E. Fahy. Variation in some biological characteristics of British sea trout, Salmo trutta L. Journal of Fish Biology, 13(1):123–138, July 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1978:BRb

[727] Anonymous. Book reviews. Journal of Fish Biology, 13(1):139–142, July 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hewett:1978:ARC

[728] C. J. Hewett and D. F. Jefferies. The accumulation of radioactive caesium from food by the plaice (Pleuronectes platessa) and the brown trout (Salmo trutta). Journal of Fish Biology, 13(2):143–153, August 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sturm:1978:ABS

[729] M. G.de L. Sturm. Aspects of the biology of Scomberomorus maculatus (Mitchill) in Trinidad. Journal of Fish Biology, 13(2):155–172, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1978:PBG

[730] A. C. Smith. Pathology and biochemical genetic variation in the milkfish, Chanos chanos. Journal of Fish Biology, 13(2):173–177, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shackley:1978:PYS

[731] S. E. Shackley and P. E. King. Protein yolk synthesis in Blennius pholis L. Journal of Fish Biology, 13(2):179–193, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caulton:1978:ETM

[732] M. S. Caulton. The effect of temperature and mass on routine metabolism in Sarotherodon (Tilapia) mossambicus (Peters). Journal of Fish Biology, 13(2):195–201, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chinniah:1978:MAF

[733] V. C. Chinniah and William Threlfall. Metazoan arasites of fish from the Smallwood Reservoir, Labrador, Canada. Journal of Fish Biology, 13(2):203–213, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nagy:1978:ICC

[734] A. Nagy, K. Rajki, L. Horvárth, and V. Csárnyi. Investigation on carp, Cyprinus carpio L. gynogenesis. Journal of Fish Biology, 13(2):215–224, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burrough:1978:IBP

[735] R. J. Burrough and C. R. Kennedy. Interaction between perch (Perca fltuviatilis), and brown trout, (Salmo trutta). Journal of Fish Biology, 13(2):225–230, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1978:TSS

[736] D. H. Jones and A. W. G. John. A three-spined stickleback, Gasterosteus aculeatus L. from the North Atlantic. Journal of Fish Biology, 13(2):231– 236, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olson:1978:PES

[737] R. E. Olson. Parasitology of the English sole, Parophrys vetulus Girard in Oregon, U.S.A. Journal of Fish Biology, 13(2):237–248, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sullivan:1978:PPB

[738] J. F. Sullivan and G. J. Atchison. Predator-prey behaviour of fathead minnows, Pimephales promelas and largemouth bass, Micropterus salmoides in a model ecosystem. Journal of Fish Biology, 13(2):249–253, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1978:AMP

[739] C. R. Kennedy. An analysis of the metazoan parasitocoenoses of brown trout Salmo trutta from British lakes. Journal of Fish Biology, 13(2):255– 263, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1978:SBR

[740] C. R. Kennedy, P. F. Broughton, and P. M. Hine. The status of brown and rainbow trout, Salmo trutta and S. gairdneri as hosts of the acanthocephalan, Pomphorhynchus laevis. Journal of Fish Biology, 13(2):265– 275, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singh:1978:AHM

[741] S. M. Singh and Peter N. Ferns. Accumulation of heavy metals in rainbow trout Salmo gairdneri (Richardson) maintained on a diet containing activated sewage sludge. Journal of Fish Biology, 13(2):277–286, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:1978:UOB

[742] C. Blake and B. F. Blake. The use of opercular bones in the study of age and growth in Labeo senegalensis from Lake Kainji, Nigeria. Journal of Fish Biology, 13(2):287–295, August 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perera:1978:SCB

[743] P. A. B. Perera and S. S. De Silva. Studies on the chemical biology of young grey mullet, Mugil cephalus L. Journal of Fish Biology, 13(3):

297–304, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanVuren:1978:SSH

[744] J. H. J. van Vuren and J. Hattingh. A seasonal study of the haematology of wild freshwater fish. Journal of Fish Biology, 13(3):305–313, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warner:1978:EDT

[745] M. C. Warner, S. A. Diehl, and A. M. Tomb. Effects of dilution and temperature of analysis on blood serum values in rainbow trout, Salmo gairdneri. Journal of Fish Biology, 13(3):315–319, September 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bhatti:1978:ACC

[746] M. Nazir Bhatti and Najim K. Al-Daham. Annual cyclical changes in the testicular activity of a freshwater teleost, Barbus luteus (Heckel) from Shatt-Al-Arab, Iraq. Journal of Fish Biology, 13(3):321–326, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murphy:1978:CZC

[747] B. R. Murphy, G. J. Atchison, Alan W. McIntosh, and D. J. Kolar. Cadmium and zinc content of fish from an industrially contaminated lake. Journal of Fish Biology, 13(3):327–335, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bortone:1978:HPR

[748] S. A. Bortone, W. K. Bradley, and J. L. Oglesby. The host — parasite relationship of two copepod species and two fish species. Journal of Fish Biology, 13(3):337–350, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Htun-Han:1978:RBDa

[749] M. Htun-Han. The reproductive biology of the dab Limanda limanda (L.) in the North Sea: Seasonal changes in the ovary. Journal of Fish Biology, 13(3):351–359, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Htun-Han:1978:RBDb

[750] M. Htun-Han. The reproductive biology of the dab Limanda limanda (L.) in the North Sea: Seasonal changes in the testis. Journal of Fish Biology, 13(3):361–367, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Htun-Han:1978:RBDc

[751] M. Htun-Han. The reproductive biology of the dab Limanda limanda (L.) in the North Sea: gonosomatic index, hepatosomatic index and condition factor. Journal of Fish Biology, 13(3):369–378, September 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:1978:SEF

[752] R. W. G. White and W. P. Williams. Studies of the ecology of fish populations in the Rye Meads sewage effluent lagoons. Journal of Fish Biology, 13(4):379–400, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sivarajah:1978:EPB

[753] K. Sivarajah, C. S. Franklin, and W. P. Williams. The effects of polychlorinated biphenyls on plasma steroid levels and hepatic microsomal enzymes in fish. Journal of Fish Biology, 13(4):401–409, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sivarajah:1978:SHE

[754] K. Sivarajah, C. S. Franklin, and W. P. Williams. Some histopathological effects of Aroclor 1254 on the liver and gonads of rainbow trout, Salmo gairdneri and carp, Cyprinus carpio. Journal of Fish Biology, 13(4):411– 414, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Junqueira:1978:BRE

[755] L. C. U. Junqueira and L. M. M. Salles. Book review: Effects of 6hydroxydopamine and 5-hydroxydopamine onlthe teleost melanophores innervation. Journal of Fish Biology, 13(4):415–419, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woo:1978:CBC

[756] N. Y. S. Woo, H. A. Bern, and R. S. Nishioka. Changes in body composition associated with smoltification and premature transfer to seawater in Coho salmon (Oncorhynchus kisutch) and king salmon (o. tschawytscha). Journal of Fish Biology, 13(4):421–428, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

WarBurton:1978:AGD

[757] K. WarBurton. Age and growth determination in a marine catfish using an otolith check technique. Journal of Fish Biology, 13(4):429–434, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardiner:1978:FSS

[758] D. M. Gardiner. Fine structure of the spermatozoon of the viviparous teleost, Cymatogaster aggregata. Journal of Fish Biology, 13(4):435–438, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gee:1978:PJA

[759] A. S. Gee, N. J. Milner, and R. J. Hemsworth. The production of juvenile Atlantic salmon, Salmo salar in the Upper Wye, Wales. Journal of Fish Biology, 13(4):439–451, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nayudu:1978:GCM

[760] P. L. Nayudu and C. R. Hunter. Genetic control of melanophore ultrastructure in Poecilia reticulata. Journal of Fish Biology, 13(4):453–456, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1978:PFR

[761] C. R. Kennedy. The parasite fauna of resident char Salvelinus alpinus from Arctic islands, with special reference to Bear Island. Journal of Fish Biology, 13(4):457–466, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1978:HBA

[762] Alwyne Wheeler. Hybrids of bleak, Alburnus alburnus, and chub, Leuciscus cephalus in English rivers. Journal of Fish Biology, 13(4):467–473, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walker:1978:SSS

[763] M. Greer Walker and L. Emerson. Sustained swimming speeds and myotomal muscle function in the trout, Salmo gairdneri. Journal of Fish Biology, 13(4):475–481, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chao:1978:SMF

[764] L. P. Chao and T. H. Birkbeck. Storage of marine fish erythrocytes in liquid nitrogen. Journal of Fish Biology, 13(4):483–488, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Went:1978:TFF

[765] A. E. J. Went. Trigger or file-fish (Balistes carolinensis) in Irish waters. Journal of Fish Biology, 13(4):489–492, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1978:NDP

[766] C. R. Goldspink. A note on the dispersion pattern of marked bream Abramis brama released into Tjeukemeer, The Netherlands. Journal of Fish Biology, 13(4):493–497, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1978:PDG

[767] C. R. Goldspink. The population density, growth rate and production of bream, Abramis brama, in Tjeukemeer, The Netherlands. Journal of Fish Biology, 13(4):499–517, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mittal:1978:NCA

[768] A. K. Mittal and M. Whitear. A note on cold anaesthesia of poikilotherms. Journal of Fish Biology, 13(4):519–520, October 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deCiechomski:1978:SGJ

[769] J. D. de Ciechomski and M. C. Cassia. Studies on the growth of juveniles of Cynoscion striatus in the sea and in aquaria. Journal of Fish Biology, 13 (5):521–526, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulbert:1978:HLE

[770] W. C. Hulbert and T. W. Moon. A histochemical, light and electron microscopic examination of eel, Anguilla rostrata red and white muscle. Journal of Fish Biology, 13(5):527–533, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mattey:1978:CPT

[771] D. L. Mattey, R. Moate, and M. Morgan. Comparison of ‘pseudobranch’ type and ‘chloride’ type cells in the pseudobranch of marine, freshwater and euryhaline teleosts. Journal of Fish Biology, 13(5):535–542, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolis:1978:VEA

[772] L. Bolis and J. C. Rankin. Vascular effects of acetylcholine, catecholamines and detergents on isolated perfused gills of pink salmon, Oncorhynchus gorbuscha Coho salmon, O. kisutch and chum salmon, O. keta. Journal of Fish Biology, 13(5):543–547, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1978:PIG

[773] J. E. Thorpe and R. I. G. Morgan. Parental influence on growth rate, smolting rate and survival in hatchery reared juvenile Atlantic salmon, Salmo salar. Journal of Fish Biology, 13(5):549–556, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valtonen:1978:CFS

[774] E. T. Valtonen and T. Valtonen. Cystidicola farionis as a swimbladder parasite of the whitefish in the Bothnian Bay. Journal of Fish Biology, 13 (5):557–561, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1978:ODM

[775] D. L. Barber and J. E. Mills Westermann. Observations on development and morphological effects of histamine liberator 48/80 on PAS-positive granular leukocytes and heterophils of Catostomus commersoni. Journal of Fish Biology, 13(5):563–573, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthiesen:1978:AOB

[776] P. Matthiesen and J. S. Johnson. Accumulation of the organophosphate blackfly larvicide Abate (Temephos) in Sarotherodon mossambicus, with reference to the larvicidal control of Simulium damnosum. Journal of Fish Biology, 13(5):575–586, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Georgi:1978:HEK

[777] T. A. Georgi and Dennis Beedle. The histology of the excretory kidney of the paddlefish, Polyodon spathula. Journal of Fish Biology, 13(5):587– 590, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waiwood:1978:ECH

[778] K. G. Waiwood and F. W. H. Beamish. The effect of copper, hardness and pH on the growth of rainbow trout, Salmo gairdneri. Journal of Fish Biology, 13(5):591–598, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milner:1978:PBT

[779] N. J. Milner, A. S. Gee, and R. J. Hemsworth. The production of brown trout, Salmo trutta in tributaries of the Upper Wye, Wales. Journal of Fish Biology, 13(5):599–612, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broughton:1978:BCL

[780] N. M. Broughton and G. Goldspink. Biochemical changes in the lateral muscles of roach, Rutilus rutilus from two habitats following exercise. Journal of Fish Biology, 13(5):613–618, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holanov:1978:PFF

[781] S. H. Holanov and J. C. Tash. Particulate and filter feeding in threadfin shad, Dorosoma petenense, at different light intensities. Journal of Fish Biology, 13(5):619–625, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payne:1978:GPD

[782] A. I. Payne. Gut pH and digestive strategies in estuarine grey mullet (Mugilidae) and tilapia (Cichlidae). Journal of Fish Biology, 13(5):627– 629, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beattie:1978:CUR

[783] J. H. Beattie and D. Pascoe. Cadmium uptake by rainbow trout, Salmo gairdneri eggs and alevins. Journal of Fish Biology, 13(5):631–637, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beach:1978:UIR

[784] M. H. Beach. The use of infra-red light and closed circuit TV to validate records from automatic fish counters. Journal of Fish Biology, 13(5):639– 644, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vandenBroek:1978:DHF

[785] W. L. F. van den Broek. Dietary habits of fish populations in the Lower Medway Estuary. Journal of Fish Biology, 13(5):645–654, November 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hawkins:1978:HAS

[786] A. D. Hawkins and A. D. F. Johnstone. The hearing of the Atlantic salmon, Salmo salar. Journal of Fish Biology, 13(6):655–673, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:1978:FFE

[787] A. K. Whitfield and S. J. M. Blaber. Food and feeding ecology of piscivorous fishes at Lake St Lucia, Zululand. Journal of Fish Biology, 13 (6):675–691, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pal:1978:ECA

[788] J. Pal, B. C. Pal, and R. Bannerjee. Epithelial carcinoma in Anabas testudineus. Journal of Fish Biology, 13(6):693–694, December 1978. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1978:ATC

[789] M. Lewis. Acute toxicity of copper, zinc and manganese in single and mixed salt solutions to juvenile longfin dace, Agosia chrysogaster. Journal of Fish Biology, 13(6):695–700, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benoit:1978:TEZ

[790] D. A. Benoit and G. W. Holcombe. Toxic effects of zinc on fathead minnows Pimephales promelas in soft water. Journal of Fish Biology, 13 (6):701–708, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vandenBroek:1978:ELB

[791] W. L. F. van den Broek. The effects of Lernaeocera branchialis on the Merlangius merlangus population in the Medway Estuary. Journal of Fish Biology, 13(6):709–715, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1978:ZCG

[792] G. M. Hughes and R. Flos. Zinc content of the gills of rainbow trout (S. gairdneri) after treatment with zinc solutions under normoxic and hypoxic conditions. Journal of Fish Biology, 13(6):717–728, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oswald:1978:UTS

[793] R. L. Oswald. The use of telemetry to study light synchronization with feeding and gill ventilation rates in Salmo trutta. Journal of Fish Biology, 13(6):729–739, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hill:1978:DES

[794] B. J. Hill and A. J. Ribbink. Depth equilibration of a shallow-water cichlid fish. Journal of Fish Biology, 13(6):741–745, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1978:BRc

[795] Anonymous. Book reviews. Journal of Fish Biology, 13(6):747–749, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1978:PJR

[796] Anonymous. Publications and journals received. Journal of Fish Biology, 13(6):749, December 1978. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:1979:FTS

[797] R. N. Campbell. Ferox trout, Salmo trutta L., and chair, Salvelinm alpinus (L.), in Scottish lochs. Journal of Fish Biology, 14(1):1–29, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1979:BCA

[798] D. H. Lewis, T. E. Eurell, M. S. Cannon, and L. C. Grumbles. T and B cell analogues from peripheral blood of immune channel catfish, Ictalurus punctatus. Journal of Fish Biology, 14(1):31–37, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezzat:1979:BVS

[799] A. A. Ezzat, M. T. Hashem, and M. M. El Gharabawy. Biometric variations in Solea vulgaris acclimatized in Lake Quarun, Upper Egypt. Journal of Fish Biology, 14(1):39–46, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madhavi:1979:OOA

[800] R. Madhavi. Observations on the occurrence of Allocreadium fasciatusi in Aplocheilus melastigma. Journal of Fish Biology, 14(1):47–58, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kapur:1979:ECC

[801] K. Kapur and H. S. Toor. The effect of clomiphene citrate on ovulation and spawning in indomethacin treated carp, Cyprinus carpio. Journal of Fish Biology, 14(1):59–66, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siddiqui:1979:HRL

[802] A. Q. Siddiqui and S. M. Naseem. The haematology of rohu, Labeo rohita. Journal of Fish Biology, 14(1):67–72, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linfield:1979:ADY

[803] R. S. J. Linfield. Age determination and year class structure in a stunted roach, Rutilus rutilus population. Journal of Fish Biology, 14(1):73–87, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wankowski:1979:MLP

[804] J. W. J. Wañkowski. Morphological limitations, prey size selectivity, and growth response of juvenile Atlantic salmon, Salmo salar. Journal of Fish Biology, 14(1):89–100, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biro:1979:AES

[805] Peter Biró. Acute effects of the sodium salt of 2,4-D on the early developmental stages of bleak, Alburnus alburnus. Journal of Fish Biology, 14 (1):101–109, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coombs:1979:DEE

[806] S. H. Coombs and A. R. Hiby. The development of the eggs and early larvae of blue whiting, Micromesistius poutassou and the effect of temperature on development. Journal of Fish Biology, 14(1):111–123, January 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jilek:1979:IHD

[807] Reid Jilek. Intestinal histology of Dorosoma cepedianum. Journal of Fish Biology, 14(2):125–126, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

John:1979:PRF

[808] M. J. John and C. L. Mahajan. The physiological response of fishes to a deficiency of cyanocobalamin and folic acid. Journal of Fish Biology, 14 (2):127–133, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1979:AEH

[809] B. Stott and B. R. Buckley. Avoidance experiments with homing shoals of minnows, Phoxinus phoxinus in a laboratory stream channel. Journal of Fish Biology, 14(2):135–146, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:1979:CBP

[810] S. Clark, D. H. Whitmore, Jr., and R. F. McMahon. Considerations of blood parameters of largemouth bass, Micropterus salmoides. Journal of Fish Biology, 14(2):147–158, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lane:1979:SHR

[811] H. C. Lane. Some haematological responses of normal and splenectomized rainbow trout Salmo gairdneri to a 12% blood loss. Journal of Fish Biology, 14(2):159–164, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carrick:1979:EAW

[812] T. R. Carrick. The effect of acid water on the hatching of salmonid eggs. Journal of Fish Biology, 14(2):165–172, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scullion:1979:MHR

[813] J. Scullion and R. W. Edwards. The movement of hatchery-reared rainbow and brown trout in a polluted river in the South Wales coalfield. Journal of Fish Biology, 14(2):173–178, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sis:1979:MAO

[814] R. F. Sis, P. J. Ives, D. M. Jones, D. H. Lewis, and W. E. Haensly. The microscopic anatomy of the oesophagus, stomach and intestine of the channel catfish, Ictalurus punctatus. Journal of Fish Biology, 14(2):179– 186, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1979:MAA

[815] S. V. Lewis. The morphology of the accessory air-breathing organs of the catfish, Clarias batrachus: a SEM study. Journal of Fish Biology, 14 (2):187–191, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ince:1979:MEP

[816] B. W. Ince. Metabolic effects of partial pancreatectomy in the northern pike, Esox indus L. Journal of Fish Biology, 14(2):193–198, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Srivastava:1979:BDT

[817] A. K. Srivastava and Shashikala Mishra. Blood dyscrasia in a teleost, Colisa fasdatus after acute exposure to sublethal concentrations of lead. Journal of Fish Biology, 14(2):199–203, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dudley:1979:CGS

[818] Richard G. Dudley. Changes in growth and size distribution of Sarotherodon macrochir and Sarotherodon andersoni from the Kafue Floodplain, Zambia, since construction of the Kafue Gorge Dam. Journal of Fish Biology, 14(2):205–223, February 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesthagen:1979:NEW

[819] I. H. Hesthagen and Jens H. Koefoed. A note on the existence of a welldeveloped swimbladder in the bottom-living goby Pomatoschistus minutus. Journal of Fish Biology, 14(3):225–228, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flowerdew:1979:SOE

[820] M. W. Flowerdew and D. J. Grove. Some observations of the effects of body weight, temperature, meal size and quality on gastric emptying time in the turbot, Scophthalmus maximus (L.) using radiography. Journal of Fish Biology, 14(3):229–238, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wankowski:1979:SDF

[821] J. W. J. Wańkowski and J. E. Thorpe. Spatial distribution and feeding in Atlantic salmon, Salmo salar L. juveniles. Journal of Fish Biology, 14(3):239–247, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingram:1979:IBT

[822] G. A. Ingram and J. B. Alexander. The immunoglobulin of the brown trout, Salmo trutta and its concentration in the serum of antigenstimulated and non-stimulated fish. Journal of Fish Biology, 14(3):249– 260, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:1979:HGRa

[823] L. G. Ross. The haemodynamics of gas resorption from the physoclist swimbladder: the structure and morphometrics of the oval in Pollachim virens (L.). Journal of Fish Biology, 14(3):261–266, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamada:1979:SMH

[824] S. Behrens Yamada, T. J. Mulligan, and S. J. Fairchild. Strontium marking of hatchery-reared Coho salmon (Oncorhynchm kisutch, Walbaum). Journal of Fish Biology, 14(3):267–275, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1979:NOR

[825] D. L. Barber, J. E. Mills Westermann, and D. N. Jensen. New observations on the rodlet cell (Rhabdospora thelohani) in the white sucker Catostomus commersoni (Lacépède): LM and EM studies. Journal of Fish Biology, 14(3):277–284, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1979:VRT

[826] G. M. Hughes and K. Nyholm. Ventilation in rainbow trout (Salmo gairdneri, Richardson) with damaged gills. Journal of Fish Biology, 14(3): 285–288, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMahon:1979:UCT

[827] Thomas E. McMahon and Jerry C. Tash. The use of chemosenses by threadfin shad, Dorosoma petenense to detect conspecifics, predators and food. Journal of Fish Biology, 14(3):289–296, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe-Jinde:1979:SOR

[828] L. Lowe-Jinde. Some observations of rainbow trout, Salmo gairdneri, Richardson, infected with Cryptobia salmositica. Journal of Fish Biology, 14(3):297–302, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pascoe:1979:RCP

[829] D. Pascoe and J. H. Beattie. Resistance to cadmium by pretreated rainbow trout alevins. Journal of Fish Biology, 14(3):303–308, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chernoff:1979:HMR

[830] B. Chernoff and J. K. Dooley. Heavy metals in relation to the biology of the mummichog, Fundulus heteroclitus. Journal of Fish Biology, 14(3): 309–328, March 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nzioka:1979:OSS

[831] R. M. Nzioka. Observations on the spawning seasons of East African reef fishes. Journal of Fish Biology, 14(4):329–342, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tandler:1979:MBC

[832] A. Tandler and F. W. H. Beamish. Mechanical and biochemical components of apparent specific dynamic action in largemouth bass, Micropterm salmoides Lacéepède. Journal of Fish Biology, 14(4):343–350, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wankowski:1979:RFP

[833] J. W. J. Wańkowski and J. E. Thorpe. The role of food particle size in the growth of juvenile Atlantic salmon (Salmo salar L.). Journal of Fish Biology, 14(4):351–370, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vandenBroek:1979:CEM

[834] W. L. F. van den Broek. Copepod ectoparasites of Merlangius merlangus and Platichthys flesus. Journal of Fish Biology, 14(4):371–380, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Daham:1979:ACO

[835] Najim K. Al-Daham and M. Nazir Bhatti. Annual changes in the ovarian activity of the freshwater teleost, Barbus luteus (Heckel) from Southern Iraq. Journal of Fish Biology, 14(4):381–387, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:1979:HGRb

[836] L. G. Ross. The haemodynamics of gas resorption from the physoclist swimbladder. II. The determination of blood flow rate in the oval of Pollachius virens (L.) using radiolabelled microspheres. Journal of Fish Biology, 14(4):389–393, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vandenBroek:1979:IEF

[837] W. L. F. van den Broek. Infection of estuarine fish populations by Cryptocotyle lingua (Creplin). Journal of Fish Biology, 14(4):395–402, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kashin:1979:DMS

[838] S. M. Kashin, A. G. Feldman, and G. N. Orlovsky. Different modes of swimming of the carp, Cyprinus carpio L. Journal of Fish Biology, 14(4): 403–405, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sauer:1979:EAP

[839] D. M. Sauer and G. Haider. Enzyme activities in the plasma of rainbow trout, Salmo gairdneri Richardson; the effects of nutritional status and salinity. Journal of Fish Biology, 14(4):407–412, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahajan:1979:SVB

[840] C. L. Mahajan and J. S. Dheer. Seasonal variations in the blood constituents of an air-breathing fish, Channa punctatus Bloch. Journal of Fish Biology, 14(4):413–417, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:1979:OBB

[841] B. E. Marshall. Observations on the breeding biology of Sarotherodon macrochir (Boulenger) in Lake Mcllwaine, Rhodesia. Journal of Fish Biology, 14(4):419–424, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reynolds:1979:PRD

[842] W. W. Reynolds and M. E. Casterlin. Photokinetic responses and diel activity of sea lamprey Petromyzon marinus ammocoete larvae. Journal of Fish Biology, 14(4):425–428, April 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gruber:1979:CIF

[843] D. Gruber, J. Cairns, Jr., K. L. Dickson, and A. C. Hendricks. A cinematographic investigation into the fish’s bioelectric breathing signal. Journal of Fish Biology, 14(5):429–436, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babiker:1979:SBRa

[844] M. M. Babiker and H. Ibrahim. Studies on the biology of reproduction in the cichlid Tilapia nilotica (L.): gonadal maturation and fecundity. Journal of Fish Biology, 14(5):437–448, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warburton:1979:GPS

[845] K. Warburton. Growth and production of some important species of fish in a Mexican coastal lagoon system. Journal of Fish Biology, 14(5):449– 464, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atkins:1979:MHC

[846] D. L. Atkins and W. L. Fink. Morphology and histochemistry of the caudal gland of Corynopoma riisei Gill. Journal of Fish Biology, 14(5): 465–469, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dartnall:1979:PMS

[847] H. J. G. Dartnall and M. Walkey. Parasites of marine sticklebacks. Journal of Fish Biology, 14(5):471–474, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shackley:1979:AAI

[848] S. E. Shackley and P. E. King. Amino acid incorporation by isolated oocytes of the marine teleost Blennius pholis L. Journal of Fish Biology, 14(5):475–480, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahajan:1979:CTP

[849] C. L. Mahajan and Jagmohan Singh Dheer. Cell types in the peripheral blood of an air-breathing fish Channa punctatus. Journal of Fish Biology, 14(5):481–487, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1979:NAC

[850] C. R. Goldspink and D. Goodwin. A note on the age composition, growth rate and food of perch Perca fluviatilis (L.) in four eutrophic lakes, England. Journal of Fish Biology, 14(5):489–505, May 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ribbink:1979:DET

[851] A. J. Ribbink and B. J. Hill. Depth equilibration by two predatory cichlid fish from Lake Malaŵi. Journal of Fish Biology, 14(6):507–510, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dewan:1979:SPF

[852] S. Dewan, J. U. Miah, A. L. Sarker, and S. N. Saha. Seasonal patterns of feeding of juvenile major carp, Labeo rohita (Ham.) in a Bangladesh pond. Journal of Fish Biology, 14(6):511–515, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rao:1979:SHT

[853] K. S. Rao and N. K. Dad. Studies of herbicide toxicity in some freshwater fishes and ectoprocta. Journal of Fish Biology, 14(6):517–522, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mong:1979:EID

[854] S. J. Mong and Tim M. Berra. The effects of increasing dosages of Xradiation on the chromosomes of the central mudminnow, Umbra limi (Kirtland) (Salmoniformes: Umbridae). Journal of Fish Biology, 14(6): 523–527, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papageorgiou:1979:LWR

[855] N. K. Papageorgiou. The length weight relationship, age, growth and reproduction of the roach Rutilus rutilus (L.) in Lake Volvi. Journal of Fish Biology, 14(6):529–538, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1979:GEP

[856] M. Jobling and P. Spencer Davies. Gastric evacuation in plaice, Pleuronectes platessa L.: effects of temperature and meal size. Journal of Fish Biology, 14(6):539–546, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jafri:1979:OIC

[857] S. I. H. Jafri and D. M. Ensor. Occurrence of an intersex condition in the roach Rutilus rutilus (L). Journal of Fish Biology, 14(6):547–549, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wallace:1979:PAO

[858] R. A. Wallace and K. Selman. Physiological aspects of oogenesis in two species of sticklebacks, Gasterosteus aculeatus L. and Apeltes quadracus (Mitchill). Journal of Fish Biology, 14(6):551–564, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofer:1979:ADEa

[859] R. Hofer. The adaptation of digestive enzymes to temperature, season and diet in roach, Rutilus rutilus L. and rudd Scardinius erythrophthalmus L.

1. Amylase. Journal of Fish Biology, 14(6):565–572, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beumer:1979:FMA

[860] J. P. Beumer. Feeding and movement of Anguilla australis and A. reinhardtii in Macleods Morass, Victoria, Australia. Journal of Fish Biology, 14(6):573–592, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jilek:1979:HDP

[861] Reid Jilek. Histopathology due to the presence of Gracilisentis gracilisentis in Dorosoma cepedianum (Le Sueur). Journal of Fish Biology, 14(6): 593–595, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pasanen:1979:SCM

[862] S. Pasanen, M. Viljanen, and E. Pulkkinen. Stress caused by the ‘markrecapture’ method to Coregonus albula (L). Journal of Fish Biology, 14(6): 597–605, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1979:ACF

[863] R. J. Wootton and L. A. Mills. Annual cycle in female minnows Phoxinus phoxinus (L.) from an upland Welsh lake. Journal of Fish Biology, 14(6): 607–618, June 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1979:SSS

[864] J. S. Alabaster, D. G. Shurben, and M. J. Mallett. The survival of smolts of salmon Salmo salar L. at low concentrations of dissolved oxygen. Journal of Fish Biology, 15(1):1–8, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:1979:BYG

[865] S. S. De Silva and E. I. L. Silva. Biology of young grey mullet, Mugil cephalus L., populatioins in a coastal lagoon in Sri Lanka. Journal of Fish Biology, 15(1):9–20, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babiker:1979:SBRb

[866] M. M. Babiker and H. Ibrahim. Studies on the biology of reproduction in the cichlid Tilapia nilotica (L.): effects of steroid and trophic hormones on ovulation and ovarian hydration. Journal of Fish Biology, 15(1):21– 30, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jilek:1979:SPD

[867] R. Jilek, B. Cassell, D. Peace, Y. Garza, L. Riley, and T. Siewart. Spawning population dynamics of smelt, Osmerus mordax. Journal of Fish Biology, 15(1):31–35, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:1979:BFF

[868] Stephen J. M. Blaber. The biology of filter feeding teleosts in Lake St Lucia, Zululand. Journal of Fish Biology, 15(1):37–59, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagua:1979:OFF

[869] V. O. Sagua. Observations on the food and feeding habits of the African electric catfish Malapterurus electricus (Gmelin). Journal of Fish Biology, 15(1):61–69, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dahlgren:1979:EPD

[870] B. T. Dahlgren. The effects of population density on fecundity and fertility in the guppy, Poecilia reticulata (Peters). Journal of Fish Biology, 15(1): 71–91, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burrough:1979:ONA

[871] R. J. Burrough and C. R. Kennedy. The occurrence and natural alleviation of stunting in a population of roach, Rutilus rutilus (L.). Journal of Fish Biology, 15(1):93–109, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beumer:1979:RCT

[872] J. P. Beumer. Reproductive cycles of two Australian freshwater fishes: the spangled perch, Therapon unicolor Günther, 1859 and the East Queensland rainbowfish, Nematocentris splendida Peters, 1866. Journal of Fish Biology, 15(1):111–134, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strasdine:1979:SAL

[873] G. A. Strasdine and J. R. McBride. Serum antibiotic levels in adult sockeye salmon as a function of route of administration. Journal of Fish Biology, 15(1):135–140, July 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Warner:1979:VSS

[874] M. C. Warner, A. M. Tomb, and Sally A. Diehl. Variability and stability of selected components in rainbow trout Salmo gairdneri serum and the precision of automated analysis in measuring these components. Journal of Fish Biology, 15(2):141–151, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rajagopal:1979:EDT

[875] P. K. Rajagopal. The embryonic development and the thermal effects on the development of the mountain whitefish, Prosopium williamsoni (Girard). Journal of Fish Biology, 15(2):153–158, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Michaelson:1979:OFG

[876] D. M. Michaelson, D. Sternberg, and L. Flshelson. Observations on feeding, growth and electric discharge of newborn Torpedo ocellata (Chondrichthyes, Batoidei). Journal of Fish Biology, 15(2):159–163, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chan:1979:FFH

[877] E. H. Chan and T. E. Chua. The food and feeding habits of greenback grey mullet, Liza subviridis (Valenciennes), from different habitats and at various stages of growth. Journal of Fish Biology, 15(2):165–171, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fouda:1979:SSS

[878] M. M. Fouda. Studies on scale structure in the common goby Pomatoschistus microps Krøyer. Journal of Fish Biology, 15(2):173–183, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Czihak:1979:SDB

[879] G. Czihak, R. Peter, B. Puschendorf, and H. Grunicke. Some data on the basic metabolism of trout eggs. Journal of Fish Biology, 15(2):185– 193, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrews:1979:HSP

[880] C. Andrews. Host specificity of the parasite fauna of perch Perca fluviatilis L. from the British Isles, with special reference to a study at Llyn Tegid (Wales). Journal of Fish Biology, 15(2):195–209, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milner:1979:RTP

[881] N. J. Milner, A. S. Gee, and R. J. Hemsworth. Recruitment and turnover of populations of brown trout, Salmo trutta, in the upper River Wye, Wales. Journal of Fish Biology, 15(2):211–222, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1979:DBC

[882] C. R. Kennedy. The distribution and biology of the cestode Eubothrium parvum in capelin, Mallotus villosus, (Pallas) in the Barents Sea, and its use as a biological tag. Journal of Fish Biology, 15(2):223–236, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ONeill:1979:IRB

[883] J. G. O’Neill. The immune response of the brown trout, Salmo trutta, L. to MS2 bacteriophage: immunogen concentration and adjuvants. Journal of Fish Biology, 15(2):237–248, August 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sivarajah:1979:SSH

[884] K. Sivarajah, C. S. Franklin, and W. P. Williams. Some studies on the hepatic microsomal enzyme activities and steroid hormone levels in carp, Cyprinm carpio exposed for six months. Journal of Fish Biology, 15(3):

249–253, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Went:1979:FTS

[885] A. E. J. Went. ‘Ferox’ trout, Salmo trutta L. of Loughs Mask and Corrib. Journal of Fish Biology, 15(3):255–262, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fouda:1979:APA

[886] M. M. Fouda and P. J. Miller. Alkaline phosphatase activity in the skin of the common goby, Pomatoschistus microps (Krøyer), in relation to cycles in scale and body growth. Journal of Fish Biology, 15(3):263–273, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linfield:1979:CRG

[887] R. S. J. Linfield. Changes in the rate of growth in a stunted roach Rutilu rutilus population. Journal of Fish Biology, 15(3):275–298, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilkins:1979:TVS

[888] N. P. Wilkins and S. Jancsar. Temporal variations in the skin of Atlantic salmon Salmo solar L. Journal of Fish Biology, 15(3):299–307, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henry:1979:ISR

[889] M. G. Henry and G. J. Atchison. Influence of social rank on the behaviour of bluegill, Lepomis macrochims Rafinesque exposed to sublethal concentrations of cadmium and zinc. Journal of Fish Biology, 15(3):309–315, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:1979:IAP

[890] D. P. Anderson, B. S. Roberson, and O. W. Dixon. Induction of antibodyproducing cells in rainbow trout, Salmo gairdneri Richardson; by flush exposure. Journal of Fish Biology, 15(3):317–322, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mrakovcic:1979:IDZ

[891] M. Mrakovčič and L. E. Haley. Inbreeding depression in the Zebra fish Brachydanio rerio (Hamilton Buchanan). Journal of Fish Biology, 15(3): 323–327, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mathur:1979:EPC

[892] Reena Mathur. The effect of porcine calcitonin on plasma calcium in Channa punctatus Bloch. Journal of Fish Biology, 15(3):329–335, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kilambi:1979:ETS

[893] R. V. Kilambi and W. R. Robison. Effects of temperature and stocking density on food consumption and growth of grass carp Ctenopharyngodon idella, Val. Journal of Fish Biology, 15(3):337–342, September 1979. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bubien:1979:REP

[894] J. K. Bubien and T. L. Meade. Relationship of environmental pCO_{2} to selected blood parameters of Oncorhynchus kisutch (Walbaum). Journal of Fish Biology, 15(3):343–347, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stagey:1979:SGI

[895] N. E. Stagey, A. F. Cook, and R. E. Peter. Spontaneous and gonadotropininduced ovulation in the goldfish, Carassius auratus l: effects of external factors. Journal of Fish Biology, 15(3):349–361, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mattey:1979:DDR

[896] D. L. Mattey, M. Morgan, and D. E. Wright. Distribution and development of rodlet cells in the gills and pseudobranch of the bass, Dicentrachus labrax (L). Journal of Fish Biology, 15(3):363–370, September 1979. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Libey:1979:FHY

[897] George S. Libey and C. F. Cole. Food habits of yellowtail flounder, Limandaferruginea (Storer). Journal of Fish Biology, 15(3):371–372, September 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofer:1979:ADEb

[898] R. Hofer. The adaptation of digestive enzymes to temperature, season and diet in roach, Rutilus rutilus and rudd Scardinius erythrophthalmus; proteases. Journal of Fish Biology, 15(4):373–379, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1979:SEM

[899] S. V. Lewis. A scanning electron microscope study of the gills of the air-breathing catfish, Clarias batrachus L. Journal of Fish Biology, 15 (4):381–384, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benda:1979:ACD

[900] R. S. Benda. Analysis of catch data from 1968 to 1976 from nine fish landings in the Kenya waters of Lake Victoria. Journal of Fish Biology, 15(4):385–387, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kratt:1979:ABA

[901] L. F. Kratt and R. J. F. Smith. Agonistic behaviour of age 0, age 1 and non-breeding adult Arctic grayling Thymallus arcticus (Pallas). Journal of Fish Biology, 15(4):389–404, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arillo:1979:GSA

[902] A. Arillo, C. Margiocco, and F. Melodia. The gill sialic acid content as an index of environmental stress in rainbow trout, Salmo gairdneri, Richardson. Journal of Fish Biology, 15(4):405–410, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niederholzer:1979:ADE

[903] R. Niederholzer and R. Hofer. The adaptation of digestive enzymes to temperature, season and diet in roach Rutilus rutilus L. and rudd Scardinius erythrophthalmus L. Cellulase. Journal of Fish Biology, 15(4):411– 416, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freadman:1979:RPS

[904] M. A. Freadman. Role partitioning of swimming musculature of striped bass Morone saxatilis Walbaum and bluefish, Pomatomus saltatrix L. Journal of Fish Biology, 15(4):417–423, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lane:1979:PCH

[905] H. C. Lane. Progressive changes in haematology and tissue water of sexually mature trout, Salmo gairdneri Richardson during the autumn and winter. Journal of Fish Biology, 15(4):425–436, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vandenBroek:1979:ABS

[906] W. L. F. van den Broek. Aspects of the biology of Sprattus sprattm (L.) in the Medway Estuary. Journal of Fish Biology, 15(4):437–447, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buddington:1979:DAM

[907] R. K. Buddington. Digestion of an aquatic macrophyte by Tilapia zillii (Gervais). Journal of Fish Biology, 15(4):449–455, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barham:1979:OAS

[908] W. T. Barham, H. Schoonbee, and G. L. Smit. The occurrence of Aeromonas and Streptococcus in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 15(4):457–460, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Otto:1979:EPE

[909] C. Otto. The effects on a pike (Esox lucius L.) population of intensive fishing in a South Swedish lake. Journal of Fish Biology, 15(4):461–468, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirji:1979:PSCa

[910] K. N. Hirji and W. A. M. Courtney. ‘pear-shaped’ cells in the digestive tract of the perch percafluviatilis (L.). Journal of Fish Biology, 15(4):469– 472, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1979:PDG

[911] C. R. Goldspink. The population density, growth rate and production of roach rutilus mtilus (L.) in Tjeukemeer, The Netherlands. Journal of Fish Biology, 15(4):473–498, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadler:1979:ETG

[912] K. Sadler. Effects of temperature on the growth and survival of the European eel, Anguilla anguilla L. Journal of Fish Biology, 15(4):499–507, October 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Batty:1979:RGL

[913] R. S. Batty and C. S. Wardle. Restoration of glycogen from lactic acid in the anaerobic swimming muscle of plaice, Pleuronectes platessa L. Journal of Fish Biology, 15(5):509–519, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fltzmaurice:1979:SPC

[914] P. Fltzmaurice. Selective predation on Cladocera by brown trout Salmo trutta L. Journal of Fish Biology, 15(5):521–525, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fryer:1979:GHB

[915] J. N. Fryer and H. A. Bern. Growth hormone binding to tissues of normal and stunted juvenile Coho salmon, Oncorhynchus kisutch. Journal of Fish Biology, 15(5):527–533, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burrough:1979:IDA

[916] R. J. Burrough, P. R. Bregazzi, and C. R. Kennedy. Interspecific dominance amongst three species of coarse fish in Slapton Ley, Devon. Journal of Fish Biology, 15(5):535–544, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adeney:1979:SLE

[917] R. J. Adeney and P. Matthiessen. Sub-lethal effects of the organophosphate blackfly larvicide, Abate (Temephos) on Sarotherodon mossambicus (Peters). Journal of Fish Biology, 15(5):545–553, November 1979. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1979:ABS

[918] R. N. Gibson and I. A. Ezzi. Aspects of the biology of the spotted dragonet Callionymus maculatus Rafinesque–Schmaltz from the west coast of Scotland. Journal of Fish Biology, 15(5):555–569, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

North:1979:ABJ

[919] E. North. Aggressive behaviour of juvenile brown trout Salmo trutta, l: an analysis of the wigwag display. Journal of Fish Biology, 15(5):571–577, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watson:1979:MPW

[920] R. A. Watson and T. A. Dick. Metazoan parasites of whitefish Coregonus clupeaformis (Mitchill) and cisco C. artedii Lesueur from Southern Indian Lake, Manitoba. Journal of Fish Biology, 15(5):579–587, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:1979:VFC

[921] D. T. Burton. Ventilation frequency compensation responses of three eurythermal estuarine fish exposed to moderate temperature increases. Journal of Fish Biology, 15(5):589–600, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrison:1979:DCE

[922] C. M. Morrison. A dense cell in the epithelium of the gill lamellae of the brook trout, Salvelinus fontinalis (Mitchill). Journal of Fish Biology, 15 (5):601–605, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jarrams:1979:EFS

[923] Peter Jarrams. Egg, fry and smolt production from salmon, Salmo salar L. and sea trout Salmo trutta L. reared entirely in fresh water. Journal of Fish Biology, 15(5):607–611, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahajan:1979:VCD

[924] C. L. Mahajan and N. K. Agrawal. Vitamin C deficiency in Channa punctatus Bloch. Journal of Fish Biology, 15(5):613–622, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1979:BRa

[925] Anonymous. Book reviews. Journal of Fish Biology, 15(5):623–628, November 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirji:1979:PSCb

[926] K. N. Hirji and W. A. M. Courtney. ‘pear-shaped’ cells in the digestive tract of the perch Perca fluviatilis (L.). Journal of Fish Biology, 15(6): 629–632, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smit:1979:HAEa

[927] G. L. Smit, J. Hattingh, and A. P. Burger. Haematological assessment of the effects of the anaesthetic MS 222 in natural and neutralized form in three freshwater fish species: interspecies differences. Journal of Fish Biology, 15(6):633–643, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smit:1979:HAEb

[928] G. L. Smit, J. Hattingh, and A. P. Burger. Haematological assessment of the effects of the anaesthetic MS 222 in natural and neutralized form in three freshwater fish species: intraspecies differences. Journal of Fish Biology, 15(6):645–653, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smit:1979:HAEc

[929] G. L. Smit, J. Hattingh, and A. P. Burger. Haematological assessment of the effects of the anaesthetic MS 222 in natural and neutralized form in three fresh water fish species: haemoglobin electrophoresis, ATP levels and corpuscular fragility curves. Journal of Fish Biology, 15(6):655–663, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zohar:1979:SKG

[930] Y. Zohar and H. Gordin. Spawning kinetics in the gilthead sea-bream, Sparus aurata L. after low doses of human chronic gonadotropin. Journal of Fish Biology, 15(6):665–670, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvajal:1979:LAO

[931] J. Carvajal, P. E. Cattan, C. Castillo, and P. Schatte. Larval anisakids and other helminths in the hake, Merluccius gayi (Guichenot) from Chile. Journal of Fish Biology, 15(6):671–677, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brill:1979:RWM

[932] R. W. Brill and A. E. Dizon. Red and white muscle fibre activity in swimming skipjack tuna, Katsuwonus pelamis (L.). Journal of Fish Biology, 15 (6):679–685, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sumpter:1979:ARC

[933] J. P. Sumpter and J. M. Dodd. The annual reproductive cycle of the female lesser spotted dogfish, Scyliorhinus canicula L., and its endocrine control. Journal of Fish Biology, 15(6):687–695, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hislop:1979:PON

[934] J. R. G. Hislop. Preliminary observations on the near-surface fish fauna of the northern North Sea in late autumn. Journal of Fish Biology, 15 (6):697–704, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1979:EDO

[935] J. S. Alabaster, D. G. Shurben, and G. Knowles. The effect of dissolved oxygen and salinity on the toxicity of ammonia to smolts of salmon, Salmo salar L. Journal of Fish Biology, 15(6):705–712, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1979:BRb

[936] Anonymous. Book review. Journal of Fish Biology, 15(6):713, December 1979. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).